### v1

# Long-time relative error analysis for linear ODEs with perturbed initial value

Abstract.
We investigate the propagation of initial value perturbations along the solution of a linear ODE $y^{\prime}(t)=Ay(t)$. This propagation is analized using the relative error rather than the absolute error. Our focus is on the long-term behavior of this relative error, which differs significantly from that of the absolute error. Understanding this long-term behavior provides insights into the growth of the relative error over all times, not just at large times. Therefore, it represents a crucial and fundamental aspect of the conditioning of linear ODEs, with applications in, for example, non-normal dynamics. The author hopes that this paper will stimulate attention to the role of relative error in dynamic contexts.
Keywords: linear ordinary differential equations, matrix exponential, relative error, asymptotic behavior, condition numbers, non-normal dynamics.
MSC2020 classification: 15A12, 15A16, 15A18, 15A21, 34A30, 34D05.
1. Introduction
Consider a linear $n-$dimensional Ordinary Differential Equation (ODE)
$\left\{\begin{array}[]{l}y^{\prime}\left(t\right)=Ay\left(t\right),\ t\in\mathbb{R},\\ y\left(0\right)=y_{0},\end{array}\right.$ (1)
where $A\in\mathbb{C}^{n\times n}$ and $y_{0},y(t)\in\mathbb{C}^{n}$. In this paper, we are interested to understand how a perturbation of the initial value $y_{0}$ is propagated along the solution $y(t)=\mathrm{e}^{tA}y_{0}$ of (1) over a long time $t$.
In the following, when a value $x$ is perturbed to $\widetilde{x}$, $\widetilde{x}$ is called the perturbed value and the difference $\widetilde{x}-x$ is called the perturbation of $x$. Moreover, we often speak about a generic situation for an element $v$ of a finite-dimensional space $V$, which refers to a condition on $v$ such that if it does not hold, then $v$ belongs to a manifold $M$ of $V$ with $\mathrm{dim}M<\mathrm{dim}V$. In other words, a generic situation for $v$ holds with probability $1$, if $v$ is randomly selected from $V$.
The next fact A) is well known.
  • A)
    In a generic situation for the perturbation of $y_{0}$ (namely, the perturbation has a nonzero projection onto the rightmost generalized eigenspaces, i.e. those associated to the rightmost eigenvalues of $A$), the perturbation of $y(t)$ asymptotically (as $t\rightarrow+\infty$) vanishes exponentially if the rightmost eigenvalues of $A$ have negative real part and asymptotically diverges exponentially if such eigenvalues have positive real part, without mentioning the case of zero real part.
Thus, if we are concerned about how large can become in the long time the perturbation of the solution, we should be reassured by knowing that the rightmost eigenvalues of $A$ have negative real part and truly concerned by knowing that these eigenvalues have positive real part.
However, in a generic situation for $y_{0}$ (namely, $y_{0}$ has a nonzero projection onto the rightmost generalized eigenspaces), also $y(t)$ asymptotically vanishes exponentially if the rightmost eigenvalues of $A$ have negative real part and asymptotically diverges exponentially if such eigenvalues have positive real part. Thus, to know that the perturbation of the solution asymptotically vanishes, or diverges, exponentially does not help to understand if it is really significant, when compared to the solution.
Aim of the present paper is to compare the perturbation of the solution to the solution itself by considering the relative error of the perturbed solution. The well-know fact A) is a description of the long-time behavior of the absolute error of the perturbed solution. In the present paper, we will give a description of the long-time behavior of the relative error. This description, as one can image, is quite different from A).
In our analysis of the relative error of the perturbed solution, we have in mind real-world systems described by mathematical models based on linear ODEs (1), which require simulation through the integration of these ODEs. In this context, if there is uncertainty in the initial value (for example due to measurement errors), it is important not only to know the long-time behavior of the absolute error of perturbed solutions as described in A), but also to understand the long-time behavior of the relative error. After all, in applied sciences, the relative error is often considered the most significant type of error, and the entire framework of computing relies on concepts related to the relative error, such as floating-point numbers and significant figures. In Subsection 1.2 below, we examine two such models and emphasize the importance of considering the relative error of the solution in cases of uncertainty in the initial value.
Moreover, from a purely mathematical perspective, even without considering potential applications (which do exist), studying the long-time behavior of the relative error of perturbed solutions can be a mathematical issue worthy of investigation.
Put differently, in this paper we are interested in the (relative) conditioning of the linear problem
$y_{0}\mapsto\mathrm{e}^{tA}y_{0}$ (2)
for large $t$, i.e. we are interested to study how a relative error in the input $y_{0}$ is propagated to the output $\mathrm{e}^{tA}y_{0}$ for large $t$. The paper gives a description of this long-time conditioning in terms of eigenvalues and generalized eigenvectors of the matrix $A$.
1.1. Literature
In the literature, the conditioning of the matrix exponential $\mathrm{e}^{tA}$ with respect to perturbations of the matrix $A$, namely the conditioning of the problem $A\mapsto\mathrm{e}^{tA},$ is a well studied topic (see e.g. [15], [21], [14], [16], [20], [2], [24] and [1]). On the other hand, the conditioning of the action of the matrix exponential $\mathrm{e}^{tA}$ on a vector with respect to perturbations of this vector, i.e. the conditioning of the problem (2), has received little attention. The papers [3] and [8] considered the conditioning of the problem $(A,y_{0})\mapsto\mathrm{e}^{tA}y_{0},$ which includes the problem (2). However, these papers were focused on computational aspects and algorithms and they did not study the conditioning from a qualitative point of view. In particular, how this conditioning depends on $t$, an issue of interest in the context of ODEs.
A study of how the conditioning of the problem (2) depends on $t$ was given in the paper [18], but the analysis was confined to the case of a normal matrix $A$ and normwise relative errors in the Euclidean norm. The present paper extends this study to a general matrix and a general norm. The conditioning of the problem $A\mapsto\mathrm{e}^{tA}y_{0}$, which is different from the problem $A\mapsto\mathrm{e}^{tA}$ since $y_{0}$ is involved, was considered in [12], for $A$ normal.
1.2. Two models
The analysis of the relative error of the perturbed solution given in the present paper can be of interest when we have a mathematical model based on a linear ODE (1) and an accurate determination of the solution $y$ is required under uncertainty in the initial value $y_{0}$. This analysis is particularly important when the solution becomes small or large, compared to the initial value, a situation where looking to the absolute error of the perturbed solution can have little significance.
Although the analysis of the propagation of perturbations in the matrix $A$, rather than in the initial value $y_{0}$, is generally considered more important, since such perturbations can lead to structural changes (e.g. a stable system becoming unstable), it is worth noting that, in many cases, the model (1) is constructed after a careful selection of the parameters appearing as entries in $A$ by the model creators. As a result, the matrix $A$ can be regarded as fixed and reliable. The model is then used repeatedly with different initial conditions, often by external users who can have little control over the accuracy of these initial conditions. In this context, where the initial conditions are inputs to a fixed model, the analysis developed here proves to be particularly useful.
We show the importance of understanding the behavior of the relative error of the perturbed solution through two mathematical models based on a linear ODE (1).
1.2.1. Gross Domestic Product and National Debt
The papers [9, 10] present a mathematical model for the Gross Domestic Product (GDP) and National Debt (ND) of a given country. This model is based on the linear system of ODEs
$\left\{\begin{array}[]{l}Q^{\prime}\left(t\right)=a_{11}Q\left(t\right)+a_{12}B\left(t\right)\\ B^{\prime}\left(t\right)=a_{21}Q\left(t\right)+a_{22}B\left(t\right),\end{array}\right.$ (3)
where $Q(t)$ is the GDP and $B(t)$ is the ND at the time $t$. The unit for time is $1\ \mathrm{yr}$ and the unit for GDP and ND is the initial GDP, i.e. $Q(0)=1$. A possible instance for the coefficients (given in [9]) is
$a_{11}=0.08,\ a_{12}=-0.07,\ a_{21}=0.03,\ a_{22}=-0.02.$ (4)
For this instance, the matrix $A$ of the ODE has the positive eigenvalues $0.05$ and $0.01$.
Suppose we are interested in simulating the growth of GDP and ND over a period of $50\ \mathrm{yr}$ by integrating (3) for a given initial ND (remind that the initial GDP is set at $1$). We assume, as in [9], that the initial ND is $0.60$. Due to uncertainty in the available data, this initial ND may not be the actual value. Suppose the actual initial ND is $0.61$. We set $B(0)$ to this actual initial ND of $0.61$, while our assumed initial value of $0.60$ for the simulation is the perturbed value $\widetilde{B}(0)$. (Conversely, we could set $B(0)=0.60$ and $\widetilde{B}(0)=0.61$. See Remark 14).
Since the eigenvalues of $A$ are positive, the perturbation of the initial ND grows exponentially in the solution. After $50\ \mathrm{yr}$, for $(Q(0),B(0))=(1,0.61)$, we have
$(Q(50),B(50))=(8.84,4.09)$
and, for the perturbed $(\widetilde{Q}(0),\widetilde{B}(0))=(1,0.60)$, we have
$(\widetilde{Q}(50),\widetilde{B}(50))=(9.02,4.15).$
Here, we denote the perturbed solution of (3) by $(\widetilde{Q}(t),\widetilde{B}(t))$.
Now, the question is:
  • after $50\ \mathrm{yr}$, is the effect of the initial uncertainty significant?
There two ways to answer this.
  • 1)
    The perturbation of the initial ND is $1$ second decimal figure and the perturbation of GDP and ND after $50\ \mathrm{yr}$ is $18$ second decimal figure and $6$ second decimal figure, respectively. This demonstrates a significant increase in the perturbation, as an effect of exponential growth. In Figure 1, we see in the plane $(Q,B)$ the points $(Q(0),B(0))$ and $(\widetilde{Q}(0),\widetilde{B}(0))$ on the left and the points $(Q(50),B(50))$ and $(\widetilde{Q}(50),\widetilde{B}(50))$ on the right. On the left and on the right, we use the same scale with both axes of length 1 GDP unit.
    Figure 1. Left: points $(Q(0),B(0))$ and $(\widetilde{Q}(0),\widetilde{B}(0))$. Right: points $(Q(50),B(50))$ and $(\widetilde{Q}(50),\widetilde{B}(50))$.
    The growth of the perturbation is evident. In Figure 2, we see the absolute error of the perturbed solution
    $e(t)=\left\|(\widetilde{Q}\left(t\right),\widetilde{B}(t))-\left(Q\left(t\right),B(t)\right)\right\|_{2},$ (5)
    i.e. the Euclidean distance between the points $(\widetilde{Q}\left(t\right),\widetilde{B}(t))$ and $\left(Q\left(t\right),B(t)\right)$, for $t\in[0,50]$.
    Figure 2. Absolute error $e(t)$ for the GDP-ND model.
    After $50\ \mathrm{yr}$, the absolute error is about $20$ times the initial absolute error.
  • 2)
    The perturbation of the initial ND is $1$ second significant figure and the perturbation of GDP and ND after $50\ \mathrm{yr}$ is $1.8$ second significant figure and $0.6$ second significant figure, respectively. This shows that the perturbed GDP and ND are still close to the unperturbed values after $50\ \mathrm{yr}$. In Figure 3, similarly to Figure 1, we see the points $(Q(0),B(0))$ and $(\widetilde{Q}(0),\widetilde{B}(0))$ on the left and the points $(Q(50),B(50))$ and $(\widetilde{Q}(50),\widetilde{B}(50))$ on the right, but now on the left we use a scale with axes of length $\left\|(Q(0),B(0))\right\|_{2}$ and on the right of length $\left\|(Q(50),B(50))\right\|_{2}$.
    Figure 3. Left: points $(Q(0),B(0))$ and $(\widetilde{Q}(0),\widetilde{B}(0))$. Right: points $(Q(50),B(50))$ and $(\widetilde{Q}(50),\widetilde{B}(50))$.
    The closeness of $(Q(50),B(50))$ and $(\widetilde{Q}(50),\widetilde{B}(50))$ is comparable with the closeness of $(Q(0),B(0))$ and $(\widetilde{Q}(0),\widetilde{B}(0))$. This closeness is also confirmed by Figure 4, where it appears that GPD and perturbed GDP, as well as ND and perturbed ND, plotted as functions of time over $50$ $\mathrm{yr}$ are barely distinguishable.
    Figure 4. GDP $Q(t)$, perturbed GDP $\widetilde{Q}(t)$, ND $B(t)$ and perturbed ND $\widetilde{B}(t)$.
    In Figure 5, we see the relative error of the perturbed solution
    $\delta(t)=\frac{\left\|(\widetilde{Q}\left(t\right),\widetilde{B}(t))-\left(Q\left(t\right),B(t)\right)\right\|_{2}}{\left\|\left(Q\left(t\right),B(t)\right)\right\|_{2}}$ (6)
    for $t\in[0,50]$. It is the distance between the points $(\widetilde{Q}\left(t\right),\widetilde{B}(t))$ and $\left(Q\left(t\right),B(t)\right)$ when, as in Figure 3, we use the length $\left\|(Q(t),B(t))\right\|_{2}$ of the axes as unit length.
    Figure 5. Relative error $\delta(t)$ for the GDP-ND model.
    Unlike the exponentially growing absolute error (5), the relative error (6) grows much less. After $50\ \mathrm{yr}$, it is about $2$ times the initial relative error.
If one believes that answer 2) is more appropriate (meaning that after 50 years, the effect of the initial perturbation is not significant because the perturbation of the solution is not much larger than its initial value, when compared to the solution), it becomes crucial to understand the behavior of the relative error (6) rather than that of the absolute error (5).
The relative error (6) refers to the relative error of the solution vector $(Q(t),B(t))$, i.e. it is a normwise relative error. One might find it more interesting to consider the relative errors of the two components $Q(t)$ and $B(t)$. In the specific integration over $50$ $\mathrm{yr}$ we are considering, the relative errors of the components and the relative error (6) have the same order of magnitude, since the components of the solution are not small compared to the Euclidean norm of the solution: we see in Figure 3 that the point $(Q(50),B(50))$ is not close to the axes. It is noteworthy that in Figure 4, $Q(t)$ and $\widetilde{Q}(t)$, as well as $B(t)$ and $\widetilde{B}(t)$, are nearly indistinguishable due to small componentwise relative errors.
About componentwise relative errors, see Subsection 2.1 below.
This mathematical model of GDP and ND will be revisited in Section 7 in light of the results presented in this paper.
1.2.2. Building heating
This example (see the book [13]) is a toy mathematical model of building heating. Similar more complex non-toy models are used in literature for modeling real building heating (see e.g. [17], [7] and [23]).
Consider a building constituted by basement, main floor and attic. Let $x_{1}(t)$, $x_{2}(t)$ and $x_{3}(t)$ be the temperatures of basement, main floor and attic, respectively, at the time $t$. By using the Newton’s law of cooling
rate of change of internal temperature
[…]
we derive the linear system of ODEs
$\left\{\begin{array}[]{l}x_{1}^{\prime}\left(t\right)=k_{g1}\left(x_{g}-x_{1}\left(t\right)\right)+k_{12}\left(x_{2}\left(t\right)-x_{1}\left(t\right)\right)+f_{1}\\ x_{2}^{\prime}\left(t\right)=k_{a2}\left(x_{a}-x_{2}\left(t\right)\right)+k_{12}\left(x_{1}\left(t\right)-x_{2}\left(t\right)\right)+k_{23}\left(x_{3}\left(t\right)-x_{2}\left(t\right)\right)+f_{2}\\ x_{3}^{\prime}\left(t\right)=k_{a3}\left(x_{a}-x_{3}\left(t\right)\right)+k_{23}\left(x_{2}\left(t\right)-x_{3}\left(t\right)\right)+f_{3},\end{array}\right.$ (7)
where
  • $x_{g}$ is temperature of the ground and $x_{a}$ is the outdoor temperature of the air;
  • $k_{g1},k_{a2},k_{a3},k_{12},k_{23}$ are positive proportionality constants in the Newton’s law depending on the thermal insulation of the floors;
  • $f_{1}$, $f_{2}$ and $f_{3}$ are constant forcing terms due to heaters in the basement, main floor and attic, respectively.
The ODE (7) has the form
$x^{\prime}(t)=Ax(t)+b,$ (8)
where
$A=\left[\begin{array}[]{ccc}-k_{g1}-k_{12}&k_{12}&0\\ k_{12}&-k_{a2}-k_{12}-k_{23}&k_{23}\\ 0&k_{23}&-k_{a3}-k_{23}\end{array}\right]\text{\ \ and\ \ }b=\left[\begin{array}[]{ccc}k_{g1}x_{g}+f_{1}\\ k_{a2}x_{a}+f_{2}\\ k_{a3}x_{a}+f_{3}\end{array}\right].$
Since the matrix $A$ is symmetric and negative diagonally dominant, it has real negative eigenvalues. Consequently, the ODE (8) has the globally asymptotically stable equilibrium point
$x_{eq}=-A^{-1}b.$
The transient
$y(t)=x(t)-x_{eq}$
satisfies the ODE
$y^{\prime}(t)=Ay(t).$ (9)
Suppose we want to simulate the transient by integrating (9), and there is uncertainty in the initial temperatures $y_{0}$ at time $t=0$. Of course, $y_{0}$ and $y(t)$ are temperatures with respect to the equilbrium temperatures.
Therefore, we have an initial value $y_{0}$, representing the actual initial temperatures, and a perturbed initial value $\widetilde{y}_{0}$, representing the initial temperatures available to us for simulation. (Alternatively, we could assume that $y_{0}$ represents the available temperatures and $\widetilde{y}_{0}$ represents the actual temperatures. See Remark 14).
For illustrating our considerations, in (7) set
$k_{g1}=0.5\ \frac{\frac{{}^{\circ}\mathrm{C}}{\mathrm{h}}}{{}^{\circ}\mathrm{C}},\ k_{a2}=0.25\ \frac{\frac{{}^{\circ}\mathrm{C}}{\mathrm{h}}}{{}^{\circ}\mathrm{C}},k_{a3}=0.25\ \frac{\frac{{}^{\circ}\mathrm{C}}{\mathrm{h}}}{{}^{\circ}\mathrm{C}},\ k_{12}=0.5\ \frac{\frac{{}^{\circ}\mathrm{C}}{\mathrm{h}}}{{}^{\circ}\mathrm{C}},\ k_{23}=1\ \frac{\frac{{}^{\circ}\mathrm{C}}{\mathrm{h}}}{{}^{\circ}\mathrm{C}}$ (10)
(temperatures are measured in Celsius degrees and the time in hours). For this particular instance, the eigenvalues of $A$ are $-0.31519\ \mathrm{h}^{-1}$, $-1.0560\ \mathrm{h}^{-1}$ and $-2.6288\ \mathrm{h}^{-1}$. Suppose
$y_{0}=\left(3.5\ ^{\circ}\mathrm{C},-4.4\ ^{\circ}\mathrm{C},2.5\ ^{\circ}\mathrm{C}\right)$ (11)
and
$\widetilde{y}_{0}=\left(4\ ^{\circ}\mathrm{C},-4\ ^{\circ}\mathrm{C},3\ ^{\circ}\mathrm{C}\right).$ (12)
In Figures 6 and 7 we show the absolute error
$e(t)=\left\|\widetilde{y}(t)-y(t)\right\|_{2}$
and the relative error
$\delta(t)=\frac{\left\|\widetilde{y}(t)-y(t)\right\|_{2}}{\left\|y(t)\right\|_{2}},$
where $\widetilde{y}$ is the perturbed (simulated) solution of (9), for $t\in\left[0,6\ \mathrm{h}\right]$.
Figure 6. Absolute error $e(t)$ for the building heating model.
Figure 7. Relative error $\delta(t)$ for the buiding heating model.
Here, the scenario is completely reversed compared to the GDP-ND model. Due to the negative eigenvalues, the absolute error decreases exponentially and the initial perturbation is reduced by about seven times in $6$ hours, reassuring us about the effect of the uncertainty in the initial temperatures when the transient is simulated by integrating (9). However, this is true only when comparing the perturbation to the initial value $y_{0}$, i.e. by looking at the absolute error. When comparing the perturbation to the solution $y(t)$, i.e. by looking at the relative error, the situation is much more concerning as the initial perturbation is magnified by about twelve times. This could significantly impact our understanding of the transient and the decisions we make based on the simulation.
For example, suppose we want to determine when the Euclidean norm of the solution drops below $0.5\ ^{\circ}\mathrm{C}$. When this occurs, we could decide that the transient phase has ended. The simulation with initial value $\widetilde{y}_{0}$ shows that this occurs at time $\widetilde{t}^{\ast}=3.0876\ \mathrm{h}$. However, at this time $\widetilde{t}^{\ast}$, the relative error of the perturbed solution is greater than 1 (see Figure 7), meaning that the norm of the perturbation is larger than the norm of the actual solution. This implies that the norm $\|\widetilde{y}(\widetilde{t}^{\ast})\|_{2}=0.5^{\circ}\mathrm{C}$ of the perturbed solution could be more than twice the norm $\|y(\widetilde{t}^{\ast})\|_{2}$ of the actual solution. Indeed, we have
$\frac{\|\widetilde{y}(\widetilde{t}^{\ast})\|_{2}}{\|y(\widetilde{t}^{\ast})\|_{2}}=2.3882.$
The norm of the actual solution, which is less than half that the norm $0.5^{\circ}\mathrm{C}$ of the perturbed solution at time $\widetilde{t}^{\ast}$, dropped below $0.5^{\circ}\mathrm{C}$ at $t^{\ast}=1.6362\ \mathrm{h}$, in roughly half the time. Therefore, the simulation with the initial temperatures $\widetilde{y}_{0}$ available to us provides a highly inaccurate estimate of the end transient time. This is not only due to poor accuracy in the initial temperatures $\widetilde{y}_{0}$ (the relative error of $\widetilde{y}_{0}$ with respect to $y_{0}$ is $\delta(0)=0.1320$), but also to inherent ill-conditioning (the relative error of $\widetilde{t}^{\ast}$ with respect to $t^{\ast}$ is $0.8871$).
Note that, to understand how the ratio $\frac{\|\widetilde{y}(t)\|_{2}}{\|y(t)\|_{2}}$ varies over time and, consequently, how different $\widetilde{t}^{\ast}$ and ${t}^{\ast}$ can be, it is natural to consider the relative error $\delta(t)$. In fact, we have
$\frac{\|\widetilde{y}(t)\|_{2}}{\|y(t)\|_{2}}=1+\xi(t),$
and, consequently,
$\|y(t^{\ast})\|_{2}=0.5^{\circ}\mathrm{C}\text{\ and\ }\|y(\widetilde{t}^{\ast})\|_{2}=0.5^{\circ}\mathrm{C}\cdot(1+\xi(\widetilde{t}^{\ast})),$
where $\xi(t)$ satisfies $|\xi(t)|\leq\delta(t)$.
The previous considerations remain valid in the case of time-varying temperatures $x_{g}(t)$ and $x_{a}(t)$ and heaters $f_{1}(t)$, $f_{2}(t)$, and $f_{3}(t)$, i.e. $b$ in (8) is time-varying, and there is a particular solution of (8) considered as an attractor.
Similarly to the previous GDP-ND model, we will revisit this building heating model in Section 7.
1.3. The asymptotic behavior of the relative error
The present paper deals with the asymptotic behavior, i.e. the long-time behavior, of the relative error of the perturbed solution.
One could observe that, in case of a solution that decays to zero or diverges, this analysis might have limited relevance since asymptotically the solution is zero or infinite and then a long-time simulation of the solution is not very interesting. However, in such a case we might be interested in simulating the solution up to a time beyond which it has become too small or too large to be of further interest. Therefore, if the relative error becomes close to, or with the same order of magnitude as, its asymptotic behavior before the solution has reached these lower and upper thresholds, then the analysis of the asymptotic behavior has interest.
In the GDP-ND and building heating models, although the solution diverges to infinity or converges to zero, we are still interested in understanding its evolution before it becomes too large or too small. The relative error after 50 years in the GDP-ND model and the relative error at the end transient time $t^{\ast}$ in the building heating model have the same order of magnitude as their asymptotic values (see Figure 15 and Figure 7).
In the present paper, in addition to study the asymptotic behavior of the relative error, we also investigate how rapidly this asymptotic behavior is attained. As expected, the non-normality of the matrix $A$ adversely affects the rapid attainment of the asymptotic behavior. However, as it will be shown, an high non-normality of $A$ does not necessarily imply a late onset of the asymptotic behavior, which could lead to a loss of interest in such behavior.
In conclusion, we can say that although the asymptotic behavior of the relative error does not fully describe the propagation to the solution of the perturbation of the initial value, since it can miss a possible initial growth of the relative error to values much larger than the asymptotic behavior, it nonetheless constitutes an important piece in the qualitative study of the conditioning of the problem (2).
However, Section 8 shows experimentally that it is very rear to have such a large initial grow of the relative error. Therefore, the asymptotic behavior might be not only an important piece of the study of conditioning, but the crucial and fundamental piece.
1.4. Plan of the paper
The asymptotic analysis of the relative error of the perturbed solution is long and the details quite technical. This is due to the fact that we address the general case involving an arbitrary matrix $A$ in (1), a choice that necessitates dealing with the Jordan Canonical Form (JCF) of $A$ and generalized eigenvectors. Moreover, the presence of complex eigenvalues further complicates matters. Indeed, if we restricted ourselves to diagonalizable matrices with only real eigenvalues, the analysis would be considerably shorter. Finally, our definition of asymptoticity also implies some effort in proving asymptotic results.
In light of this, in the present paper we limit ourselves to illustrate only the results of the analysis, avoiding all the development that led to them. This development can be found in [19], a manuscript supplementing the present paper.
To investigate the relative error of the perturbed solution, Section 2 introduces two condition numbers for the problem (2). Sections 3 and 4 describe the asymptotic behavior of these condition numbers and Section 5 explores how rapidly this asymptotic behavior is attained. Section 6 is essentially a summary of the results on the asymptotic behavior of the relative error. In Section 7, the GDP-ND and building heating models are revisited and five additional examples illustrate the results presented in the paper. Section 8 shows that the asymptotic behavior of the relative error can also provide insight into the non-asymptotic behavior in most cases. Section 9 illustrates how the results of the present paper can be applied in non-normal dynamics. Section 10 presents the conclusions and can even be read now to gain a better idea of the contribution of this paper.
Given the length of the paper, it is useful to note that a fundamental understanding of the most important results can be gained by reading only the following parts.
  • Section 2 without 2.1.
  • Section 3: only 3.2 and the last paragraph of 3.3.
  • Section 4: 4.1, 4.2 and, in 4.3, only 4.3.1, Theorem 9 and the preceding paragraph, Theorem 10 and the preceding paragraph.
  • Sections 5 and 6.
  • Section 7: only 7.1, 7.2 and 7.5.
  • Sections 8 and 9.
The paper presents several numerical experiments in which a large number of random instances of the ODE (1) are generated by sampling the entries of $A$ and the components of $y_{0}$ from the standard normal distribution, except in one case where a uniform distribution is used. This means that all entries of $A$ and all components of $y_{0}$ in every instance are independently sampled. It is worth noting that, the entries of $A$ are drawn from the standard normal distribution rather than from the normal distribution with mean zero and standard deviation $\frac{1}{\sqrt{n}}$, which is often employed. We choose not to use the $\frac{1}{\sqrt{n}}$ scaling because we aim to include a broad range of values for the norm of the matrix $A$, thereby generating potentially stiffer or more extreme instances of the ODE. Nevertheless, performing the experiments with the scaled version yields essentially the same results and, consequently, leads to the same conclusions.
All numerical experiments are carried out in MATLAB, with the matrix exponential values $\mathrm{e}^{tA}$ computed using the expm function.
2. The condition numbers
[…]
$\varepsilon=\frac{\left\|\widetilde{y}_{0}-y_{0}\right\|}{\left\|y_{0}\right\|}$ (13)
of the perturbed initial value $\widetilde{y}_{0}$ to the normwise relative error
$\delta\left(t\right)=\frac{\left\|\widetilde{y}\left(t\right)-y\left(t\right)\right\|}{\left\|y\left(t\right)\right\|}$ (14)
of the perturbed solution $\widetilde{y}(t)$. Observe that $\delta(0)=\varepsilon$. Moreover, in the case of the Euclidean norm as vector norm, observe that $\delta(t)$ is the distance between $\widetilde{y}(t)$ and $y(t)$ when $\|y(t)\|_{2}$ is used as scale unit (see Figure 8).
Remark 1.
Observe that the normwise relative error (14) is invariant under multiplication of the vector norm by a constant. Therefore, these relative errors remain the same for both the $p$-norm and the mean $p$-norm,
$\|x\|_{p,\text{mean}}=\left(\sum\limits_{i=1}^{n}\frac{1}{n}|x_{i}|^{p}\right)^{\frac{1}{p}},\quad x\in\mathbb{C}^{n}.$
Using such mean norms avoids the expected growth of the $p$-norms with the dimension $n$.
Figure 8. The relative error $\delta(t)$ when $\|y(t)\|_{2}$ is the scale unit.
By writing the perturbed initial value as
$\widetilde{y}_{0}=y_{0}+\varepsilon\|y_{0}\|\widehat{z}_{0},$
where $\widehat{z}_{0}\in\mathbb{C}^{n}$ is a unit vector (i.e. $\|\widehat{z}_{0}\|=1$) indicating the direction of perturbation, we obtain
$\delta\left(t\right)=K\left(t,y_{0},\widehat{z}_{0}\right)\cdot\varepsilon,$ (15)
where
$K\left(t,y_{0},\widehat{z}_{0}\right):=\frac{\left\|\mathrm{e}^{tA}\widehat{z}_{0}\right\|}{\left\|\mathrm{e}^{tA}\widehat{y}_{0}\right\|}$ (16)
with $\widehat{y}_{0}:=\frac{y_{0}}{\left\|y_{0}\right\|}$ the normalized initial value. We define $K(t,y_{0},\widehat{z}_{0})$ as the directional condition number of the problem (2).
In general, we know nothing about the direction $\widehat{z}_{0}$ of the perturbation of $y_{0}$. Therefore, it is useful to introduce
$K\left(t,y_{0}\right):=\max\limits_{\begin{subarray}{c}\widehat{z}_{0}\in\mathbb{C}^{n}\\ \|\widehat{z}_{0}\|=1\end{subarray}}K\left(t,y_{0},\widehat{z}_{0}\right)=\frac{\left\|\mathrm{e}^{tA}\right\|}{\left\|\mathrm{e}^{tA}\widehat{y}_{0}\right\|},$ (17)
[…]
The paper [18] analyzed the condition numbers (16) and (17) for $A$ normal and the Euclidean norm as vector norm. The present paper looks at the general case where $A$ and the vector norm are arbitrary.
2.1. Normwise and componentwise relative errors
We have introduced the normwise relative errors (13) and (14), but it may also be worthwhile to consider the componentwise relative errors
$\varepsilon_{\ell}=\frac{\left|\widetilde{y}_{0\ell}-y_{0l}\right|}{\left|y_{0\ell}\right|},\ \ell\in\{1,\ldots,n\},$
of the perturbed initial value and
$\delta_{\ell}\left(t\right)=\frac{\left|\widetilde{y}_{\ell}\left(t\right)-y_{\ell}\left(t\right)\right|}{\left|y_{\ell}\left(t\right)\right|},\ \ell\in\{1,\ldots,n\},$
of the perturbed solution, where $y_{0\ell}$, $\widetilde{y}_{0\ell}$, $y_{\ell}(t)$ and $\widetilde{y}_{\ell}(t)$, $\ell\in\{1,\ldots,n\}$, are the components of $y_{0}$, $\widetilde{y}_{0}$, $y(t)$ and $\widetilde{y}(t)$, respectively. Componentwise relative errors were considered in [11], for $A$ diagonalizable.
[…]
  • 1)
    $\varepsilon\leq\max\limits_{\ell\in\{1,\ldots,n\}}\varepsilon_{l}$
  • 2)
    $\delta(t)\leq\max\limits_{\ell\in\{1,\ldots,n\}}\delta_{\ell}\left(t\right)$
  • 3)
    $\delta_{\ell}\left(t\right)\leq\frac{\|y(t)\|}{|y_{\ell}(t)|}\delta(t),\ \ell\in\{1,\ldots,n\}.$
Observe that 1), 2) and 3) imply, respectively:
  • 1’)
    if all the componentwise relative errors $\varepsilon_{\ell}$ are small, then the normwise relative error $\varepsilon$ is small;
  • 2’)
    if the normwise relative error $\delta(t)$ is large, then some componentwise relative error $\delta_{\ell}(t)$ is large;
  • 3’)
    if the normwise relative error $\delta(t)$ is small, then the componentwise relative error $\varepsilon_{\ell}$ is small for components $y_{\ell}(t)$ with a non-large ratio $\frac{\|y(t)\|}{|y_{\ell}(t)|}$.
[…]
$\frac{\|y(t)\|}{|y_{\ell}(t)|}\leq n^{\frac{1}{p}}\frac{\|y(t)\|_{\infty}}{|y_{\ell}(t)|}.$ (18)
Example 2.
Consider the GDP-ND model. We look at 10 0001000010\ 00010 000 instances of (3), where the coefficients $a_{11}$, $-a_{12}$, $a_{21}$ and $-a_{11}$ are sampled from the uniform distribution on the interval $[0.01,0.10]$ and the initial value has $Q(0)=1$ and $B(0)$ sampled from the uniform distribution on $[0,1]$. Figure 9 shows the maximum ratio
$R=\max\left\{\frac{\|(Q(50),B(50))\|_{2}}{|Q(50)|},\frac{\|(Q(50),B(50))\|_{2}}{|B(50)|}\right\}$ (19)
in logarithmic scale for these 10 0001000010\ 00010 000 instances.
Figure 9. Maximum ratio $R$ in (19) for 10 0001000010\ 00010 000 random instances of (3).
There are few instances with a large $R$. The percentage of instances with $R$ larger than $10$, $100$ and $1000$ is $13.43\%$, $1.34\%$ and $0.06\%$, respectively. The median $R$ is $2.63$. Observe that $R\geq\sqrt{2}$.
Example 3.
Figure 10. Maximum ratio $R(t)$ in (20) for 10 0001000010\ 00010 000 random instances of (1) with $n=100$.
In Figure 10, we see for $t=0.1,1,10$ the maximum ratio
$R(t)=\max\limits_{\ell\in\{1,\ldots,n\}}\frac{\|y(t)\|_{\infty}}{|y_{\ell}(t)|}$ (20)
in logarithmic scale, for 10 0001000010\ 00010 000 instances of (1), where $A$ of order $n=100$ and $y_{0}$ have entries sampled from the standard normal distribution. Here the picture is completely different from the previous example, since $R(t)$ is large for almost all instances.
However, in Figure 11, we see, for all 10 0001000010\ 00010 000 instances, the fraction
$r(t,M)=\frac{\text{number of components $y_{\ell}(t)$, $\ell\in\{1,\ldots,n\}$, such that $\frac{\|y(t)\|_{\infty}}{|y_{\ell}(t)|}>M$}}{n}$ (21)
for $M=10$ in the left side and $M=100$ in the right side. For $M=10$, the fraction is less than $50\%$ in all instances and around $25\%$ in the average. For $M=100$, it is less than $10\%$ in all instances and under $5\%$ in the average.
Figure 11. Fraction $r(t,M)$ in (21) for the 10 0001000010\ 00010 000 random instances of Figure 10.
In conclusion, analyzing the normwise relative error can indeed provide valuable insights into the componentwise relative errors. In fact, we have evidence that a substantial percentage of the componentwise relative errors of the solution have an order of magnitude not larger than that of the normwise relative error, when the $\infty$-norm is used as vector norm. For other $p$-norms, consider (18).
3. The asymptotic behavior of the condition numbers
In this section and the next two, the asymptotic (long-time) behavior of the condition numbers (16) and (17) is illustrated. In light of (15), this also illustrates the asymptotic behavior of the normwise relative error $\delta(t)$. The proofs for all results presented here can be found in [19], as previously remarked.
3.1. Notations
We need some notations now introduced.
  • $\Lambda=\{\lambda_{1},\ldots,\lambda_{p}\}$, where $\lambda_{1},\ldots,\lambda_{p}$ are distinct, is the set of the eigenvalues of $A$ (the spectrum of $A$). The imaginary part of $\lambda_{i}\in\Lambda$ is denoted by $\omega_{i}$.
  • We assume to have a Jordan basis
    $v^{(i,j,k)},\ i\in\{1,\ldots,p\},\ j\in\{1,\ldots,d_{i}\}\text{\ and\ }k\in\{1,\ldots,m_{ij}\},$ (22)
    of $\mathbb{C}^{n}$ constituted by $d_{i}$ Jordan chains
    $v^{(i,j,1)},\ldots,v^{(i,j,m_{ij})},\ j\in\{1,\ldots,d_{i}\},$
    for each eigenvalue $\lambda_{i}\in\Lambda$, where $d_{i}$ is the geometric multeplicity of $\lambda_{i}$. The lenghts $m_{ij}$, $j\in\{1,\ldots,d_{i}\}$, of the Jordan chains corresponding to $\lambda_{i}$ are the orders of the $d_{i}$ Jordan mini-blocks corresponding to $\lambda_{i}$ in the JCF of $A$.
  • Let $V$ be the $n\times n$ matrix whose columns are the elements (22) of the Jordan basis, ordered lexicographically by means of the indeces $(i,j,k)$. The rows of the inverse $V^{-1}$ are denoted by
    $w^{(i,j,k)},\ i\in\{1,\ldots,p\},\ j\in\{1,\ldots,d_{i}\}\text{\ and\ }k\in\{1,\ldots,m_{ij}\},$
    also ordered lexicographically by means of the indeces $(i,j,k)$. The columns $v^{(i,j,k)}$ of $V$ and the rows $w^{(i,j,k)}$ of $V^{-1}$ are (right) generalized eigenvectors and left generalized eigenvectors, respectively, of $A$ corresponding to the eigenvalue $\lambda_{i}$. The columns $v^{(i,j,1)}$ and the rows $w^{(i,j,1)}$ are (right) eigenvectors and left eigenvectors, respectively, corresponding to $\lambda_{i}$.
  • For any $u\in\mathbb{C}^{n}$, the components of $u$ in the Jordan basis are denoted by
    $\alpha_{ijk}(u),\ i\in\{1,\ldots,p\},\ j\in\{1,\ldots,d_{i}\}\text{\ and\ }k\in\{1,\ldots,m_{ij}\},$
    where $\alpha_{ijk}(u)$ is the component along the element $v^{(i,j,k)}$ of the Jordan basis.
  • The spectrum $\Lambda$ of $A$ is partitioned by decreasing real parts. Let $r_{1},\ldots,r_{q}$ be the distinct decreasing real parts of the eigenvalues of $A$. For any $j\in\{1,\ldots,q\}$, let $\Lambda_{j}$ be the set of the eigenvalues with real part $r_{j}$ (see Figure 12). Thus, $\Lambda_{1}$ is the set of the rightmost eigenvalues of $A$.
    Figure 12. Partition of the eigenvalues $\lambda_{1},\ldots,\lambda_{p}$ of $A$ by decreasing real part. The eigenvalues are marked by “x”.
In the asymptotic behavior of the condition numbers, a crucial role is played by the $n\times n$ matrix
$Q_{1}(t)=\sum\limits_{\begin{subarray}{c}\lambda_{i}\in\Lambda_{1}\\ j\in\{1,\ldots,d_{i}\}\\ m_{ij}=M_{1}\end{subarray}}\mathrm{e}^{\sqrt{-1}\ \omega_{i}t}v^{\left(i,j,1\right)}w^{\left(i,j,M_{1}\right)},$ (23)
where $\sqrt{-1}$ is the imaginary unit and
$M_{1}=\max\limits_{\lambda_{i}\in\Lambda_{1}}\max\limits_{j\in\{1,\ldots,d_{i}\}}m_{ij}$
is the maximum length of Jordan chains (maximum order of Jordan mini-blocks) corresponding to rightmost eigenvalues of $A$. In (23), the sum is over all Jordan chains, corresponding to rightmost eigenvalues, whose length $m_{ij}$ is equal to the maximum length $M_{1}$. In other words, over all Jordan mini-blocks, corresponding to rightmost eigenvalues, whose order is equal to the maximum order $M_{1}$. Observe that, for $u\in\mathbb{C}^{n}$, we have
$Q_{1}(t)u=\sum\limits_{\begin{subarray}{c}\lambda_{i}\in\Lambda_{1}\\ j\in\{1,\ldots,d_{i}\}\\ m_{ij}=M_{1}\end{subarray}}\mathrm{e}^{\sqrt{-1}\ \omega_{i}t}\alpha_{ijM_{1}}(u)v^{\left(i,j,1\right)}.$ (24)
Definition 4.
We say that $u\in\mathbb{C}^{n}$ satisfies the Rightmost Last Generalized Eigenvectors (RLGE) condition if some coefficient $\alpha_{ijM_{1}}(u)$ in (24) is non-zero, i.e. $u\in\mathbb{C}^{n}$ has some non-zero component along the last generalized eigenvectors $v^{(i,j,M_{1})}$ of the Jordan chains, corresponding to rightmost eigenvalues, with the maximum length $M_{1}$.
Note that if $u$ does not satisfy the RLGE condition, then $Q_{1}(t)u=0$ for all $t\in\mathbb{R}$.
Two important properties of the matrix $Q_{1}(t)$ are
$\inf\limits_{t\in\mathbb{R}}\|Q_{1}(t)\|>0,$ (25)
where $\|Q_{1}(t)\|$ is the matrix norm induced by the vector norm on $\mathbb{C}^{n}$, and
$\inf\limits_{t\in\mathbb{R}}\|Q_{1}(t)u\|>0$ (26)
for $u$ satisfying the RLGE condition.
3.2. Asymptotic equivalence
For the description of the asymptotic behavior of the condition numbers, we use the following notion of asymptotic equivalence. Let $a(t)$ and $b(t)$ be real functions of $t\in\mathbb{R}$. We say that $a(t)$ and $b(t)$ are asymptotically equivalent and write
$a\left(t\right)\sim b\left(t\right),\ t\rightarrow+\infty,$
if
$\lim\limits_{t\rightarrow+\infty}\frac{a(t)}{b(t)}=1.$
In other words, we can say that the asymptotic equivalence holds if
$\lim\limits_{t\rightarrow+\infty}\chi(t)=0,$
where $\chi(t)$ is the relative error of $a(t)$ with respect to $b(t)$, i.e. we have
$a(t)=b(t)\left(1+\chi(t)\right).$
3.3. The asymptic behavior of $K(t,y_{0},\widehat{z}_{0})$ and $K(t,y_{0})$
The next theorem describes the asymptotic behaviors of the condition numbers $K(t,y_{0},\widehat{z}_{0})$ and $K(t,y_{0})$.
Theorem 5.
For $y_{0}$ and $\widehat{z}_{0}$ satisfying the RLGE condition, we have
$K(t,y_{0},\widehat{z}_{0})\sim K_{\infty}(t,y_{0},\widehat{z}_{0}):=\frac{\left\|Q_{1}(t)\widehat{z}_{0}\right\|}{\left\|Q_{1}(t)\widehat{y}_{0}\right\|},\ t\rightarrow+\infty,$
and
$K(t,y_{0})\sim K_{\infty}(t,y_{0}):=\frac{\left\|Q_{1}(t)\right\|}{\left\|Q_{1}(t)\widehat{y}_{0}\right\|},\ t\rightarrow+\infty.$
We can observe the following.
  • 1.
    The RLGE condition is a generic condition for $y_{0}$ and $\widehat{z}_{0}$.
  • 2.
    By (26), we have
    $0<\inf\limits_{t\in\mathbb{R}}K_{\infty}(t,y_{0},\widehat{z}_{0})\text{\ \ and\ \ }\sup\limits_{t\in\mathbb{R}}K_{\infty}(t,y_{0},\widehat{z}_{0})<+\infty$
    and, by (25) and (26), we have
    $0<\inf\limits_{t\in\mathbb{R}}K_{\infty}(t,y_{0})\text{\ \ and\ \ }\sup\limits_{t\in\mathbb{R}}K_{\infty}(t,y_{0})<+\infty.$
  • 3.
    If $\Lambda_{1}$ consists of a (unique) real eigenvalue $\lambda_{1}$, then
    $Q_{1}(t)=Q_{1}=\sum\limits_{\begin{subarray}{c}j\in\{1,\ldots,d_{1}\}\\ m_{1j}=M_{1}\end{subarray}}v^{\left(1,j,1\right)}w^{\left(1,j,M_{1}\right)},$
    where
    $M_{1}=\max\limits_{j\in\{1,\ldots,d_{1}\}}m_{1j},$
    is a constant function of $t$ and then
    $K_{\infty}(t,y_{0},\widehat{z}_{0})=K_{\infty}(y_{0},\widehat{z}_{0})=\frac{\left\|Q_{1}\widehat{z}_{0}\right\|}{\left\|Q_{1}\widehat{y}_{0}\right\|}$
    and
    $K_{\infty}(t,y_{0})=K_{\infty}(y_{0})=\frac{\left\|Q_{1}\right\|}{\left\|Q_{1}\widehat{y}_{0}\right\|}$
    are constant functions of $t$.
In view of points 1-2-3 and (15), we can state a fact B) regarding the relative error of the perturbed solution, analogous to fact A) concerning the absolute error of the perturbed solution and presented at the beginning of the paper.
  • B)
    In a generic situation for the initial value and the perturbation of the initial value (namely they satisfy the RLGE condition), the relative error of the perturbed solution asymptotically (as $t\rightarrow+\infty$) neither diverges nor decays to zero. Instead, it converges to a non-zero constant, in case of a real eigenvalue as rightmost eigenvalue, and to an oscillating function bounded and uniformly away from zero, otherwise.
For any $y_{0}$ and $\widehat{z}_{0}$, not only for $y_{0}$ and $\widehat{z}_{0}$ satisfying the RGLE condition, there are functions
$t\rightarrow K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right)\text{\ \ and\ \ }t\rightarrow K_{\infty}\left(t,y_{0}\right),\ t\in\mathbb{R},$
such that
$K(t,y_{0},\widehat{z}_{0})\sim K_{\infty}(t,y_{0},\widehat{z}_{0})\text{\ \ and\ \ }K(t,y_{0})\sim K_{\infty}(t,y_{0}),\ t\rightarrow+\infty.$
They are called the asymptotic directional condition number and the asymptotic condition number, respectively, of the problem (2). Such functions for the general case are described in [19].
3.4. Is $K_{\infty}(t,y_{0})$ the worst $K_{\infty}(t,y_{0},\widehat{z}_{0})$?
By definition, the condition number $K\left(t,y_{0}\right)$ is the worst condition number $K\left(t,y_{0},\widehat{z}_{0}\right)$ as $\widehat{z}_{0}$ varies, at a specific time $t$. An interesting question is the following. Does this fact hold asymptotically as $t\rightarrow+\infty$? Specifically, is the asymptotic condition number $K_{\infty}\left(t,y_{0}\right)$ the worst asymptotic condition number $K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right)$ as $\widehat{z}_{0}$ varies? In other terms, do
[…]
commute? The answer is positive, since we have
$\max\limits_{\begin{subarray}{c}\widehat{z}_{0}\in\mathbb{C}^{n}\\ \|\widehat{z}_{0}\|=1\end{subarray}}\limsup\limits_{t\rightarrow+\infty}\frac{K\left(t,y_{0},\widehat{z}_{0}\right)}{K\left(t,y_{0}\right)}=\max\limits_{\begin{subarray}{c}\widehat{z}_{0}\in\mathbb{C}^{n}\\ \|\widehat{z}_{0}\|=1\end{subarray}}\limsup\limits_{t\rightarrow+\infty}\frac{K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right)}{K_{\infty}\left(t,y_{0}\right)}=1.$ (27)
In line with (27), we observe the following. If $y_{0}$ satisfies the RLGE condition and $\Lambda_{1}$ consists of a real eigenvalue, we have (see point 3 after Theorem 5)
$K_{\infty}(y_{0})=\max\limits_{\begin{subarray}{c}\widehat{z}_{0}\in\mathbb{C}^{n}\\ \|\widehat{z}_{0}\|=1\\ \widehat{z}_{0}\text{\ satisfying the RLGE condition}\end{subarray}}K_{\infty}\left(y_{0},\widehat{z}_{0}\right).$
4. The real case
Now, we consider a real ODE (1). In this case, we replace $\mathbb{C}$ by $\mathbb{R}$ and we have $A\in\mathbb{R}^{n\times n}$ and $y_{0},y(t),\widetilde{y}_{0},\widetilde{y}(t)\in\mathbb{R}^{n}$.
Our results concern the generic situation for the real matrix $A$ where the set $\Lambda_{1}$ of the rightmost eigenvalues of $A$ consists of either a real simple eigenvalue or a single pair of complex conjugate simple eigenvalues.
4.1. Notations
We introduce some notations.
  • 1)
    In the case where $\Lambda_{1}$ consists of a real simple eigenvalue $\lambda_{1}$, let $w^{(1)}$ (a row real vector) be a left eigenvector corresponding to $\lambda_{1}$ and let $\widehat{w}^{(1)}=\frac{w^{(1)}}{\|w^{(1)}\|}$ be its normalization. Here,
    $\|w^{\left(1\right)}\|=\max\limits_{\begin{subarray}{c}u\in\mathbb{R}^{n}\\ \|u\|=1\end{subarray}}|w^{\left(1\right)}u|$
    is the real induced norm of the real matrix $w^{(1)}$.
    We say that $\Lambda_{1}$ is real when $\Lambda_{1}$ consists of a real simple eigenvalue.
  • 2)
    In the case where $\Lambda_{1}$ consists of a single pair of complex conjugate simple eigenvalues $\lambda_{1}$ and $\overline{\lambda_{1}}$, with $\lambda_{1}$ having positive imaginary part $\omega_{1}$, let $w^{(1)}$ (a complex row vector) and $v^{(1)}$ (a complex column vector) be left and right, respectively, eigenvectors corresponding to $\lambda_{1}$ and let $\widehat{w}^{(1)}=\frac{w^{(1)}}{\|w^{(1)}\|}$ and $\widehat{v}^{(1)}=\frac{v^{(1)}}{\|v^{(1)}\|}$ be their normalizations. Here,
    $\|w^{\left(1\right)}\|=\max\limits_{\begin{subarray}{c}u\in\mathbb{C}^{n}\\ \|u\|=1\end{subarray}}|w^{\left(1\right)}u|$
    is the complex induced norm of the complex matrix $w^{(1)}$.
    Moreover, given the polar forms
    $\widehat{v}_{k}^{\left(1\right)}=\left|\widehat{v}_{k}^{\left(1\right)}\right|\mathrm{e}^{\sqrt{-1}\alpha_{1k}},\ k\in\{1,\ldots,n\},$
    $\widehat{w}_{l}^{\left(1\right)}=\left|\widehat{w}_{l}^{\left(1\right)}\right|\mathrm{e}^{\sqrt{-1}\beta_{1l}},\ l\in\{1,\ldots,n\},$
    of the components of the complex vectors $\widehat{v}^{\left(1\right)}$ and $\widehat{w}^{\left(1\right)}$ and the polar form
    $\widehat{w}^{\left(1\right)}u=\left|\widehat{w}^{\left(1\right)}u\right|\mathrm{e}^{\sqrt{-1}\gamma_{1}\left(u\right)}$
    of the complex scalar $\widehat{w}^{\left(1\right)}u$, where $u\in\mathbb{R}^{n}$, we introduce the vector
    $\widehat{\Theta}_{1}\left(t,u\right):=\left(\left|\widehat{v}_{k}^{\left(1\right)}\right|\cos\left(\omega_{1}t+\alpha_{1k}+\gamma_{1}\left(u\right)\right)\right)_{k=1,\ldots,n}\in\mathbb{R}^{n}$
    and the matrix
    $\widehat{\Theta}_{1}\left(t\right)=\left(\left|\widehat{v}_{k}^{\left(1\right)}\right|\left|\widehat{w}_{l}^{\left(1\right)}\right|\cos\left(\omega_{1}t+\alpha_{1k}+\beta_{1l}\right)\right)_{k,l=1,\ldots,n}\in\mathbb{R}^{n\times n}.$
    We say that $\Lambda_{1}$ is complex when $\Lambda_{1}$ consists of a single complex conjugate pair of simple eigenvalues.
4.2. The asymptic behavior of $K(t,y_{0},\widehat{z}_{0})$ and $K(t,y_{0})$
Next two results, derived by Theorem 5, consider the generic situation for $A$, described at the beginning of this Section 4, and a generic situation for $y_{0}$ and $\widehat{z}_{0}$. The generic situation for $y_{0}$ and $\widehat{z}_{0}$ is the RLGE condition, which becomes $w^{(1)}y_{0}\neq 0$ and $w^{(1)}\widehat{z}_{0}\neq 0$ in the generic situation for $A$ we are considering.
Theorem 6.
Assume that $A$ is a real matrix such that $\Lambda_{1}$ is real. For $y_{0}$ and $\widehat{z}_{0}$ such that $w^{(1)}y_{0}\neq 0$ and $w^{(1)}\widehat{z}_{0}\neq 0$, we have (recall point 3 after Theorem 5)
$K_{\infty}\left(y_{0},\widehat{z}_{0}\right)=\frac{\left|\widehat{w}^{(1)}\widehat{z}_{0}\right|}{\left|\widehat{w}^{(1)}\widehat{y}_{0}\right|}\text{\ \ and\ \ }K_{\infty}\left(y_{0}\right)=\frac{1}{\left|\widehat{w}^{(1)}\widehat{y}_{0}\right|}.$
Theorem 7.
Assume that $A$ is a real matrix such that $\Lambda_{1}$ is complex. For $y_{0}$ and $\widehat{z}_{0}$ such that $w^{(1)}y_{0}\neq 0$ and $w^{(1)}\widehat{z}_{0}\neq 0$, we have
$K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right)=\mathrm{OSF}(y_{0},\widehat{z}_{0})\cdot\mathrm{OT}(t,y_{0},\widehat{z}_{0})\text{\ \ and\ \ }K_{\infty}\left(t,y_{0}\right)=\mathrm{OSF}(y_{0})\cdot\mathrm{OT}(t,y_{0}),$
where
$\mathrm{OSF}(y_{0},\widehat{z}_{0}):=\frac{\left|\widehat{w}^{(1)}\widehat{z}_{0}\right|}{\left|\widehat{w}^{(1)}\widehat{y}_{0}\right|}\text{\ \ and\ \ }\mathrm{OSF}(y_{0}):=\frac{1}{\left|\widehat{w}^{(1)}\widehat{y}_{0}\right|}$ (28)
and
$\mathrm{OT}(t,y_{0},\widehat{z}_{0}):=\frac{\left\|\widehat{\Theta}_{1}\left(t,\widehat{z}_{0}\right)\right\|}{\left\|\widehat{\Theta}_{1}\left(t,\widehat{y}_{0}\right)\right\|}\text{\ \ and\ \ }\mathrm{OT}(t,y_{0}):=\frac{\left\|\widehat{\Theta}_{1}\left(t\right)\right\|}{\left\|\widehat{\Theta}_{1}\left(t,\widehat{y}_{0}\right)\right\|}.$ (29)
The constants $\mathrm{OSF}(y_{0},\widehat{z}_{0})$ and $\mathrm{OSF}(y_{0})$ are called oscillation scale factors and $\mathrm{OT}(t,y_{0},\widehat{z}_{0})$ and $\mathrm{OT}(t,y_{0})$, which are periodic functions of $t$ of period $\frac{\pi}{\omega_{1}}$, are called oscillating terms. Note that the oscillation scale factors depend on the moduli, and the oscillating terms on the angles, in the polar forms of the complex numbers $\widehat{w}^{(1)}\widehat{y}_{0}$ and $\widehat{w}^{(1)}\widehat{z}_{0}$.
4.3. The case $\Lambda_{1}$ complex and the Euclidean norm
In this subsection, in the case of $\Lambda_{1}$ complex, we analyze the oscillation scale factors and the oscillating terms when the vector norm is the Euclidean norm. It will be the norm used throughout the subsection. As in Section 3, the proofs for all results presented here can be found in [19].
The protagonists of this subsection are the numbers $V_{1}$ and $W_{1}$, which are now introduced.
4.3.1. $V_{1}$ and $W_{1}$
We introduce
$V_{1}:=\left|\left(\widehat{v}^{(1)}\right)^{T}\widehat{v}^{(1)}\right|\text{\ \ and\ \ }W_{1}:=\left|\widehat{w}^{(1)}\left(\widehat{w}^{(1)}\right)^{T}\right|,$
moduli of the complex numbers $\left(\widehat{v}^{(1)}\right)^{T}\widehat{v}^{(1)}$ and $\widehat{w}^{(1)}\left(\widehat{w}^{(1)}\right)^{T}$. Here T, unlike H, denote pure transposition without conjugation. We have $V_{1},W_{1}\in[0,1)$.
In Figure 13, we see, for $n=5,25,100$, the set of the pairs $(V_{1},W_{1})$ for 50 0005000050\ 00050 000 random matrices $A$ of order $n$ such that $\Lambda_{1}$ is complex. The elements of $A$ are sampled from the standard normal distribution. Observe that the distribution of the pairs in the square $[0,1)^{2}$ is not uniform, since the pairs tend to accumulate around the diagonal $V_{1}=W_{1}$.
Figure 13. Pairs $(V_{1},W_{1})$ for 50 0005000050\ 00050 000 random matrices $A$ of order $n=5,25,100$.
4.3.2. The set $\mathcal{W}_{1}$ and the ellipse $\mathcal{E}_{1}$
As we have seen above in Theorem 7, the asymptotic behavior of the condition numbers is determined by the complex numbers $\widehat{w}^{(1)}\widehat{y}_{0}$ and $\widehat{w}^{(1)}\widehat{z}_{0}$, with $\widehat{y}_{0}$ and $\widehat{z}_{0}$ unit real vectors. The moduli determine the oscillation scale factors and the angles in the polar forms determine the oscillating terms. Therefore, we now consider the set
$\mathcal{W}_{1}:=\{\widehat{w}^{(1)}u:u\in\mathbb{R}^{n}\text{\ and\ }\left\|u\right\|_{2}=1\}$
of the complex numbers $\widehat{w}^{(1)}u$, by varying the unit real vector $u$.
Next theorem describes the set $\mathcal{W}_{1}$.
Theorem 8.
Let
$R_{1}:=\left[\begin{array}[]{c}\mathrm{Re}(\widehat{w}^{(1)})\\ \mathrm{Im}(\widehat{w}^{(1)})\end{array}\right]\in\mathbb{R}^{2\times n},$ (30)
let $\sigma_{1}$ and $\mu_{1}$, with $\sigma_{1}\geq\mu_{1}\geq 0$, be the singular values of $R_{1}$ and let $\alpha^{(1)},\beta^{(1)}\in\mathbb{R}^{2}$ the corresponding left singular vectors of $R_{1}$. Remind that the left singular vectors of $R_{1}$ constitute an orthonormal basis of $\mathbb{R}^{2}$. We have
$\sigma_{1}=\sqrt{\frac{1+W_{1}}{2}}\text{\ \ and\ \ }\mu_{1}=\sqrt{\frac{1-W_{1}}{2}}.$
Consider, in the complex plane, the ellipse $\mathcal{E}_{1}$ centered at the origin with semi-axes the vectors $\sigma_{1}\alpha^{(1)}$ and $\mu_{1}\beta^{(1)}$. If $n=2$, the set $\mathcal{W}_{1}$ is the ellipse $\mathcal{E}_{1}$. If $n>2$, the set $\mathcal{W}_{1}$ is the union of the ellipse $\mathcal{E}_{1}$ and its interior.
Moreover, for $u\in\mathbb{R}^{n}$, we have
$\widehat{w}^{(1)}u=\sigma_{1}c_{1}(u)\alpha^{(1)}+\sqrt{-1}\mu_{1}d_{1}(u)\beta^{(1)},$ (31)
where $c_{1}(u)$ and $d_{1}(u)$ are the components of $u$ along the first right singular vector of $R_{1}$ (corresponding to the singular value $\sigma_{1}$) and the second right singular vector (corresponding to the singular value $\mu_{1}$), respectively. Remind that the right singular vectors of $R_{1}$ constitute an orthonormal basis of $\mathbb{R}^{n}$.
4.3.3. $\mathrm{OSF}(y_{0},\widehat{z}_{0})$ and $\mathrm{OT}(t,y_{0},\widehat{z}_{0})$
The oscillation scale factor $\mathrm{OSF}(y_{0},\widehat{z}_{0})$ in (28) is the ratio of the moduli of the complex numbers $\widehat{w}^{(1)}\widehat{z}_{0}$ and $\widehat{w}^{(1)}\widehat{y}_{0}$. Such complex numbers can be arbitrary non-zero points of the set $\mathcal{W}_{1}$. Thus, as $y_{0}$ and $\widehat{z}_{0}$ vary, $\mathrm{OSF}(y_{0},\widehat{z}_{0})$ can take any value in $\left[\frac{\mu_{1}}{\sigma_{1}},\frac{\sigma_{1}}{\mu_{1}}\right]$ for $n=2$, and any value in $(0,+\infty)$ for $n>2$. Moreover, by (31), we have
$\mathrm{OSF}(y_{0},\widehat{z}_{0})=\sqrt{\frac{(1+W_{1})c_{1}(\widehat{z}_{0})^{2}+(1-W_{1})d_{1}(\widehat{z}_{0})^{2}}{(1+W_{1})c_{1}(\widehat{y}_{0})^{2}+(1-W_{1})d_{1}(\widehat{y}_{0})^{2}}}.$
The next theorem specifies upper and lower bounds of the oscillating term $\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)$ in (29). Remind that $\gamma_{1}(u)$, $u\in\mathbb{R}^{n}$, is the angle in the polar form of $w^{(1)}u$. Moreover, remind that $V_{1}\in[0,1)$, as well as $W_{1}\in[0,1)$.
Theorem 9.
For $y_{0}$ and $\widehat{z}_{0}$ such that $w^{(1)}y_{0}\neq 0$ and $w^{(1)}\widehat{z}_{0}\neq 0$, we have
$\sqrt{\frac{1-V_{1}}{1+V_{1}}}\leq\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)\leq\sqrt{\frac{1+V_{1}}{1-V_{1}}},\ t\in\mathbb{R}.$
Moreover, if $\gamma_{1}(\widehat{z}_{0})-\gamma_{1}(\widehat{y}_{0})$ is an odd multiple of $\frac{\pi}{2}$, we have
$\min\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)=\sqrt{\frac{1-V_{1}}{1+V_{1}}}\text{\ \ and\ \ }\max\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)=\sqrt{\frac{1+V_{1}}{1-V_{1}}}.$ (32)
Regarding the previous theorem, we can make the following observations.
  • 1.
    If $V_{1}$ is not close to 1, then, for any $y_{0}$ and $\widehat{z}_{0}$, $\mathrm{OT}(t,y_{0},\widehat{z}_{0})$ does not assume large or small values as $t$ varies. Therefore, if $V_{1}$ is not close to 1, then, for any $y_{0}$ and $\widehat{z}_{0}$, the values of $\mathrm{OT}(t,y_{0},\widehat{z}_{0})$ as $t$ varies have the order of magnitude $1$. Consequently, the values of $K_{\infty}(t,y_{0},\widehat{z}_{0})$ as $t$ varies have the same order of magnitude as $\mathrm{OSF}(y_{0},\widehat{z}_{0})$.
  • 2.
    If $V_{1}$ is close to $1$, then, for any $y_{0}$ and for $\widehat{z}_{0}$ such that $\gamma_{1}(\widehat{z}_{0})-\gamma_{1}(\widehat{y}_{0})$ is an odd multiple of $\frac{\pi}{2}$, $\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)$ assumes both large and small values as $t$ varies.
  • 3.
    The condition that $\gamma_{1}(\widehat{z}_{0})-\gamma_{1}(\widehat{y}_{0})$ is an odd multiple of $\frac{\pi}{2}$ is equivalent to saying that the vectors $\widehat{w}^{(1)}\widehat{y}_{0}$ and $\widehat{w}^{(1)}\widehat{z}_{0}$ are orthogonal in the complex plane, i.e.
    $(1+W_{1})c_{1}(\widehat{y}_{0})c_{1}(\widehat{z}_{0})+(1-W_{1})d_{1}(\widehat{y}_{0})d_{1}(\widehat{z}_{0})=0$ (33)
    by (31).
  • 4.
    Two particular situations are worth considering.
    • If $y_{0}$ and $\widehat{z}_{0}$ are orthogonal to the first and second, respectively, right singular vectors of the matrix $R_{1}$ (i.e. $c_{1}(\widehat{y}_{0})=0$ and $d_{1}(\widehat{z}_{0})=0$ and thus (33) holds), then we have
[…]
      If $W_{1}$ is close to $1$, it is expected to have $\mathrm{OSF}(y_{0},\widehat{z}_{0})$ large. If $V_{1}$ is close to $1$, then $\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)$ assumes both large and small values as $t$ varies.
      Regarding the asymptotic condition number
      $K_{\infty}(t,y_{0},\widehat{z}_{0})=\mathrm{OSF}(y_{0},\widehat{z}_{0})\cdot\mathrm{OT}(t,y_{0},\widehat{z}_{0}),$
      it is expected that it assumes large values as $t$ varies, if $V_{1}$ is close to $1$ or $W_{1}$ is close to $1$.
    • If $y_{0}$ and $\widehat{z}_{0}$ are orthogonal to the second and first, respectively, right singular vectors of $R_{1}$ (i.e. $d_{1}(\widehat{y}_{0})=0$ and $c_{1}(\widehat{z}_{0})=0$ and thus (33) holds), then we have
[…]
      If $W_{1}$ is close to $1$, it is expected to have $\mathrm{OSF}(y_{0},\widehat{z}_{0})$ small. If $V_{1}$ is close to $1$, then $\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)$ assumes both large and small values as $t$ varies.
      Regarding the asymptotic condition number
      $K_{\infty}(t,y_{0},\widehat{z}_{0})=\mathrm{OSF}(y_{0},\widehat{z}_{0})\cdot\mathrm{OT}(t,y_{0},\widehat{z}_{0}),$
      it is expected that it assumes large values as $t$ varies, if $\sqrt{\frac{1-W_{1}}{1-V_{1}}}$ is large.
4.3.4. $\mathrm{OSF}(y_{0})$ and $\mathrm{OT}(t,y_{0})$
The oscillation scale factor $\mathrm{OSF}(y_{0})$ in (28) is the reciprocal of the modulus of the complex number $\widehat{w}^{(1)}\widehat{y}_{0}$. Such complex number can be an arbitrary non-zero point in the set $\mathcal{W}_{1}$. Thus, as $y_{0}$ varies, $\mathrm{OSF}(y_{0})$ can take any value in $\left[\frac{1}{\sigma_{1}},\frac{1}{\mu_{1}}\right]$ for $n=2$, and any value in $\left[\frac{1}{\sigma_{1}},+\infty\right)$ for $n>2$. Moreover, by (31), we have
$\mathrm{OSF}(y_{0})=\sqrt{\frac{2}{(1+W_{1})c_{1}(\widehat{y}_{0})^{2}+(1-W_{1})d_{1}(\widehat{y}_{0})^{2}}}.$
The next theorem specifies upper and lower bounds of the oscillating term $\mathrm{OT}\left(t,y_{0}\right)$ in (29). Let
$a_{\min}(V_{1},W_{1}):=\left\{\begin{array}[]{l}\sqrt{\frac{(1+W_{1})(1-V_{1})}{2(1+V_{1})}}\text{\ if\ }V_{1}\leq W_{1}\\ \\ \sqrt{\frac{1-W_{1}}{2}}\text{\ if\ }V_{1}\geq W_{1},\end{array}\right.$ (34)
let
$a_{\max}(V_{1},W_{1}):=\sqrt{\frac{(1+W_{1})(1+V_{1})}{2(1-V_{1})}}$ (35)
and let $\theta_{1}$ be the angle in the complex plane between the real axis and the major semi-axis $\sigma_{1}\alpha^{(1)}$ of the ellipse $\mathcal{E}_{1}$ introduced in Theorem 8.
Theorem 10.
For $y_{0}$ such that $w^{(1)}y_{0}\neq 0$, we have
$a_{\min}(V_{1},W_{1})\leq\mathrm{OT}(t,y_{0})\leq a_{\max}(V_{1},W_{1}),\ t\in\mathbb{R}.$
Moreover, if $\gamma_{1}(\widehat{y}_{0})-\theta_{1}$ is an odd multiple of $\frac{\pi}{2}$, then
$\min\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0}\right)=a_{\min}(V_{1},W_{1})\text{\ \ and\ \ }\max\limits_{t\in\mathbb{R}}\mathrm{OT}\left(t,y_{0}\right)=a_{\max}(V_{1},W_{1}).$ (36)
Regarding the previous theorem, we can make the following observations.
  • 1.
    If $V_{1}$ is not close to $1$, then, for any $y_{0}$, $\mathrm{OT}(t,y_{0})$ does not assume large values as $t$ varies. If $V_{1}$ is not close to $1$ or $W_{1}$ is not close to $1$, then, for any $y_{0}$, $\mathrm{OT}(t,y_{0})$ does not assume small values as $t$ varies. Therefore, if $V_{1}$ is not close to 1, then, for any $y_{0}$, the values of $\mathrm{OT}(t,y_{0})$ as $t$ varies have the order of magnitude $1$. Consequently, the values of $K_{\infty}(t,y_{0})$ as $t$ varies have the same order of magnitude as $\mathrm{OSF}(y_{0})$.
  • 2.
    The condition $\gamma_{1}(\widehat{y}_{0})-\theta_{1}$ is an odd multiple of $\frac{\pi}{2}$ is equivalent to saying that the vector $\widehat{w}^{(1)}\widehat{y}_{0}$ is orthogonal in the complex plane to the major semi-axis $\sigma_{1}\alpha^{(1)}$ of the ellipse $\mathcal{E}_{1}$, i.e. $c_{1}(\widehat{y}_{0})=0$ by (31). This indicates that $y_{0}$ is orthogonal to the first right singular vector of the matrix $R_{1}$. In this case, we have
[…]
    If $W_{1}$ is close to $1$, then $\mathrm{OSF}(y_{0})$ is large. If $V_{1}$ is close to $1$, then $\mathrm{OT}\left(t,y_{0}\right)$ assumes large values as $t$ varies. If $V_{1}$ is close to $1$ and $W_{1}$ is close $1$, then $\mathrm{OT}\left(t,y_{0}\right)$ assumes small values as $t$ varies.
    Regarding the asymptotic condition number
    $K_{\infty}(t,y_{0})=\mathrm{OSF}(y_{0})\cdot\mathrm{OT}(t,y_{0}),$
    it assumes large values as $t$ varies, if $V_{1}$ is close to $1$ or $W_{1}$ is close to $1$.
4.3.5. Is $V_{1}$ close to $1$?
We have seen in points 1, following Theorems 9 and 10, that the asymptotic condition numbers have the same order of magnitude as the oscillation scale factors when $V_{1}$ is not close to 1, since, in this case, the oscillating terms have the order of magnitude $1$. Therefore, it is of interest to understand how frequently $V_{1}$ is close to 1.
The following table shows, for the 50 0005000050\ 00050 000 random matrices of Figure 13, the percentages of cases with $V_{1}$ greater than $0.9$, $0.99$, $0.999$ and $0.9999$.
$V_{1}>0.9$ $V_{1}>0.99$ $V_{1}>0.999$ $V_{1}>0.9999$
$n=5$ 6% 0.6% 0.05% 0.002%
$n=25$ 4% 0.4% 0.03% 0.002%
$n=100$ 3% 0.3 % 0.03% 0.004%
We have strong evidence that $V_{1}$ is rarely close to 1.
Regarding the order of magnitude of the oscillating terms when $V_{1}$ is close to $1$, note that
$\frac{1}{k}\leq\sqrt{\frac{1-V_{1}}{1+V_{1}}}\leq\mathrm{OT}(t,y_{0},\widehat{z}_{0})\leq\sqrt{\frac{1+V_{1}}{1-V_{1}}}\leq k$
and
$\frac{1}{k}\leq a_{\mathrm{min}}(V_{1},W_{1})\leq\mathrm{OT}(t,y_{0})\leq a_{\mathrm{max}}(V_{1},W_{1})\leq k$
where
$k=\sqrt{\frac{2}{1-V_{1}}}.$
The values of $k$ for $V_{1}=0.9,0.99,0.999,0.9999$ in the table (37) above are
$V_{1}=0.9$ $V_{1}=0.99$ $V_{1}=0.999$ $V_{1}=0.9999$
$k$ $4.4721$ $14.1421$ $44.7214$ $141.4214$
5. The onset of the asymptotic behavior
As mentioned in the introduction, it is important to understand when the condition numbers $K(t,y_{0},\widehat{z}_{0})$ and $K(t,y_{0})$ begin to get close to their asymptotic versions $K_{\infty}(t,y_{0},\widehat{z}_{0})$ and $K_{\infty}(t,y_{0})$.
In this section, we investigate such a question by considering, as in the previous section, a real ODE (1). We assume that the real matrix $A$ satisfies a condition a little bit stronger than the generic condition of the previous section, where we assumed that $\Lambda_{1}$ consists of either a real simple eigenvalue or a single pair of complex conjugate simple eigenvalues. Here, we assume that, for any $j\in\{1,\ldots,q\}$, $\Lambda_{j}$ (see Figure 12) consists of either a real simple eigenvalue or a single pair of complex conjugate simple eigenvalues. This latter condition, nonetheless, remains a generic condition for the real matrix $A$.
Once again, the proofs of the results presented in this section can be found in [19].
5.1. Notations
We introduce for $\Lambda_{j}$, $j\in\{1,\ldots,q\}$, notations similar to the notations 1) and 2) introduced for $\Lambda_{1}$ in Subsection 4.1.
  • 1)
    In the case where $\Lambda_{j}$ consists of a real simple eigenvalue $\lambda_{j}$, let $w^{(j)}$ and $v^{(j)}$ be left and right, respectively, eigenvectors corresponding to $\lambda_{j}$. The normalized vector $\widehat{w}^{(j)}$ is defined similarly to $\widehat{w}^{(1)}$.
    We say that $\Lambda_{j}$ is real when $\Lambda_{j}$ consists of a real simple eigenvalue.
  • 2)
    In the case where $\Lambda_{j}$ consists of a single pair of complex conjugate simple eigenvalues $\lambda_{j}$ and $\overline{\lambda_{j}}$, with $\lambda_{j}$ having a positive imaginary part $\omega_{j}$, let $w^{(j)}$ and $v^{(j)}$ be left and right, respectively, eigenvectors corresponding to $\lambda_{j}$.
    Moreover, we introduce the normalized vectors $\widehat{w}^{(j)}$ and $\widehat{v}^{(j)}$, the vector $\widehat{\Theta}_{j}(t,u)\in\mathbb{R}^{n}$, with $u\in\mathbb{R}^{n}$, the matrix $\widehat{\Theta}_{j}(t)\in\mathbb{R}^{n\times n}$ and the scalars $V_{j},W_{j}\in[0,1)$, defined similarly to $\widehat{w}^{(1)}$, $\widehat{v}^{(1)}$, $\widehat{\Theta}_{1}(t,u)$, $\widehat{\Theta}_{1}(t)$, $V_{1}$, and $W_{1}$.
    We say that $\Lambda_{j}$ is complex when $\Lambda_{j}$ consists of a single complex conjugate pair of simple eigenvalue.
We also introduce the values
$f_{j}:=\|w^{(j)}\|\cdot\|v^{(j)}\|\in[1,+\infty),\ j\in\{1,\ldots,q\}.$ (38)
5.2. Approximation with a given precision
The closeness of the condition numbers to their asympototic version is measured by using the following notion of approximation with a given precision. For $a,b\in\mathbb{R}$ and $\epsilon\geq 0$, we write
$a\approx b\text{\ with precision\ }\epsilon$
if
$a=b(1+\chi)\text{\ with\ }|\chi|\leq\epsilon,$
i.e. the relative error $\chi$ of $a$ with respect to $b$ has modulus not larger than $\epsilon$.
5.3. Closeness to the asymptotic behavior
The next theorem describes how the condition numbers get closer to their asymptotic versions, as $t$ increases.
Theorem 11.
Assume that $A$ is a real matrix such that, for any $j\in\{1,\ldots,q\}$, $\Lambda_{j}$ is real or $\Lambda_{j}$ is complex. For $y_{0}$ and $\widehat{z}_{0}$ such that $w^{(1)}y_{0}\neq 0$ and $w^{(1)}\widehat{z}_{0}\neq 0$, we have
$K\left(t,y_{0},\widehat{z}_{0}\right)\approx K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right)\text{\ with precision\ }\frac{\epsilon(t,\widehat{z}_{0})+\epsilon(t,\widehat{y}_{0})}{1-\epsilon(t,\widehat{y}_{0})}$
and
$K\left(t,y_{0}\right)\approx K_{\infty}\left(t,y_{0}\right)\text{\ with precision\ }\frac{\epsilon(t)+\epsilon(t,\widehat{y}_{0})}{1-\epsilon(t,\widehat{y}_{0})}$
whenever $\epsilon(t,\widehat{y}_{0})<1$, where
$\epsilon(t,u):=\sum\limits_{j=2}^{q}\mathrm{e}^{\left(r_{j}-r_{1}\right)t}\frac{f_{j}}{f_{1}}\cdot\frac{\left|\widehat{w}^{(j)}u\right|}{\left|\widehat{w}^{(1)}u\right|}G_{j}\left(t,u\right),\ u=\widehat{y}_{0},\widehat{z}_{0},$
and
$\epsilon(t):=\sum\limits_{j=2}^{q}\mathrm{e}^{\left(r_{j}-r_{1}\right)t}\frac{f_{j}}{f_{1}}G_{j}(t).$
Here, $G_{j}(t,u)$ and $G_{j}(t)$, $j\in\{2,\ldots,q\}$, are defined as follows.
  • 1)
    If both $\Lambda_{j}$ and $\Lambda_{1}$ are real, then
    $G_{j}(t,u)=1\text{\ \ and\ \ }G_{j}(t)=1.$
  • 2)
    If $\Lambda_{j}$ is complex and $\Lambda_{1}$ is real, then
    $G_{j}(t,u)=2\left\|\widehat{\Theta}_{j}(t,u)\right\|\text{\ \ and\ \ }G_{j}(t)=2\left\|\widehat{\Theta}_{j}(t)\right\|.$
  • 3)
    If $\Lambda_{j}$ is real and $\Lambda_{1}$ is complex, then
    $G_{j}(t,u)=\frac{1}{2\left\|\widehat{\Theta}_{1}(t,u)\right\|}\text{\ \ and\ \ }G_{j}(t)=\frac{1}{2\left\|\widehat{\Theta}_{1}(t)\right\|}.$
  • 4)
    If both $\Lambda_{j}$ and $\Lambda_{1}$ are complex, then
    $G_{j}(t,u)=\frac{\left\|\widehat{\Theta}_{j}(t,u)\right\|}{\left\|\widehat{\Theta}_{1}(t,u)\right\|}\text{\ \ and\ \ }G_{j}(t)=\frac{\left\|\widehat{\Theta}_{j}(t)\right\|}{\left\|\widehat{\Theta}_{1}(t)\right\|}.$
Regarding points 2), 3), and 4) in the previous theorem, the following can be stated.
  • 2bis)
    If $\Lambda_{j}$ is complex and $\Lambda_{1}$ is real and the vector norm is a $p$-norm, then
    $G_{j}(t,u)\leq 2\text{\ \ and\ \ }G_{j}(t)\leq 2.$
  • 3bis)
    If $\Lambda_{j}$ is real and $\Lambda_{1}$ is complex and the vector norm is the Euclidean norm, then
    $G_{j}(t,u)\leq\sqrt{\frac{1}{2(1-V_{1})}}\text{\ \ and\ \ }G_{j}(t)\leq\sqrt{\frac{1}{a(V_{1},W_{1})}},$
    where
    $a(V_{1},W_{1}):=\left\{\begin{array}[]{l}(1-V_{1})(1+W_{1})\text{\ if\ }V_{1}\leq W_{1}\\ \\ (1+V_{1})(1-W_{1})\text{\ if\ }V_{1}\geq W_{1}.\end{array}\right.$
  • 4bis)
    If both $\Lambda_{j}$ and $\Lambda_{1}$ are complex and the vector norm is the Euclidean norm, then
    $G_{j}(t,u)\leq\sqrt{\frac{1+V_{j}}{1-V_{1}}}\text{\ \ and\ \ }G_{j}(t)\leq\sqrt{\frac{(1+W_{j})(1+V_{j})}{a(V_{1},W_{1})}}.$
Observe that $\epsilon(t,u)\rightarrow 0$ and $\epsilon(t)\rightarrow 0$, as $t\rightarrow+\infty$, since $r_{j}-r_{1}<0$, $j\in\{2,\ldots,q\}$, in the exponential functions. Thus, we retrieve
$K(t,y_{0},\widehat{z}_{0})\sim K_{\infty}(t,y_{0},\widehat{z}_{0})\text{\ \ and\ \ }K(t,y_{0})\sim K_{\infty}(t,y_{0}),\ t\rightarrow+\infty.$
In conclusion, the onset of the asymptotic behavior, i.e. when the condition numbers begin to get close to their asymptotic versions, is determined by how rapidly $\epsilon(t,u)$, $u=\widehat{y}_{0},\widehat{z}_{0}$, and $\epsilon(t)$ reach small values. The factors influencing this can be found in the expressions for $\epsilon(t,u)$ and $\epsilon(t)$ in the previous Theorem 11. They are, for $j\in\{2,\ldots,q\}$:
  • the decreasing exponentials $\mathrm{e}^{(r_{j}-r_{1})t}$;
  • the ratios $\frac{f_{j}}{f_{1}}$;
  • the ratios $\frac{\left|\widehat{w}^{(j)}u\right|}{\left|\widehat{w}^{(1)}u\right|}$;
  • the numbers $G_{j}(t,u)$ and $G_{j}(t)$.
An exploration of such factors influencing how rapidly $K(t,y_{0})$ get close to $K_{\infty}(t,y_{0})$ is presented in the next section.
6. Some considerations
This section discusses two important issues concerning the asymptotic condition number $K_{\infty}(t,y_{0})$. It is the most important between the two asymptotic condition numbers since, in general, we have no information about the direction of perturbation $\widehat{z}_{0}$.
As in the previous section, we assume the generic situation of a real matrix $A$ in (1) such that, for any $j\in\{1,\ldots,q\}$, $\Lambda_{j}$ is real or $\Lambda_{j}$ is complex, and of an initial value $y_{0}$ such that $w^{(1)}y_{0}\neq 0$.
It is of interest to know:
  • AWC)
    for which initial values $y_{0}$ the problem (2) is asymptotically well-conditioned, i.e. $K_{\infty}(t,y_{0})$ does not assume large values as $t$ varies;
  • ONS)
    the onset of the asymptotic behavior, i.e. when $K(t,y_{0})$ begins to be close to $K_{\infty}(t,y_{0})$; as we have seen, this happens as soon as $\epsilon(t,\widehat{y}_{0})$ and $\epsilon(t)$ are small.
6.1. The case $\Lambda_{1}$ real
Suppose that $\Lambda_{1}$ is real. We have
$K_{\infty}(y_{0})=\frac{1}{\left|\widehat{w}^{(1)}\widehat{y}_{0}\right|}$
Thus, regarding the issue AWC) for $\Lambda_{1}$ real, we can conclude that:
  • AWCr)
    the problem (2) is asymptotically well-conditioned if and only if $\left|\widehat{w}^{(1)}\widehat{y}_{0}\right|$ is not small.
Concerning the onset of the asymptotic behavior, by Theorem 11 and the point 2bis) immediately after the theorem, we have, when the vector norm is a $p$-norm,
$\max\{\epsilon(t,\widehat{y}_{0}),\epsilon(t)\}\leq 2\sum\limits_{j=2}^{q}\mathrm{e}^{\left(r_{j}-r_{1}\right)t}\frac{f_{j}}{f_{1}}\max\left\{1,\frac{\left|\widehat{w}^{(j)}\widehat{y}_{0}\right|}{\left|\widehat{w}^{(1)}\widehat{y}_{0}\right|}\right\}$
and then we have $\max\{\epsilon(t,\widehat{y}_{0}),\epsilon(t)\}\leq\epsilon$, where $\epsilon>0$, if
$\frac{t}{\widehat{t}}\geq\max\limits_{j\in\{2,\ldots,q\}}\frac{1}{\left(r_{1}-r_{j}\right)\widehat{t}}\left(\log\frac{2(q-1)}{\epsilon}+\log\frac{f_{j}}{f_{1}}+\max\left\{0,\log\frac{\left|\widehat{w}^{(j)}\widehat{y}_{0}\right|}{\left|\widehat{w}^{(1)}\widehat{y}_{0}\right|}\right\}\right),$
where $\widehat{t}$ is the time used as time unit. If $r_{1}\neq 0$, one possible choice for the time unit is the characteristic time $\frac{1}{|r_{1}|}$. Thus, regarding the issue ONS) for $\Lambda_{1}$ real and a $p$-norm as vector norm, by Theorem 11 we can conclude that:
  • ONSr)
    for $\overline{\epsilon}>0$, we have $K(t,y_{0})\approx K_{\infty}(y_{0})$ with precision $\overline{\epsilon}$ if
    $\frac{t}{\widehat{t}}\geq\max\limits_{j\in\{2,\ldots,q\}}\frac{1}{\left(r_{1}-r_{j}\right)\widehat{t}}\left(\log 2+\log\frac{2+\overline{\epsilon}}{\overline{\epsilon}}+\log(q-1)\right.$
    $\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\left.+\log\frac{f_{j}}{f_{1}}+\max\left\{0,\log\frac{\left|\widehat{w}^{(j)}\widehat{y}_{0}\right|}{\left|\widehat{w}^{(1)}\widehat{y}_{0}\right|}\right\}\right).$
    Therefore, the smaller the right-hand side in (39), the earlier the onset of the asymptotic behavior.
6.2. The case $\Lambda_{1}$ complex
Suppose that $\Lambda_{1}$ is complex and the vector norm is the Euclidean norm.
We have
$K_{\infty}(t,y_{0})=\mathrm{OSF}(y_{0})\cdot\mathrm{OT}(t,y_{0}),$
where the oscillation scale factor $\mathrm{OSF}(y_{0})$ is given by
$\mathrm{OSF}(y_{0})=\frac{1}{\left|\widehat{w}^{(1)}\widehat{y}_{0}\right|}.$
If $V_{1}$ is not close to $1$, then the values of $K_{\infty}(t,y_{0})$ as $t$ varies have the same order of magnitude as $\mathrm{OSF}(y_{0})$, since the values of $\mathrm{OT}(t,y_{0})$ as $t$ varies have the order of magnitude $1$ (see point 1 after Theorem 10). Thus, regarding the issue AWC) for $\Lambda_{1}$ complex and the Euclidean norm as vector norm, we can conclude that:
  • AWCc)
    if $V_{1}$ is not close to $1$, then the problem (2) is asymptotically well-conditioned if and only if $\left|\widehat{w}^{(1)}\widehat{y}_{0}\right|$ is not small.
If $V_{1}$ is close to $1$, then $\left|\widehat{w}^{(1)}\widehat{y}_{0}\right|$ does not determine the asymptotic well-conditioning of the problem (2), unlike when $V_{1}$ is not close to $1$, since it no longer guaranteed that the values of $\mathrm{OT}(t,y_{0})$ as $t$ varies maintain the order of magnitude of $1$. Indeed, if $V_{1}$ is close to $1$ and $y_{0}$ is orthogonal to the first right singular vector of the matrix $R_{1}$ in (30), then $\mathrm{OT}(t,y_{0})$ assume large values as $t$ varies (see point 2 after Theorem 10). On the other hand, if $W_{1}$ is not close to $1$, then, for such $y_{0}$,
$\mathrm{OSF}(y_{0})=\frac{1}{\left|\widehat{w}^{(1)}\widehat{y}_{0}\right|}=\sqrt{\frac{2}{1-W_{1}}}\cdot\frac{1}{|d(\widehat{y}_{0})|}$
is not large if $|d(\widehat{y}_{0})|$ is not small, e.g. if $y_{0}$ is the second right singular vector of the matrix $R_{1}$. If this occurs, we have a situation where $K_{\infty}(t,y_{0})$ assumes large values as $t$ varies, i.e. the problem (2) is not asymptotically well-conditioned, although $\left|\widehat{w}^{(1)}\widehat{y}_{0}\right|$ is not small.
However, by recalling Subsection 4.3.5, observe that $V_{1}$ is rarely close to $1$.
Concerning the onset of the asymptotic behavior, by Theorem 11 and points 3bis) and 4bis) immediately after the theorem, we have
$\max\{\epsilon(t,\widehat{y}_{0}),\epsilon(t)\}\leq\frac{2}{\sqrt{1-V_{1}}}\sum\limits_{j=2}^{q}\mathrm{e}^{\left(r_{j}-r_{1}\right)t}\frac{f_{j}}{f_{1}}\max\left\{1,\frac{\left|\widehat{w}^{(j)}\widehat{y}_{0}\right|}{\left|\widehat{w}^{(1)}\widehat{y}_{0}\right|}\right\}.$
(The term $\frac{2}{\sqrt{1-V_{1}}}$ is obtained by comparing all the upper bounds for $G_{j}(t,u)$ and $G_{j}(t)$, $j\in\{2,\ldots,q\}$, provided in 3bis) and 4bis) and by noting that $a(V_{1},W_{1})\geq 1-V_{1}$). Thus, regarding the issue ONS) for $\Lambda_{1}$ complex and the Euclidean norm as vector norm, by Theorem 11 we can conclude that:
  • ONSc)
    for $\overline{\epsilon}>0$, we have $K(t,y_{0})\approx K_{\infty}(t,y_{0})$ with precision $\overline{\epsilon}$ if
    $\frac{t}{\widehat{t}}\geq\max\limits_{j\in\{2,\ldots,q\}}\frac{1}{\left(r_{1}-r_{j}\right)\widehat{t}}\left(\log 2+\log\frac{2+\overline{\epsilon}}{\overline{\epsilon}}+\log(q-1)\right.$
    $\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\left.+\frac{1}{2}\log\frac{1}{1-V_{1}}+\log\frac{f_{j}}{f_{1}}+\max\left\{0,\log\frac{\left|\widehat{w}^{(j)}\widehat{y}_{0}\right|}{\left|\widehat{w}^{(1)}\widehat{y}_{0}\right|}\right\}\right).$
    Therefore, the smaller the right-hand side in (40), the earlier the onset of the asymptotic behavior.
By comparing (39) and (40), we observe the presence of the term $\frac{1}{2}\log\frac{1}{1-V_{1}}$ in (40). If $V_{1}$ is close to $1$, we could observe a delayed onset of the asymptotic behavior, compared to the case where $V_{1}$ is not close to $1$ or to the case where $\Lambda_{1}$ is real.
6.3. Non-normal matrices
Suppose that the vector norm is the Euclidean norm. For a normal matrix $A$, we have $f_{j}=1,\ j\in\{1,\ldots,q\}$, and, in the case of $\Lambda_{1}$ complex, $V_{1}=0$. On the other hand, a non-normal matrix $A$ can exhibit $f_{j}$ values that are arbitrarily large, as well as $V_{1}$ values that are arbitrarily close to $1$.
It is important to understand the impact of the non-normality of the matrix $A$ on the two aforementioned issues AWC) and ONS).
  • AWCnn)
    Regarding the asymptotic well-conditioning of the problem (2), we can say that the non-normality of $A$ has an impact only if $\Lambda_{1}$ is complex. In fact, if $\Lambda_{1}$ is real, or $\Lambda_{1}$ is complex and $V_{1}$ is not close to $1$, then the problem (2) is asymptotically well-conditioned if and only if $|\widehat{w}^{(1)}\widehat{y}_{0}|$ is not small, where the quantity $|\widehat{w}^{(1)}\widehat{y}_{0}|$ is unrelated to the non-normality. On the other hand, if $\Lambda_{1}$ is complex and $V_{1}$ is close to $1$, a case possible only for $A$ non-normal, the quantity $|\widehat{w}^{(1)}\widehat{y}_{0}|$ does not determine the asymptotic well-conditioning of the problem (2) (recall the previous Subsection 6.2).
  • ONSnn)
    Regarding the onset of the asymptotic behavior, if some ratio $\frac{f_{j}}{f_{1}}$, $j\in\{2,\ldots,q\}$, is large, or $\Lambda_{1}$ is complex and $V_{1}$ is close to 1, cases possible only for $A$ non-normal, the onset of the asymptotic behavior could be delayed (recall (39) and (40)).
It’s worth noting that in ONSnn) the focus is on the ratios $\frac{f_{j}}{f_{1}}$, $j\in\{2,\ldots,q\}$, which can be much smaller than the large values of $f_{j}$ characterizing an high non-normality of the matrix $A$. More details are given in the next example.
Example 12.
In Figure 14, for the Euclidean norm as vector norm, we see the maxima
$M=\max\limits_{j\in\{2,\ldots,n\}}\log f_{j}\text{\ \ and\ \ }\widehat{M}=\max\limits_{j\in\{2,\ldots,n\}}\log\frac{f_{j}}{f_{1}},$ (41)
for 10 0001000010\ 00010 000 matrices $A$ of order $n=500$, whose elements are sampled from the standard normal distribution. It appears that the values $\widehat{M}$ are smaller than the values $M$. The following table provides some statistical details:
maximum value minimum value median
$M$ $10.06$ $3.41$ $4.35$
$\widehat{M}$ $8.00$ $0.00$ $2.95$
Figure 14. Maxima $M$ and $\widehat{M}$ in (41) for 10 0001000010\ 00010 000 random instances of $A$ of order $500$.
Moreover, the example in Subsection 7.5 also highlights that the ratios $\frac{f_{j}}{f_{1}}$ can be significantly smaller than the values of $f_{j}$.
6.4. Normal matrices
Once again, assume that the vector norm is the Euclidean norm and consider $A$ normal.
In the case where $\Lambda_{1}$ is complex, we have $V_{1}=W_{1}=0$. Consequently, the oscillating term $\mathrm{OT}(t,y_{0})$ has a constant in time value of $\sqrt{\frac{1}{2}}$ and $K_{\infty}(t,y_{0})$ has a constant in time value of $\frac{1}{|\widehat{w}^{(1)}\widehat{y}_{0}|}\sqrt{\frac{1}{2}}$. Note that, in the case where $\Lambda_{1}$ is real, the constant in time $K_{\infty}(y_{0})$ has a value of $\frac{1}{|\widehat{w}^{(1)}\widehat{y}_{0}|}$.
Moreover, since $\frac{f_{j}}{f_{1}}=1$ for $j\in\{2,\ldots,q\}$, and $V_{1}=0$ in the case where $\Lambda_{1}$ is complex, an earlier onset of the asymptotic behavior for $K(t,y_{0})$ is expected compared to non-normal matrices (see (39) and (40)).
Finally, the constant in time asymptotic condition number is not only the limit as $t\rightarrow+\infty$ of $K(t,y_{0})$, but it is also the supremum of $K(t,y_{0})$ for $t\geq 0$ (see [18]).
7. Examples
To illustrate the analysis given in this paper, we consider seven examples of a real ODE (1), where the first two are the models of Section 1 revisited.
7.1. Gross Domestic Product and National Debt
The first example is the ODE (3) of the GDP-ND model of Section 1. Assume, as in the instance (4), that there are two real eigenvalues. Moreover, since we set $Q(0)=1$, there is uncertainty only on $B(0)$ and then the direction of perturbation is $\widehat{z}_{0}=(0,\pm 1)$ (in any vector $p$-norm). In the Euclidean norm, we have
$K_{\infty}(y_{0},\widehat{z}_{0})=\frac{\left|\widehat{w}^{(1)}\widehat{z}_{0}\right|}{\left|\widehat{w}^{(1)}\widehat{y}_{0}\right|}=\sqrt{1+B(0)^{2}}\cdot\frac{1}{\left|\frac{w^{(1)}_{1}}{w^{(1)}_{2}}+B(0)\right|}.$
In the instance (4), we have $\frac{w^{(1)}_{1}}{w^{(1)}_{2}}=-1$ and so
$K_{\infty}(y_{0},\widehat{z}_{0})=\sqrt{1+B(0)^{2}}\cdot\frac{1}{|-1+B(0)|}.$ (42)
For the initial ND $B(0)=0.61$ considered in Section 1, we have $K_{\infty}(y_{0},\widehat{z}_{0})=3.0035$.
In Figure 15, which extends Figure 5 to the interval $[0,150]$ (with a characteristic time $20$ in this context), we can observe how the relative error $\delta(t)$ tends to asymptotically approach the value $0.025641$. We have, remind (15) and $\varepsilon=\delta(0)$,
$\delta(t)=K(t,y_{0},\widehat{z}_{0})\cdot\delta(0)\text{\ \ and\ \ }\lim\limits_{t\rightarrow+\infty}\delta(t)=K_{\infty}(y_{0},\widehat{z}_{0})\cdot\delta(0)=3.0035\cdot\delta(0).$
The expression for $K_{\infty}(y_{0},\widehat{z}_{0})$ in (42) has a singularity at $B(0)=1$. In this case, where the initial value $(Q(0),B(0))=(1,1)$ has zero component along the rightmost eigenvector, the asymptotic condition number $K_{\infty}(t,y_{0},\widehat{z}_{0})$ grows exponentially in time (see [19]).
For $(Q(0),B(0))=(1,1)$ and $(\widetilde{Q}(0),\widetilde{B}(0))=(1,0.99)$, Figures 16, 17, and 18 depict the same information as Figures 3, 4, and 5 do for $(Q(0),B(0))=(1,0.61)$ and $(\widetilde{Q}(0),\widetilde{B}(0))=(1,0.60)$. The differences between the two cases are quite evident. After $50\ \mathrm{yr}$, the relative error is about $12$ times the initial relative error.
Figure 15. Relative error $\delta(t)$ for the GDP-ND model.
Figure 16. $B(0)=1$ and $\widetilde{B}(0)=0.99$. Left: points $(Q(0),B(0))$ and $(\widetilde{Q}(0),\widetilde{B}(0))$. Right: points $(Q(50),B(50))$ and $(\widetilde{Q}(50),\widetilde{B}(50))$.
Figure 17. $B(0)=1$ and $\widetilde{B}(0)=0.99$. GDP $Q(t)$, perturbed GDP $\widetilde{Q}(t)$, ND $B(t)$ and perturbed ND $\widetilde{B}(t)$.
Figure 18. $B(0)=1$ and $\widetilde{B}(0)=0.99$. Relative error $\delta(t)$ for the GDP-ND model.
7.2. Building heating
The second example is the ODE (9) of the building heating model. Assume, as in the instance (10), that there are three (negative) eigenvalues. We have
$K_{\infty}(y_{0},\widehat{z}_{0})=\frac{\left|\widehat{w}^{(1)}\widehat{z}_{0}\right|}{\left|\widehat{w}^{(1)}\widehat{y}_{0}\right|}\text{\ and\ }K_{\infty}(y_{0})=\frac{1}{\left|\widehat{w}^{(1)}\widehat{y}_{0}\right|}.$ (43)
For the particular instance (10) and the Euclidean norm as vector norm, we have the normalized rightmost left eigenvector
$\widehat{w}^{(1)}=\left[\begin{array}[]{c}-0.4462\ -0.6111\ -0.6538\end{array}\right].$
For the initial value and perturbed initial value in (11) and (12), we have $K_{\infty}(y_{0},\widehat{z}_{0})$ $=11.8648$. In Figure 7, we observe the relative error $\delta(t)=K(t,y_{0},\widehat{z}_{0})\cdot\delta(0)$ over approximately two characteristic times (6 hours is roughly twice the characteristic time of 3.17 hours). The relative error asymptotically approaches
$\lim\limits_{t\rightarrow+\infty}\delta(t)=K_{\infty}(y_{0},\widehat{z}_{0})\cdot\delta(0)=11.8648\cdot\delta(0)$
as $t$ increases. Moreover, we have $K_{\infty}(y_{0})=12.1330$. This indicates that the perturbation in (12) is nearly the worst-case scenario.
The expression for $K_{\infty}(y_{0})$ in (43) has a singularity when $y_{0}$ satisfies $\widehat{w}^{(1)}y_{0}=0$. For the instance (10), this occurs when (11) is replaced with
$y_{0}=\left(3.5\ ^{\circ}\mathrm{C},-5.2298\ ^{\circ}\mathrm{C},2.5\ ^{\circ}\mathrm{C}\right).$ (44)
In this case, where $y_{0}$ has zero component along the rightmost eigenvector, the asymptotic condition number $K_{\infty}(t,y_{0})$ grows exponentially in time (see [19]).
In Figure 19, for $y_{0}$ given in (44) and
$\widetilde{y}_{0}=\left(4\ ^{\circ}\mathrm{C},-5\ ^{\circ}\mathrm{C},3\ ^{\circ}\mathrm{C}\right),$
we observe the relative error $\delta(t)$ for $t\in[0,6\mathrm{h}]$. The difference compared to Figure 7 is quite evident. After $6\ \mathrm{h}$, the relative error is 235 times the initial relative error.
Figure 19. $\widehat{w}^{(1)}y_{0}=0$. Relative error $\delta(t)$ for the heating building model.
Remark 13.
In the instance (10), all the components of the left rightmost eigenvector $w^{(1)}$ have the same sign. This fact permits the following observation. If all the components of $y_{0}$ have the same sign, then
$K_{\infty}(y_{0})=\frac{\|y_{0}\|_{2}}{\left|\widehat{w}^{(1)}_{1}\right||y_{01}|+\left|\widehat{w}^{(1)}_{2}\right||y_{02}|+\left|\widehat{w}^{(1)}_{3}\right||y_{03}|}$
and then
$K_{\infty}(y_{0})\leq\frac{1}{\min\left\{\left|\widehat{w}^{(1)}_{1}\right|,\left|\widehat{w}^{(1)}_{2}\right|,\left|\widehat{w}^{(1)}_{3}\right|\right\}}=\frac{1}{0.4462}=2.2411.$
This shows that if the temperatures of the three floors are all larger or smaller than the equilibrium temperatures, then the relative error of the perturbed solution is not much larger than the initial relative error. This is not the case for the initial values (11) and (44), where the relative error grows significantly compared to its initial value.
Remark 14.
In the building heating model, as well as in the GDP-ND model, we have considered $y_{0}$ as the actual initial value and $\widetilde{y}_{0}$ as the initial value available to us for the simulation. Therefore, we have considered $\delta(t)$ as the relative error of the simulated solution with respect to the actual solution.
However, to obtain a computable asymptotic condition number $K_{\infty}(y_{0})$ for the heating building model, or a computable asymptotic condition number $K_{\infty}(y_{0},\widehat{z}_{0})$ for the GDP-ND model, one may consider $y_{0}$ as the available initial value and $\widetilde{y}_{0}$ as the actual initial value. Therefore, now we are considering $\delta(t)$ as the relative error of the actual solution with respect to the simulated solution.
In the case of the building heating model, we have $K_{\infty}(y_{0})=4.9195$ by considering (12) as $y_{0}$, instead of $K_{\infty}(y_{0})=12.1330$ by considering (11) as $y_{0}$.
In the case of the GDP-ND model, we have $K_{\infty}(y_{0},\widehat{z}_{0})=2.29$ for $B(0)=0.60$, instead of $K_{\infty}(y_{0},\widehat{z}_{0})=3.00$ for $B(0)=0.61$.
7.3. Another bulding heating model
The paper [23] presents a buildings heating model of dimension $4$, where the first state component is the indoor temperature and the other three are the temperatures in three layers of the wall separating the indoor space from the outdoor space. The matrix of the system of differential equations is
$A=\left[\begin{array}[]{rrrr}-5.7215&5.7215&0&0\\ 0.23076&-0.39276&0.162&0\\ 0&0.081&-0.162&0.081\\ 0&0&0.162&-0.91116\end{array}\right]$ (45)
when the time is measured in hours. The matrix has the four real negative eigenvalues $-0.038938\ \mathrm{h}^{-1}$, $-0.26104\ \mathrm{h}^{-1}$, $-0.92864\ \mathrm{h}^{-1}$ and $-5.9588\ \mathrm{h}^{-1}$. The characteristic time is $25.7\ \mathrm{h}$. The rightmost left eigenvector is
$w^{(1)}=\left[-0.029943\ -0.73735\ -1.1058\ -0.1027\right]$
Consider initial temperatures $y_{0}$ (these are temperatures with respect to the equilibrium temperatures) with $y_{02}=y_{03}=y_{04}=0$. We have, for a vector $p$-norm,
$K_{\infty}(y_{0})=\frac{1}{|\widehat{w}^{(1)}\widehat{y}_{0}|}=\frac{\|w^{(1)}\|}{|w^{(1)}_{1}|}.$
In Figure 20, for the vector $\infty$-norm, we see the condition numbers $K(t,y_{0})$ and $K_{\infty}(y_{0})=65.987$ for $t\in[0,30\ \mathrm{h}]$. The asymptotic behavior of the condition number is achieved within a characteristic time.
Figure 20. Condition numbers $K(t,y_{0})$ and $K_{\infty}(y_{0})$ for the matrix (45) and $y_{0}$ with $y_{02}=y_{03}=y_{04}=0$. The vector norm is the $\infty$-norm.
7.4. Charged particle subject to a magnetic fields with viscous frictional force
The first three examples described situations in which the matrix $A$ in (1) had real eigenvalues. This fourth example examines a situation in which the rightmost eigenvalues of $A$ form a complex conjugate pair.
We consider a charged particle subject to a magnetic field and moving in a medium with an anisotropic viscous frictional force (see the paper [6], which also considers Brownian motion and an electric field). Let $m$ and $q$ be the mass and the charge, respectively, of the particle, let $B=(B_{x},B_{y},B_{z})$ the constant magnetic field and let $\Gamma_{x}$, $\Gamma_{y}$ and $\Gamma_{z}$ be the constant coefficients of the viscous frictional force in the three spatial directions. The motion of the particle is described by the first order linear ODE for its velocity $v=(v_{x},v_{y},v_{z})$:
$v^{\prime}(t)=Av(t),$
where
$A=\frac{1}{m}\left[\begin{array}[]{rrr}-\Gamma_{x}&qB_{z}&-qB_{y}\\ -qB_{z}&-\Gamma_{y}&qB_{x}\\ qB_{y}&-qB_{x}&-\Gamma_{z}\end{array}\right].$
We examine the situation (comparable with the examples given in [6]) where, in suitable units,
$m=1$
$qB_{x}=-0.5,\ qB_{y}=-1.3,\ qB_{z}=-0.4,$
$\Gamma_{x}=0.4,\ \Gamma_{y}=0.8,\ \Gamma_{z}=0.2.$
In such a situation, the matrix $A$ has the eigenvalues
$\lambda_{1}=-0.3433+\sqrt{-1}\cdot 1.4326,\ \overline{\lambda_{1}}=-0.3433-\sqrt{-1}\cdot 1.4326,\ \lambda_{2}=-0.7133,$
with a complex conjugate pair $\lambda_{1}$ and $\overline{\lambda_{1}}$ as rightmost eigenvalues.
In the Euclidean norm, the normalized rightmost left eigenvector corresponding to $\lambda_{1}$ is
$\widehat{w}^{(1)}=\left[\begin{array}[]{rrr}0.6543+\sqrt{-1}\cdot 0.0111&-0.2821+\sqrt{-1}\cdot 0.1288&-0.1031-\sqrt{-1}\cdot 0.6819\end{array}\right].$
The oscillation scale factor of the asymptotic condition number $K_{\infty}(t,v_{0})$ is
$\mathrm{OSF}(v_{0})=\frac{1}{\left|\widehat{w}^{(1)}\widehat{v}_{0}\right|}=\frac{\|v_{0}\|_{2}}{\sqrt{\left(\mathrm{Re}(\widehat{w}^{(1)})v_{0}\right)^{2}+\left(\mathrm{Im}(\widehat{w}^{(1)})v_{0}\right)^{2}}},$
where $v_{0}$ is the initial velocity.
In Figure 21, we see, for $v_{0}=(0.5,1,0.5)$, over $6$ characteristic times $\frac{1}{-r_{1}}=2.9125$: the condition numbers $K(t,v_{0})$ and $K_{\infty}(t,v_{0})$, the oscillation scale factor $\mathrm{OSF}(v_{0})=5.92$ and the lower and upper bounds for $K_{\infty}(t,y_{0})$ derived by the lower and upper bounds (34)-(35) for the oscillating term. The asymptotic behavior is achieved in approximately three characteristic times.
Since $V_{1}=0.0587$ and $W_{1}=0.0937$ in this example, the oscillations of $K_{\infty}(t,y_{0})$ are tight. In fact, the oscillating term $\mathrm{OT}(t,y_{0})$ varies only slightly around $\sqrt{\frac{1}{2}}$ (the constant value of $\mathrm{OT}(t,y_{0})$ for $V_{1}=W_{1}=0$).
Figure 21. Condition numbers $K_{\infty}(t,v_{0})$ and $K_{\infty}(t,v_{0})$, $t\in[0,6\text{ characteristic time}]$, oscillating scale factor $\mathrm{OSF}(v_{0})$ and lower and upper bounds for $K_{\infty}(t,v_{0})$. The vector norm is the Euclidean norm.
7.5. An highly non-normal matrix $A$
In this example, we consider in the ODE (1) the highly non-normal matrix $A=VDV^{-1}$, where $V$ is the Hilbert matrix of order $8$ and
$D=\mathrm{diag}(-0.1,-0.2,\ldots,-0.8).$
The Euclidean norm is used as vector norm.
The high non-normality of the matrix $A$ appears in the large values $f_{j}$ defined in (38): we have
$f_{1}=5.2554\cdot 10^{5},\ f_{2}=1.677\cdot 10^{7},\ f_{3}=1.6347\cdot 10^{8},\ f_{4}=7.1819\cdot 10^{8},$
$f_{5}=1.6407\cdot 10^{9},\ f_{6}=2.0252\cdot 10^{9},\ f_{7}=1.2815\cdot 10^{9},\ f_{8}=3.2603\cdot 10^{8}.$
In Figure 22, we see the norm $\|\mathrm{e}^{tA}\|_{2}$ for $t\in[0,100]$, i.e. for $t$ over ten characteristic times $\frac{1}{-r_{1}}=10$. Note that $\|\mathrm{e}^{tA}\|_{2}$ is the condition number of the problem (2), when we consider absolute errors instead of relative errors. Despite $\|\mathrm{e}^{tA}\|_{2}\rightarrow 0$, $t\rightarrow+\infty$, this fact holds little significance. This is because, owing to the high non-normality of the matrix $A$, the norm remains large for numerous characteristic times and it only drops below $1$ after 13 characteristic times.
Hence, it is of little interest to ascertain that the absolute error of a perturbation in the initial value in the long-time decays to zero, as this “long-time” may be very distant. What we observe within a possible time span of our interest is a very large amplification of the absolute error of the perturbation.
Figure 22. Norm $\|\mathrm{e}^{tA}\|_{2}$ for $A=VDV^{-1}$, where $V$ is the Hilbert matrix of order $8$.
On the other hand, in Figure 23 we see, for two different initial values $y_{0}$ and for $t\in[0,100]$, the condition number $K(t,y_{0})$ of the problem (2), when we consider relative errors instead of absolute errors, and its asymptotic value
$K_{\infty}(y_{0})=\frac{1}{|\widehat{w}^{(1)}\widehat{y}_{0}|}.$
Similar pictures are obtained by randomly varying the initial value.
(a) $y_{0}=(1,1,1,1,1,1,1,1)$
(b) $y_{0}=(1,1,1,1,-1,-1,-1,-1)$
Figure 23. Condition numbers $K(t,y_{0})$ and $K_{\infty}(y_{0})$ for $A=VDV^{-1}$, where $V$ is the Hilbert matrix of order $8$. The vector norm is the Euclidean norm.
Note that the high non-normality does not have a critical impact, as is the case with $\|\mathrm{e}^{tA}\|_{2}$. In fact, we can observe the following.
  • 1)
    Although $K(t,y_{0})$ oscillates similarly to $\|\mathrm{e}^{tA}\|_{2}$ as $t$ varies, the maximum of $K(t,y_{0})$ has the same order of magnitude as $K_{\infty}(y_{0})$. The high value of $K_{\infty}(y_{0})$ in the top figure is due to a small value of $|\widehat{w}^{(1)}\widehat{y}_{0}|$, rather than high non-normality.
  • 2)
    The constant asymptotic behavior is achieved after a few (approximately five) characteristic times, similar to many examples where the matrix is not highly non-normal. For comparison, consider the values of $\|\mathrm{e}^{tA}\|_{2}$ in Figure 22 after five characteristic times. Indeed, the fact that the high non-normality does not critically impact the attainment of asymptotic behavior can be explained by observing that, when the characteristic time $\frac{1}{-r_{1}}$ is used as the time unit $\widehat{t}$, we have
    $\max\limits_{j\in\{2,\ldots,8\}}\frac{1}{(r_{1}-r_{j})\widehat{t}}\log\frac{f_{j}}{f_{1}}=\max\limits_{j\in\{2,\ldots,8\}}\frac{1}{\underset{=j}{\underbrace{\frac{-r_{j}}{-r_{1}}}}-1}\log\frac{f_{j}}{f_{1}}=3.4629.$
    in (39). Therefore, the high non-normality affects the achievement of the asymptotic behavior for only $3.4629$ characteristic times.
Whereas the fact that the absolute condition number $\|\mathrm{e}^{tA}\|_{2}$ decays to zero in the long-time is not significant, the fact that the relative condition number $K(t,y_{0})$ asymptotically becomes $K_{\infty}(y_{0})$, and always remains of the same order of magnitude as $K_{\infty}(y_{0})$, is highly significant.
7.6. An highly oscillating condition number $K_{\infty}(t,y_{0})$
Here, we consider a matrix $A$ of order 3 with a pair of imaginary conjugate simple eigenvalues as rightmost eigenvalues, and both $V_{1}$ and $W_{1}$ close to 1.
Let
$A=V\mathrm{diag}(\sqrt{-1},-\sqrt{-1},-1)V^{-1},$
where
$V=\left[\begin{array}[]{rrr}1+\sqrt{-1}&1-\sqrt{-1}&1\\ 1+\sqrt{-1}&1-\sqrt{-1}&0\\ 1+0.9\sqrt{-1}&1-0.9\sqrt{-1}&0\end{array}\right].$
The matrix $A$ is
$A=\left[\begin{array}[]{rrr}-1&20&-20\\ 0&19&-20\\ 0&18.1&-19\end{array}\right].$ (46)
Using the Euclidean norm as vector norm, we have $V_{1}=0.9988$ and $W_{1}=0.9986$.
Figure 24. Condition numbers $K(t,y_{0})$ and $K_{\infty}(t,y_{0})$, $t\in[0,4\pi]$, and oscillations scale factor $\mathrm{OSF}(y_{0})$, for $A$ in (46) and $y_{0}=(0,-0.7245,-0.6892)$ equal to the second right singular vector of $R_{1}$. The vector norm is the Euclidean norm.
In Figure 24, for $y_{0}$ equal to the second right singular vector of the matrix $R_{1}$ in (30) given by
$R_{1}=\left[\begin{array}[]{c}\mathrm{Re}(\widehat{w}^{(1)})\\ \mathrm{Im}(\widehat{w}^{(1)})\end{array}\right]=\left[\begin{array}[]{rrr}0&-0.4611&0.5123\\ 0&-0.5123&0.5123\end{array}\right],$
we present the condition numbers $K(t,y_{0})$ and $K_{\infty}(t,y_{0})$ along with the oscillation scale factor $\mathrm{OSF}(y_{0})$ for $t\in[0,4\pi]$, covering four periods of the function $K_{\infty}(t,y_{0})$. We observe large oscillations for the asymptotic condition number $K_{\infty}(t,y_{0})$.
Recall point 2 after Theorem 10. For this particular $y_{0}$, the oscillating scale factor is
$\mathrm{OSF}(y_{0})=\sqrt{\frac{2}{1-W_{1}}}=38.1\ .$
It is large, since $W_{1}$ is close to $1$. The oscillating term $\mathrm{OT}(t,y_{0})$ attains, as $t$ varies, the lower bound (34) and the upper bound (35): we have $a_{\mathrm{min}}(V_{1},W_{1})=0.0263\text{\ and \ }a_{\mathrm{max}}(V_{1},W_{1})=41.0.$ The lower bound is small and the upper bound is large, since $V_{1}$ and $W_{1}$ are close to 1. The minimum value and the maximum value for $K_{\infty}(t,y_{0})$ are
$\mathrm{OSF}(y_{0})\cdot a_{\mathrm{min}}(V_{1},W_{1})=1$
and
$\mathrm{OSF}(y_{0})\cdot a_{\mathrm{max}}(V_{1},W_{1})=\sqrt{\frac{1+W_{1}}{1-W_{1}}\cdot\frac{1+V_{1}}{1-V_{1}}}=1563,$
respectively.
Figure 25. Condition numbers $K(t,y_{0})$ and $K_{\infty}(t,y_{0})$, $t\in[0,4\pi]$, and oscillations scale factor $\mathrm{OSF}(y_{0})=1.003$, for $A$ in (46) and $y_{0}=(0,0.6892,-0.7245)$ equal to the first right singular vector of $R_{1}$. The vector norm is the Euclidean norm.
On the other hand, Figure 25 shows that even though $V_{1}$ and $W_{1}$ are close to 1, by choosing a different initial value $y_{0}$, we can ensure that $K_{\infty}(t,A)$ does not reach large values as $t$ varies. When $y_{0}$ is equal to the first right singular vector of the matrix $R_{1}$, instead of the second one, the condition number remains close to 1 (for details see [19]).
In both Figures 24 and 25, the asymptotic behavior is quickly attained. Regarding this, note that $f_{1}=23.5245$ and $\frac{f_{2}}{f_{1}}=0.0601$, and in (40) we have
$\frac{1}{(r_{1}-r_{2})\widehat{t}}\log\frac{f_{2}}{f_{1}}=-2.8115\text{\ \ and\ \ }\frac{1}{(r_{1}-r_{2})\widehat{t}}\cdot\frac{1}{2}\log\frac{1}{\sqrt{1-V_{1}}}=1.6844$
using $\widehat{t}=1$ as time unit.
7.7. A non-diagonalizable matrix $A$
We consider a real non-diagonalizable matrix $A$ of order $3$ with two distinct real eigenvalues: the defective eigenvalue $r_{1}$ with one Jordan mini-block of dimension $2$ and the non-defective eigenvalue $r_{2}$, where $r_{2}<r_{1}$. By recalling point 3 after Theorem 5, under the RLGE condition $w^{\left(1,1,2\right)}y_{0}\neq 0$ we have
$K_{\infty}(y_{0})=\frac{1}{\left|\widehat{w}^{\left(1,1,2\right)}\widehat{y}_{0}\right|},$
where
$\widehat{w}^{(1,1,2)}=\frac{w^{(1,1,2)}}{\|w^{(1,1,2)}\|}$
and $\|w^{(1,1,2)}\|$ is the induced norm of the row $w^{(1,1,2)}$.
We consider the case where $r_{1}=0$, $r_{2}=-1$,
$V=\left[\begin{array}[]{rrr}0&0&1\\ 1&0&-1\\ -1&1&0\end{array}\right]\text{\ \ and\ then\ \ }V^{-1}=\left[\begin{array}[]{ccc}1&1&0\\ 1&1&1\\ 1&0&0\end{array}\right].$
The matrix $A$ is
$A=\left[\begin{array}[]{rrr}-1&0&0\\ 2&1&1\\ -1&-1&-1\end{array}\right].$ (47)
[…]
$K_{\infty}(y_{0})=\frac{\|y_{0}\|_{1}}{|y_{01}+y_{02}+y_{03}|}=\frac{|y_{01}|+|y_{02}|+|y_{03}|}{|y_{01}+y_{02}+y_{03}|}$
for $t\in\left[0,50\right]$, for three different initial values $y_{0}$. We can observe an approach of $K(t,y_{0})$ to $K_{\infty}(y_{0})$, as $t\rightarrow+\infty$, slower than the previous examples. The curves $K(t,y_{0})$ appear, for $t$ close to $50$, nearly parallel to the lines representing their asymptotic value, with a small gap remaining between them. Indeed, we have an $\frac{1}{t}$-convergence of $K(t,y_{0})$ to $K_{\infty}(y_{0})$ as $t\rightarrow+\infty$ (see [19]), instead of the exponential convergence seen in the previous examples.
(a) $y_{0}=(-0.8,-2.9,1.4)$
(b) $y_{0}=(-1,1,0.05)$
(c) $y_{0}=(1,2,3)$
Figure 26. Condition numbers $K(t,y_{0})$ and $K_{\infty}(y_{0})$, $t\in[0,50]$ for the non-diagonalizable matrix (47). The vector norm is the $1$-norm.
8. Asymptotic and non-asymptotic conditioning
In this paper, we have studied the asymptotic conditioning of the problem (2), i.e. the conditioning for large $t$. On the other hand, we are also interested in the conditioning of (2) for any $t$, not only for large $t$.
Of course, if the problem (2) is asymptotically ill-conditioned, i.e $K_{\infty}(t,y_{0})$ assumes large values as $t$ varies, then the problem (2) is also ill-conditioned, i.e. $K(t,y_{0})$ assumes large values as $t$ varies.
Viceversa, can we say that if the problem (2) is asymptotically well-conditioned, i.e $K_{\infty}(t,y_{0})$ does not assume large values as $t$ varies, then the problem (2) is also well-conditioned, i.e. $K(t,y_{0})$ does not assume large values as $t$ varies?
Of course, the asymptotic behavior $K_{\infty}(t,y_{0})$ of $K(t,y_{0})$ cannot take into account for an initial increase of $K(t,y_{0})$ to values much larger than those of $K_{\infty}(t,y_{0})$. Thus, the answer to the previous question appears to be negative.
However, we now show experimentally for a real ODE (1) that such an initial increase in $K(t,y_{0})$ is very rare, and then the answer can indeed be considered positive. This stresses the crucial and central importance of the asymptotic condition number.
For $n=5,25,100$, we consider 10 0001000010\ 00010 000 instances of a real ODE (1) of dimension $n$, where the entries of $A$ and the components of the initial value $y_{0}$ are sampled from the standard normal distribution. Of these 10 0001000010\ 00010 000 instances, $5000$ have a matrix $A$ with a rightmost real eigenvalue and $5000$ have a matrix $A$ with a rightmost pair of complex conjugate eigenvalues. Since these are random instances, the rightmost eigenvalues are always simple.
We compute the ratios
$R=\frac{\max\limits_{t\in[0,T]}K(t,y_{0})}{\max\limits_{t\in[0,T]}K_{\infty}(t,y_{0})}$ (48)
for these 10 0001000010\ 00010 000 instances, where $T$ is fifty times the characteristic time $\widehat{t}=\frac{1}{|r_{1}|}$. The maxima are determined using $1000$ equally spaced sampling points over the interval $[0,T]$, i.e. with a stepsize $\frac{1}{20}$ of the characteristic time. Observe that, for the case of a real rightmost eigenvalue, we have $\max_{t\in[0,T]}K_{\infty}(t,y_{0})=K_{\infty}(y_{0})$, since $K_{\infty}(t,y_{0})=K_{\infty}(y_{0})$ remains constant over time $t$.
In the case of the Euclidean norm as vector norm, we see in Figures 27 and 28 the ratio $R$ for the $5000$ instances with a rightmost real eigenvalue and for the $5000$ instances with a rightmost pair of complex conjugate eigenvalues, respectively.
Figure 27. Ratio $R$ in (48) for $5000$ instances of $A$ with a rightmost real eigenvalue for $n=5,25,100$. The vector norm is the Euclidean norm.
Figure 28. Ratio $R$ in (48) for $5000$ instances of $A$ with a rightmost pair of complex conjugate eigenvalues for $n=5,25,100$. The vector norm is the Euclidean norm.
The next tables give statistical details for the ratio $R$: the first for the 5000 instances with the rightmost real eigenvalue, and the second for the 5000 instances with the rightmost pair of complex conjugate eigenvalues.
In the case of the $\infty$-norm and the $1$-norm as vector norms, we do not present figures but instead report the statistical details for $R$ in two similat tables as above.
$\infty$-norm:
$1$-norm:
From this experimentation, we can conclude that, at least for the vector $p$-norm, where $p\in\{1,2,\infty\}$, the ratio $R$ is very rarely much larger than $1$. Indeed, it is never much larger than $1$ in these 10 0001000010\ 00010 000 instances. Therefore, we have strong evidence that in the vast majority of cases
$\max\limits_{t\geq 0}K(t,y_{0})\ \text{and}\ \max\limits_{t\geq 0}K_{\infty}(t,y_{0})$ (55)
are of the same order of magnitude. Consequently, in the vast majority of cases, the problem (2) is well-conditioned if and only if it is asymptotically well-conditioned.
Thus, for the vector $p$-norm, $p\in\{1,2,\infty\}$, we have strong evidence that in the vast majority of cases of a rightmost real eigenvalue
$\max\limits_{t\geq 0}K(t,y_{0})\ \text{and}\ K_{\infty}(y_{0})=\frac{1}{|\widehat{w}^{(1)}\widehat{y}_{0}|}=\frac{\left\|w^{(1)}\right\|\left\|y_{0}\right\|_{2}}{|w^{(1)}y_{0}|}$ (56)
are of the same order of magnitude. Consequently, in the vast majority of cases of a rightmost real eigenvalue, the problem (2) is well-conditioned if and only if $|\widehat{w}^{1}\widehat{y}_{0}|$ is not small.
Moreover, for the Euclidean norm, by recalling Subsection 4.3.5 we have strong evidence that in the vast majority of cases of a rightmost complex conjugate pair of eigenvalues
$\max\limits_{t\geq 0}K(t,y_{0})\ \text{and}\ \max\limits_{t\geq 0}K_{\infty}(t,y_{0})\text{\ \ and\ \ }\mathrm{OSF}(y_{0})=\frac{1}{|\widehat{w}^{(1)}\widehat{y}_{0}|}=\frac{\left\|w^{(1)}\right\|_{2}\left\|y_{0}\right\|_{2}}{|w^{(1)}y_{0}|}$
are of the same order of magnitude. Consequently, in the vast majority of cases of a rightmost complex conjugate pair of eigenvalues, the problem (2) is well-conditioned if and only if $|\widehat{w}^{1}\widehat{y}_{0}|$ is not small.
8.1. Highly non-normal matrices
One might get the impression that the previous conclusion regarding what happens in the vast majority of cases does not hold for highly non-normal matrices, since sampling the entries of the matrix $A$ from the standard normal distribution rarely produces highly non-normal matrices. However, this is not the case, as demonstrated by the example of a highly non-normal matrix in Subsection 7.5, where the quantities in (56) are of the same order of magnitude.
To reinforce this conclusion, we consider, for $n=5,25,100$, a total of $5000$ instances of the real ODE (1) of dimension $n$, where the initial value $y_{0}$ has components drawn from the standard normal distribution, and the matrix $A$ is given by $A=QUQ^{T}$: $Q$ is the orthogonal matrix obtained from the QR factorization of a $n\times n$ random matrix with entries from the standard normal distribution, while $U$ is a $n\times n$ random upper triangular matrix with entries also drawn from the standard normal distribution. Generating matrices $A$ in this manner yields highly non-normal matrices with real eigenvalues.
Figure 29 and the following tables are analogous to Figures 27 and 28 and the subsequent tables. We now observe larger values of the ratio $R$, but they are rare. Overall, the global picture remains unchanged compared to the previous experiment.
Euclidean norm:
$\infty$-norm:
$1$-norm:
Figure 29. Ratio $R$ in (48) for the $5000$ instances of $A=QUQ^{T}$. The vector norm is the Euclidean norm.
9. Non-normal dynamics
Suppose we want to simulate the transient phase of a real-world system, where this transient $y$ satisfies an ODE (1) with a non-normal stable matrix $A$, where stable means that all the eigenvalues of $A$ have negative real part. For a non-linear real-world system, this ODE is the linearization around an asymptotically stable equilibrium of a non-linear ODE.
Due to the non-normality of the matrix $A$, it is expected that the transient exhibits an initial (possibly large) growth before decaying to zero. This phenomenon is what characterizes non-normal dynamics and can destabilize a non-linear system subject to small perturbation from the asymptotically stable equilibrium (see [22] and [4]).
Suppose that, through simulation, we need to determine the maximum growth of the transient, i.e. $\max_{t\geq 0}\|y(t)\|$. Additionally, suppose there is uncertainty in the initial value of our simulation. Here, we consider $y_{0}$ as the initial value available for the simulation and $\widetilde{y}_{0}$ as the actual initial value.
Since there are uncertainties in the initial value, it is important to assess how close the simulation result $\max_{t\geq 0}\|y(t)\|$ is to the actual maximum growth of the transient, i.e. $\max_{t\geq 0}\|\widetilde{y}(t)\|$.
Information on relative closeness can be obtained as follows. Since
$\frac{\|\widetilde{y}(t)\|}{\|y(t)\|}=1+\chi(t),$
where
$|\chi(t)|\leq\delta(t)=K(t,y_{0},\widehat{z}_{0})\varepsilon\leq K(t,y_{0})\varepsilon$
(see (15)), we have
$\left|\frac{\max\limits_{t\geq 0}\|\widetilde{y}(t)\|}{\max\limits_{t\geq 0}\|y(t)\|}-1\right|\leq E$
whenever $E<1$, where
$E=\max\limits_{t\geq 0}K(t,y_{0})\cdot\varepsilon.$ (60)
The quantity $E$ is an upper bound for the relative error of the actual maximum transient growth with respect to the simulated maximum transient growth, demonstrating how the condition number $K(t,y_{0})$ introduced in the present paper can be useful in the study of non-normal dynamics.
In the case of the vector $p$-norm, $p\in\{1,2,\infty\}$, the available evidence of Section 8, in particular that in Subsection 8.1, shows that in the vast majority of cases the quantities in (55) are of the same order of magnitude. Consequently, we conclude that, in the vast majority of cases, $E$ in (60) and
$E_{\infty}=\max\limits_{t\geq 0}K_{\infty}(t,y_{0})\cdot\varepsilon$ (61)
are of the same order of magnitude. Therefore, also the asymptotic condition number $K_{\infty}(t,y_{0})$ introduced and analyzed in this paper can be useful in the study of non-normal dynamics.
Example 15.
Consider an ODE (1) with
$A=\left[\begin{array}[]{rr}-1&a\\ 0&-2\end{array}\right],$ (62)
where $a>0$ is large. In this example, the $\infty$-norm is used as vector norm.
The matrix $A$ represents a classical example of an highly non-normal stable matrix (see the “quiz” at the beginning of [22]), exibiting a large initial growth of the solution of (1). In Figure 30, for $a=50,500,5000$, we see that $\|\mathrm{e}^{tA}\|_{\infty}$ initially grows significantly over time before eventually decaying to zero.
Since
$\left|\max_{t\geq 0}\|\widetilde{y}(t)\|_{\infty}-\max_{t\geq 0}\|y(t)\|_{\infty}\right|\leq\max_{t\geq 0}\|\mathrm{e}^{tA}\|_{\infty}\|\widetilde{y}_{0}-y_{0}\|_{\infty},$
it is expected that $\max_{t\geq 0}\|\widetilde{y}(t)\|_{\infty}$ will not be as close to $\max_{t\geq 0}\|y(t)\|_{\infty}$ as $\widetilde{y}_{0}$ is to $y_{0}$, when the closeness is measured in terms of absolute error.
On the other hand, as we will see below, this is not the case when the closeness is measured in terms of relative error instead of absolute error. Therefore, if controlling the relative error of $\max_{t\geq 0}\|\widetilde{y}(t)\|_{\infty}$ is important, considering the growth of the matrix exponential $\mathrm{e}^{tA}$ can be misleading, as it pertains to the absolute error rather than the relative error.
Figure 30. $\|\mathrm{e}^{tA}\|_{\infty}$, $t\in[0,10]$, for $a=50,500,5000$ in (62).
For $a=50,500,5000$, we present in the table below statistical details of the ratio
$R=\frac{\max\limits_{t\in[0,T]}K(t,y_{0})}{K_{\infty}(y_{0})}$
computed for 10 0001000010\ 00010 000 instance of $y_{0}$ sampled from the standard normal distribution. As in Section 8 (see the ratio (48)), we set $T$ to fifty times the characteristic time, i.e. $T=50$, and the maximum is determined using $1000$ equally spaced sampling points.
median ninth decile 99th percentile maximum value
$a=50$ $1.2629$ $3.3095$ $16.2544$ $46.806$
$a=500$ $1.039$ $1.1803$ $9.5118$ $275.9292$
$a=5000$ $1.0039$ $1.0159$ $1.1308$ $53.393$
This table confirms that, in the vast majority of cases for $y_{0}$, the maxima in (56) are of the same order of magnitude. Consequently, it validates that $E$ and $E_{\infty}$ in (60) and (61) are of the same order of magnitude.
We can explain this example in more detail. We have
$V=\left[\begin{array}[]{rr}1&-a\\ 0&1\end{array}\right]\text{\ \ and\ \ }V^{-1}=\left[\begin{array}[]{rr}1&a\\ 0&1\end{array}\right].$
Therefore,
$E_{\infty}=\frac{1}{|\widehat{w}^{(1)}\widehat{y}_{0}|}\cdot\varepsilon=\frac{\left\|w^{(1)}\right\|_{\infty}}{|w^{(1)}y_{0}|}\cdot\|\widetilde{y}_{0}-y_{0}\|_{\infty}=\frac{1+a}{|y_{01}+ay_{02}|}\cdot\max\{|\widetilde{y}_{01}-y_{01}|,|\widetilde{y}_{02}-y_{02}|\}.$
Since $a$ is large, for $\left|\frac{y_{01}}{y_{02}}\right|$ non-large (this is true in the vast majority of cases for $y_{0}$), $E_{\infty}$ is close to
$\frac{1}{|y_{02}|}\cdot\max\{|\widetilde{y}_{01}-y_{01}|,|\widetilde{y}_{02}-y_{02}|\}=\max\left\{\left|\frac{y_{01}}{y_{02}}\right|\cdot\frac{|\widetilde{y}_{01}-y_{01}|}{|y_{01}|},\frac{|\widetilde{y}_{02}-y_{02}|}{|y_{02}|}\right\},$
which has the same order of magnitude as the initial relative errors of the components. We conclude that, in the vast majority of cases for $y_{0}$, the relative error of $\max_{t\geq 0}\|\widetilde{y}(t)\|_{\infty}$ with respect to $\max_{t\geq 0}\|y(t)\|_{\infty}$ is not much larger than the initial relative errors of the components.
To confirm this conclusion, consider
$y_{0}=(1,1)\quad\text{and}\quad\widetilde{y}_{0}=(1.01,0.99).$
The following table presents the absolute and relative errors of $\max_{t\geq 0}\|\widetilde{y}(t)\|_{\infty}$ with respect to $\max_{t\geq 0}\|y(t)\|_{\infty}$.
Absolute error Relative error
$a=50$ $-0.1198$ $-0.0092$
$a=500$ $-1.2450$ $-0.0099$
$a=5000$ $-12.4949$ $-0.0100$
While the absolute error grows significantly with respect to the absolute errors ±$\pm 0.01$ of the components of $\widetilde{y}_{0}$, due to the large values attained by $\|\mathrm{e}^{tA}\|_{\infty}$ , the relative error, as expected, remains close to the relative errors ±$\pm 0.01$ of the components of $\widetilde{y}_{0}$.
10. Conclusion
The present paper explores how a perturbation of the initial value of the ODE (1) is propagated to the solution by examining the relative error, i.e. by comparing the perturbation of the solution to the solution itself. Considering the relative error, rather than the absolute error, provides a clearer perspective on the perturbation as the solution $y(t)$ evolves with time $t$: at any given time $t$, the perturbation of the solution $y(t)$ can be large (small) when compared to the initial value $y_{0}$, and simultaneously small (large) when compared to $y(t)$.
Understanding the behavior of the relative error of the perturbation can offer new insights and impact our approach to understanding the propagation of uncertainties. For example, the non-normality of the matrix $A$ appears to have a lesser effect on the relative error compared to the absolute error.
When we examine the absolute conditioning of the problem (2), the condition number is
$K_{\mathrm{abs}}(t)=\|\mathrm{e}^{tA}\|.$
Therefore, in the context of absolute error, a perturbation of the initial value grows in the worst-case scenario as a solution of $y^{\prime}(t)=Ay(t)$ grows in the worst-case scenario. There is nothing new: by knowing how the solutions grow, we also know how the perturbations grow, since the perturbations are themselves solutions. Moreover, the initial value $y_{0}$ does not play any role, since the condition number is independent of $y_{0}$.
On the other hand, when we consider the relative conditioning of the problem (2), the condition number is
$K_{\mathrm{rel}}(t,y_{0})=K(t,y_{0})=\frac{\|\mathrm{e}^{tA}\|\|y_{0}\|}{\|\mathrm{e}^{tA}y_{0}\|}.$
Consequently, in the context of relative error, the perturbations do not grow in the same manner as the solutions, but in a completely different way: compare fact A) at page 2 with fact B) at page 20. Moreover, unlike the absolute conditioning, in the relative conditioning the initial value $y_{0}$ plays a role, since the condition number depends on the direction of $y_{0}$.
These considerations represent the novelties in focusing on relative error rather than absolute error, offering new perspectives in linear dynamics, particularly in non-normal dynamics, when simulating the transient behavior of a real system and there is uncertainty in the initial value of the transient.
The mathematical results of the paper, included in Sections 3 to 6, analyze the asymptotic (long-time) behavior of the relative error of the perturbed solution. This asymptotic behavior also provides information on the non-asymptotic behavior (for all times, not only for long-time). In fact, the strong experimental evidence in Section 8 suggests that, in the vast majority of cases, the maximum values over time $t$ of the condition number $K(t,y_{0})$ and its asymptotic behavior $K_{\infty}(t,y_{0})$ share the same order of magnitude.
The main result of the paper can be summarized as follows. Consider a real ODE (1) and the Euclidean norm as vector norm. In the vast majority of cases for (1), the relative error of the perturbed initial value is magnified in the perturbed solution, in the worst-case scenario for the initial perturbation, by a factor of the order of magnitude of
$\frac{\left\|w^{(1)}\right\|_{2}\left\|y_{0}\right\|_{2}}{|w^{(1)}y_{0}|},$
where the row vector $w^{(1)}$ is a left eigenvector of the matrix $A$ corresponding to the rightmost real or complex eigenvalue.
The author hopes that this paper will stimulate attention to the role of relative error in dynamic contexts.
Acknowledgements: the research was supported by the INdAM Research group GNCS (Gruppo Nazionale di Calcolo Scientifico).