### v2

# Long-time relative error analysis for linear ordinary differential equations with perturbed initial value

Abstract.
We investigate the propagation of initial value perturbations along the solution of a linear ordinary differential equation $y^{\prime}(t)=Ay(t)$. This propagation is analyzed using the relative error rather than the absolute error. Our focus is on the long-term behavior of this relative error, which differs significantly from that of the absolute error. The present paper is a practical sequel to the theoretical papers [18, 19] on the long-time behavior of the relative error: it includes applicative examples and important issues not addressed in [18, 19]. In addition, the present paper shows that understanding the long-term behavior provides insights into the growth of the relative error over all times, not just at large times. Therefore, it represents a crucial and fundamental aspect of the conditioning of linear ordinary differential equations, with applications in, for example, non-normal dynamics.
Keywords: linear ordinary differential equations, matrix exponential, relative error, asymptotic behavior, condition numbers, non-normal dynamics.
MSC2020 classification: 15A12, 15A16, 15A18, 15A21, 34A30, 34D05.
1. Introduction
Consider a linear $n$-dimensional Ordinary Differential Equation (ODE)
$\left\{\begin{array}[]{l}y^{\prime}\left(t\right)=Ay\left(t\right),\ t\in\mathbb{R},\\ y\left(0\right)=y_{0},\end{array}\right$. (1)
where $A\in\mathbb{R}^{n\times n}$ and $y_{0},y(t)\in\mathbb{R}^{n}$. In the present paper, we are interested in understanding how a perturbation of the initial value $y_{0}$ is propagated along the solution $y(t)=\mathrm{e}^{tA}y_{0}$ of (1) over a long time interval.
The next fact A) is well known.
  • A)
    For a generic perturbation of $y_{0}$, the perturbation of $y(t)$ asymptotically (as $t\rightarrow+\infty$) vanishes exponentially if the rightmost eigenvalues of $A$ have negative real part and asymptotically diverges exponentially if such eigenvalues have positive real part.
Thus, if we are concerned about how large the perturbation can become for large $t$, we should be reassured by knowing that the rightmost eigenvalues of $A$ have negative real part and truly concerned by knowing that these eigenvalues have positive real part.
However, for a generic $y_{0}$, also $y(t)$ asymptotically vanishes exponentially if the rightmost eigenvalues of $A$ have negative real part and asymptotically diverges exponentially if such eigenvalues have positive real part. Thus, to know that the perturbation of the solution asymptotically vanishes, or diverges, exponentially does not help us understand if it is really significant, when compared to the solution.
Comparing the perturbation of the solution to the solution itself means considering the relative error of the perturbed solution. The well-known fact A) is a description of the long-time behavior of the absolute error of the perturbed solution. It is important also to give a similar description of the long-time behavior of the relative error. This description is fact B) on page 17.
Understanding the long-time behavior of the relative error of the perturbed solution is fundamental in real-world systems described by mathematical models based on linear ODEs (1), which require simulation through the integration of these ODEs in the presence of uncertainty in the initial value (for example due to measurement errors). This is particularly true when the solution becomes small or large, compared to the initial value, a situation where looking at the absolute error of the perturbed solution can have little significance. In Subsection 1.2 below, we examine two such models and emphasize the importance of considering the relative error of the perturbed solution in cases of uncertainty in the initial value.
Put differently, we are interested in the relative conditioning of the problem
$y_{0}\mapsto\mathrm{e}^{tA}y_{0}$ (2)
for large $t$, i.e. we are interested in studying how a relative error in the input $y_{0}$ is propagated to the output $\mathrm{e}^{tA}y_{0}$ for large $t$.
The present paper is a sequel to the theoretical papers [18, 19] on the long-time relative conditioning of the problem (2) and it contains applications to real-world systems, experimental tests, and other practical issues related to the results in [18, 19].
1.1. Literature
In the literature, the relative conditioning of the matrix exponential function is a well-studied topic (see, e.g., [14], [21], [13], [15], [20], [2], [24], [3], [8], and [1]). Some of these papers are mainly focused on computational aspects and algorithms, while others consider how the relative conditioning of the problem $A\mapsto\mathrm{e}^{tA}$ depends on $t$. However, for this problem, a general characterization of the long-time behavior of the relative conditioning is still not known.
In contrast, the relative conditioning of the action of the matrix exponential $\mathrm{e}^{tA}$ on a vector with respect to perturbations of this vector, i.e., the relative conditioning of the problem (2), has received little attention. The reason could be that (2) is perceived as an easy linear problem, without all the complications involved in the non-linearity of the matrix exponential function problem. However, once one examines the dynamics of the relative conditioning, the issue is no longer straightforward (see [18, 19]).
A study of how the conditioning of the problem (2) depends on $t$ was given in the paper [17], but the analysis was confined to the case of a normal matrix $A$. The paper [18] extends this study to a general complex linear ODE and [19] delves into the results of [18] for the real case.
1.2. Two models
In many cases, a mathematical model based on a linear ODE (1) is constructed after a careful selection of the parameters appearing in the entries of the matrix $A$ by the model creators. As a result, the matrix $A$ can be regarded as fixed and reliable. The model is then used repeatedly with different initial conditions, often by external users who can have little control over the accuracy of these initial conditions. In this context, where the initial conditions are inputs to a fixed model, the analysis of the relative error of the perturbed solution can be of interest.
We show the importance of understanding the behavior of the relative error of the perturbed solution through two mathematical models based on a linear ODE (1).
1.2.1. Gross Domestic Product and National Debt
The papers [9, 10] present a mathematical model for the Gross Domestic Product (GDP) and National Debt (ND) of a given country. This model is based on the linear system of ODEs
$\left\{\begin{array}[]{l}Q^{\prime}\left(t\right)=a_{11}Q\left(t\right)+a_{12}B\left(t\right)\\ B^{\prime}\left(t\right)=a_{21}Q\left(t\right)+a_{22}B\left(t\right),\end{array}\right$. (3)
where $Q(t)$ is the GDP and $B(t)$ is the ND at the time $t$. The unit for time is 1yr1\ $\mathrm${yr} and the unit for GDP and ND is the initial GDP, i.e., $Q(0)=1$. A possible instance for the coefficients (given in [9]) is
$a_{11}=0.08,\ a_{12}=-0.07,\ a_{21}=0.03,\ a_{22}=-0.02.$ (4)
For this instance, the matrix $A$ of the ODE has the positive eigenvalues $0.05$ and $0.01$.
Suppose we are interested in simulating the growth of GDP and ND over a period of 50yr50\ $\mathrm${yr} by integrating (3) for a given initial ND (remember that the initial GDP is set at $1$). We assume, as in [9], that the initial ND is $0.60$. Due to uncertainty in the available data, this initial ND may not be the actual value. Suppose the actual initial ND is $0.61$. We set $B(0)$ to this actual initial ND of $0.61$, while our assumed initial value of $0.60$ for the simulation is the perturbed value $\widetilde{B}(0)$. (Conversely, we could set $B(0)=0.60$ and $\widetilde{B}(0)=0.61$. See Remark 14).
Since the eigenvalues of $A$ are positive, the perturbation of the initial ND grows exponentially in the solution. After 50yr50\ $\mathrm${yr}, for $(Q(0),B(0))=(1,0.61)$, we have
$(Q(50),B(50))=(8.84,4.09)$
and, for the perturbed $(\widetilde{Q}(0),\widetilde{B}(0))=(1,0.60)$, we have
$(\widetilde{Q}(50),\widetilde{B}(50))=(9.02,4.15).$
Here, we denote the perturbed solution of (3) by $(\widetilde{Q}(t),\widetilde{B}(t))$.
Now, the question is:
  • after 50yr50\ $\mathrm${yr}, is the effect of the initial uncertainty significant?
There are two ways to answer this.
  • 1)
    The perturbation of the initial ND is $1$ second decimal figure (from $61$ to $60$) and the perturbation of GDP and ND after 50yr50\ $\mathrm${yr} is $18$ second decimal figure and $6$ second decimal figure (from $884$ to $902$ and from $409$ to $415$), respectively . This demonstrates a significant increase in the perturbation, as an effect of exponential growth. In Figure 1, we see in the plane $(Q,B)$ the points $(Q(0),B(0))$ and $(\widetilde{Q}(0),\widetilde{B}(0))$ on the left and the points $(Q(50),B(50))$ and $(\widetilde{Q}(50),\widetilde{B}(50))$ on the right. On the left and on the right, we use the same scale with both axes of length 1 GDP unit.
    Figure 1. Left: points $(Q(0),B(0))$ and $(\widetilde{Q}(0),\widetilde{B}(0))$. Right: points $(Q(50),B(50))$ and $(\widetilde{Q}(50),\widetilde{B}(50))$.
    The growth of the perturbation is evident. In Figure 2, we see the absolute error of the perturbed solution
    $e(t)=\left\|(\widetilde{Q}\left(t\right),\widetilde{B}(t))-\left(Q\left(t\right),B(t)\right)\right\|_{2},$ (5)
    i.e. the Euclidean distance between the points $(\widetilde{Q}\left(t\right),\widetilde{B}(t))$ and $\left(Q\left(t\right),B(t)\right)$, for $t\in[0,50]$.
    Figure 2. Absolute error $e(t)$ for the GDP-ND model.
    After 50yr50\ $\mathrm${yr}, the absolute error is about $20$ times the initial absolute error.
  • 2)
    The perturbation of the initial ND is $1$ second significant figure (from $61$ to $60$) and the perturbation of GDP and ND after 50yr50\ $\mathrm${yr} is $1.8$ second significant figure and $0.6$ second significant figure (from $88.4$ to $90.2$ and from $40.9$ to $41.5$), respectively. This shows that the perturbed GDP and ND are still close to the unperturbed values after 50yr50\ $\mathrm${yr}. In Figure 3, similarly to Figure 1, we see the points $(Q(0),B(0))$ and $(\widetilde{Q}(0),\widetilde{B}(0))$ on the left and the points $(Q(50),B(50))$ and $(\widetilde{Q}(50),\widetilde{B}(50))$ on the right, but now on the left we use a scale with axes of length $\left\|(Q(0),B(0))\right\|_{2}$ and on the right we use a scale with axes of length $\left\|(Q(50),B(50))\right\|_{2}$.
    Figure 3. Left: points $(Q(0),B(0))$ and $(\widetilde{Q}(0),\widetilde{B}(0))$. Right: points $(Q(50),B(50))$ and $(\widetilde{Q}(50),\widetilde{B}(50))$.
    The closeness of $(Q(50),B(50))$ and $(\widetilde{Q}(50),\widetilde{B}(50))$ is comparable to the closeness of $(Q(0),B(0))$ and $(\widetilde{Q}(0),\widetilde{B}(0))$. This closeness is also confirmed by Figure 4, where it appears that GDP and perturbed GDP, as well as ND and perturbed ND, plotted as functions of time over $50$ $\mathrm{yr}$ are barely distinguishable.
    Figure 4. GDP $Q(t)$, perturbed GDP $\widetilde{Q}(t)$, ND $B(t)$ and perturbed ND $\widetilde{B}(t)$.
    In Figure 5, we see the relative error of the perturbed solution
    $\delta(t)=\frac{\left\|(\widetilde{Q}\left(t\right),\widetilde{B}(t))-\left(Q\left(t\right),B(t)\right)\right\|_{2}}{\left\|\left(Q\left(t\right),B(t)\right)\right\|_{2}}$ (6)
    for $t\in[0,50]$. It is the distance between the points $(\widetilde{Q}\left(t\right),\widetilde{B}(t))$ and $\left(Q\left(t\right),B(t)\right)$ when, as in Figure 3, we use the length $\left\|(Q(t),B(t))\right\|_{2}$ of the axes as unit length.
    Figure 5. Relative error $\delta(t)$ for the GDP-ND model.
    Unlike the exponentially growing absolute error (5), the relative error (6) grows much less. After 50yr50\ $\mathrm${yr}, it is about $2$ times the initial relative error.
If one believes that answer 2) is more appropriate (meaning that after 50 years, the effect of the initial perturbation is not significant because the perturbation of the solution is not much larger than its initial value, when compared to the solution), it becomes crucial to understand the behavior of the relative error (6) rather than that of the absolute error (5).
The relative error (6) refers to the relative error of the solution vector $(Q(t),B(t))$, i.e., it is a normwise relative error. One might find it more interesting to consider the relative errors of the two components $Q(t)$ and $B(t)$. In the specific integration over $50$ $\mathrm{yr}$ we are considering, the relative errors of the components do not have order of magnitude larger than that of the relative error (6), since the components of the solution are not small compared to the Euclidean norm of the solution: we see in Figure 3 that the point $(Q(50),B(50))$ is not close to the axes. It is noteworthy that in Figure 4, $Q(t)$ and $\widetilde{Q}(t)$, as well as $B(t)$ and $\widetilde{B}(t)$, are nearly indistinguishable due to small componentwise relative errors.
About componentwise relative errors, see Subsection 2.1 below.
This mathematical model of GDP and ND will be revisited in Section 6 in light of the results presented in this paper.
1.2.2. Building heating
This example (see the book [12]) is a toy mathematical model of building heating. Similar more complex non-toy models are used in literature for modeling real building heating (e.g., see [16], [7] and [23]).
Consider a building constituted by basement, main floor and attic. Let $x_{1}(t)$, $x_{2}(t)$ and $x_{3}(t)$ be the temperatures of basement, main floor and attic, respectively, at the time $t$. By using Newton’s law of cooling
rate of change of internal temperature
$\propto\text{external temperature $-$ internal temperature,}$
we derive the linear system of ODEs
$\left\{\begin{array}[]{l}x_{1}^{\prime}\left(t\right)=k_{g1}\left(x_{g}-x_{1}\left(t\right)\right)+k_{12}\left(x_{2}\left(t\right)-x_{1}\left(t\right)\right)+f_{1}\\ x_{2}^{\prime}\left(t\right)=k_{o2}\left(x_{o}-x_{2}\left(t\right)\right)+k_{12}\left(x_{1}\left(t\right)-x_{2}\left(t\right)\right)+k_{23}\left(x_{3}\left(t\right)-x_{2}\left(t\right)\right)+f_{2}\\ x_{3}^{\prime}\left(t\right)=k_{o3}\left(x_{o}-x_{3}\left(t\right)\right)+k_{23}\left(x_{2}\left(t\right)-x_{3}\left(t\right)\right)+f_{3},\end{array}\right$. (7)
where
  • $x_{g}$ is temperature of the ground and $x_{o}$ is the outdoor temperature of the air;
  • $k_{g1},k_{o2},k_{o3},k_{12},k_{23}$ are positive proportionality constants in the Newton’s law depending on the thermal insulation of the floors;
  • $f_{1}$, $f_{2}$ and $f_{3}$ are constant forcing terms due to heaters in the basement, main floor and attic, respectively.
The ODE (7) has the form
$x^{\prime}(t)=Ax(t)+b,$ (8)
where
$A=\left[\begin{array}[]{ccc}-k_{g1}-k_{12}&k_{12}&0\\ k_{12}&-k_{a2}-k_{12}-k_{23}&k_{23}\\ 0&k_{23}&-k_{a3}-k_{23}\end{array}\right]\text{\ \ and\ \ }b=\left[\begin{array}[]{ccc}k_{g1}x_{g}+f_{1}\\ k_{a2}x_{a}+f_{2}\\ k_{a3}x_{a}+f_{3}\end{array}\right].$
Since the matrix $A$ is symmetric and strictly diagonally dominant with negative diagonal entries, it has real negative eigenvalues. Consequently, the ODE (8) has a globally asymptotically stable equilibrium point
$x_{eq}=-A^{-1}b.$
The transient
$y(t)=x(t)-x_{eq}$
satisfies the ODE
$y^{\prime}(t)=Ay(t).$ (9)
Suppose we want to simulate the transient by integrating (9), and there is uncertainty in the initial temperatures $y_{0}$ at time $t=0$. Of course, $y_{0}$ and $y(t)$ are temperatures with respect to the equilbrium temperatures.
Therefore, we have an initial value $y_{0}$, representing the actual initial temperatures, and a perturbed initial value $\widetilde{y}_{0}$, representing the initial temperatures available to us for simulation. (Alternatively, we could assume that $y_{0}$ represents the available temperatures and $\widetilde{y}_{0}$ represents the actual temperatures. See Remark 14).
For illustrating our considerations, in (7) set
$k_{g1}=0.5\ \frac{\frac{{}^{\circ}\mathrm{C}}{\mathrm{h}}}{{}^{\circ}\mathrm{C}},\ k_{o2}=0.25\ \frac{\frac{{}^{\circ}\mathrm{C}}{\mathrm{h}}}{{}^{\circ}\mathrm{C}},k_{o3}=0.25\ \frac{\frac{{}^{\circ}\mathrm{C}}{\mathrm{h}}}{{}^{\circ}\mathrm{C}},\ k_{12}=0.5\ \frac{\frac{{}^{\circ}\mathrm{C}}{\mathrm{h}}}{{}^{\circ}\mathrm{C}},\ k_{23}=1\ \frac{\frac{{}^{\circ}\mathrm{C}}{\mathrm{h}}}{{}^{\circ}\mathrm{C}}$ (10)
[…]
$y_{0}=\left(3.5\ ^{\circ}\mathrm{C},-4.4\ ^{\circ}\mathrm{C},2.5\ ^{\circ}\mathrm{C}\right)$ (11)
and
$\widetilde{y}_{0}=\left(4\ ^{\circ}\mathrm{C},-4\ ^{\circ}\mathrm{C},3\ ^{\circ}\mathrm{C}\right).$ (12)
In Figures 6 and 7 we show the absolute error
$e(t)=\left\|\widetilde{y}(t)-y(t)\right\|_{2}$
and the relative error
$\delta(t)=\frac{\left\|\widetilde{y}(t)-y(t)\right\|_{2}}{\left\|y(t)\right\|_{2}},$
where $\widetilde{y}$ is the perturbed (simulated) solution of (9), for $t\in\left[0,6\ \mathrm{h}\right]$.
Figure 6. Absolute error $e(t)$ for the building heating model.
Figure 7. Relative error $\delta(t)$ for the building heating model.
Here, the scenario is completely reversed compared to the GDP-ND model. Due to the negative eigenvalues, the absolute error decreases exponentially and the initial perturbation is reduced by about seven times in $6$ hours, reassuring us about the effect of the uncertainty in the initial temperatures, when the transient is simulated by integrating (9). However, this is true only when comparing the perturbation to the initial value $y_{0}$, i.e., by looking at the absolute error. When comparing the perturbation to the solution $y(t)$, i.e., by looking at the relative error, the situation is much more concerning, as the initial perturbation is magnified by about twelve times. This could significantly impact our understanding of the transient and the decisions we make based on the simulation.
For example, suppose we want to determine when the Euclidean norm of the solution drops below 0.5C0.5\ $^{\circ}\mathrm${C}. When this occurs, we could decide that the transient phase has ended. The simulation with initial value $\widetilde{y}_{0}$ shows that this occurs at time $\widetilde{t}^{\ast}=3.0876\ \mathrm{h}$. However, at this time $\widetilde{t}^{\ast}$, the relative error of the perturbed solution is greater than 1 (see Figure 7), meaning that the norm of the perturbation is larger than the norm of the actual solution. This implies that the norm $\|\widetilde{y}(\widetilde{t}^{\ast})\|_{2}=0.5^{\circ}\mathrm{C}$ of the perturbed solution could be more than twice the norm $\|y(\widetilde{t}^{\ast})\|_{2}$ of the actual solution. Indeed, we have
$\frac{\|\widetilde{y}(\widetilde{t}^{\ast})\|_{2}}{\|y(\widetilde{t}^{\ast})\|_{2}}=2.3882.$
The norm of the actual solution, which is less than half the norm 0.5C0.5\ $^{\circ}\mathrm${C} of the perturbed solution at time $\widetilde{t}^{\ast}$, dropped below 0.5C0.5\ $^{\circ}\mathrm${C} at $t^{\ast}=1.6362\ \mathrm{h}$, in roughly half the time. Therefore, the simulation with the initial temperatures $\widetilde{y}_{0}$ available to us provides a highly inaccurate estimate of the end transient time. This is not only due to poor accuracy in the initial temperatures $\widetilde{y}_{0}$ (the relative error of $\widetilde{y}_{0}$ with respect to $y_{0}$ is $\delta(0)=0.1320$), but also to inherent ill-conditioning (the relative error of $\widetilde{t}^{\ast}$ with respect to $t^{\ast}$ is $0.8871$).
Note that, to understand how the ratio $\frac{\|\widetilde{y}(t)\|_{2}}{\|y(t)\|_{2}}$ varies over time and, consequently, how different $\widetilde{t}^{\ast}$ and ${t}^{\ast}$ can be, it is natural to consider the relative error $\delta(t)$. In fact, we have
$\frac{\|\widetilde{y}(t)\|_{2}}{\|y(t)\|_{2}}=1+\xi(t),$
where $\xi(t)$ satisfies $|\xi(t)|\leq\delta(t)$. Consequently, $t^{\ast}$ and $\widetilde{t}^{\ast}$ are such that
$\|y(t^{\ast})\|_{2}=0.5\ ^{\circ}\mathrm{C}\text{\ and\ }\|y(\widetilde{t}^{\ast})\|_{2}=\frac{0.5\ ^{\circ}\mathrm{C}}{1+\xi(\widetilde{t}^{\ast})}$, (13)
[…]
Similarly to the previous GDP-ND model, we will revisit this building heating model in Section 6.
1.3. The asymptotic behavior of the relative error
The present paper deals with the asymptotic behavior, i.e., the long-time behavior, of the relative error of the perturbed solution.
One could observe that, in case of a solution that decays to zero or diverges, this analysis might have limited relevance since asymptotically the solution is zero or infinite and then a long-time simulation of the solution is not very interesting. However, in such a case we might be interested in simulating the solution only up to a certain size threshold, beyond which it has become too small or too large to be of further interest. Therefore, if the relative error becomes close to, or of the same order of magnitude as, its asymptotic behavior before the solution has reached this threshold, then the analysis of the asymptotic behavior has interest.
In the GDP-ND and building heating models, although the solution diverges to infinity or converges to zero, we are still interested in understanding its evolution before it becomes too large or too small. The relative error after 50 years in the GDP-ND model and the relative error at the end transient time $\widetilde{t}^{\ast}$ in the building heating model (see (13)) have the same order of magnitude as their asymptotic values: see Figure 13 and Figure 7.
In the present paper, in addition to studying the asymptotic behavior of the relative error, we also investigate how rapidly this asymptotic behavior is attained. As expected, the non-normality of the matrix $A$ adversely affects the rapid attainment of the asymptotic behavior. However, as it will be shown, a high non-normality of $A$ does not necessarily imply a late onset of the asymptotic behavior, which could lead to a loss of interest in such behavior.
In conclusion, we can say that although the asymptotic behavior of the relative error does not fully describe the propagation to the solution of the perturbation of the initial value, since it can miss a possible initial growth of the relative error to values much larger than the asymptotic behavior, it nonetheless constitutes an important piece in the qualitative study of the relative conditioning of the problem (2).
However, Section 8 shows experimentally that it is very rare to have such a large initial growth of the relative error. Therefore, the asymptotic behavior might be not only an important piece of the study of conditioning, but the crucial and fundamental piece.
1.4. Plan of the paper
The asymptotic analysis of the relative error of the perturbed solution given in the papers [18, 19] is long and the details quite technical. This is due to the fact that [18] addressed the general case involving an arbitrary matrix $A$ in (1), a choice that necessitates dealing with the Jordan Canonical Form of $A$ and generalized eigenvectors. Moreover, the presence of complex eigenvalues further complicates matters in [19]. Indeed, by considering only diagonalizable matrices with real eigenvalues, the analysis would be considerably shorter. Finally, the definition of asymptoticity used in [18] also implies some effort in proving asymptotic results.
Due to their lengths, the papers [18, 19] only include theoretical results. All the non-theoretical practical issues are moved to the present paper, which is organized as follows.
Section 2 introduces two condition numbers presented in [18] for the problem (2). Section 3 recalls the results of [19] regarding the asymptotic behavior of these condition numbers in a generic case for the ODE (1). Section 4 recalls the results of [19] on the closeness of the condition numbers to their asymptotic behavior, as a function of time. Section 5 focuses on the more important of the two condition numbers and addresses fundamental questions such as asymptotic well-conditioning, the onset of asymptotic behavior, and the effect of the non-normality of the matrix $A$. In Section 6, the GDP-ND and building heating models are revisited and three additional examples illustrate the contents of the previous sections. Section 7 shows that the asymptotic behavior of the relative error can also provide insight into the non-asymptotic behavior in most cases. Section 8 illustrates how the results of the present paper can be applied in non-normal dynamics. Section 9 presents the conclusions and can even be read now to gain a better idea of the subject of this paper.
The paper presents several numerical experiments in which a large number of random instances of the ODE (1) are generated by sampling the entries of $A$ and the components of $y_{0}$ from the standard normal distribution. This means that all entries of $A$ and all components of $y_{0}$ in every instance are independently sampled. It is worth noting that the entries of $A$ are drawn from the standard normal distribution rather than from the normal distribution with mean zero and standard deviation $\frac{1}{\sqrt{n}}$, which is often employed. We choose not to use the $\frac{1}{\sqrt{n}}$ scaling because we aim to include a broad range of values for the norm of the matrix $A$, thereby generating potentially stiffer or more extreme instances of the ODE. Nevertheless, performing the experiments with the scaled version yields essentially the same results and, consequently, leads to the same conclusions.
All numerical experiments are carried out in MATLAB, with the matrix exponential values $\mathrm{e}^{tA}$ computed using the expm function.
2. The condition numbers
Assume that the initial value $y_{0}\neq 0$ of (1) is perturbed to $\widetilde{y}_{0}$ and then the solution $y$ is perturbed to $\widetilde{y}$. Fixed an arbitrary vector norm $\left\|\ \cdot\ \right\|$, we are interested in relating the normwise relative error
$\varepsilon:=\frac{\left\|\widetilde{y}_{0}-y_{0}\right\|}{\left\|y_{0}\right\|}$ (14)
of the perturbed initial value $\widetilde{y}_{0}$ to the normwise relative error
$\delta\left(t\right):=\frac{\left\|\widetilde{y}\left(t\right)-y\left(t\right)\right\|}{\left\|y\left(t\right)\right\|}$ (15)
of the perturbed solution $\widetilde{y}(t)$. Observe that $\delta(0)=\varepsilon$. Moreover, in the case of the Euclidean norm as vector norm, observe that $\delta(t)$ is the distance between $\widetilde{y}(t)$ and $y(t)$ when $\|y(t)\|_{2}$ is used as scale unit (see Figure 8).
Remark 1.
Observe that the normwise relative error (15) is invariant under multiplication of the vector norm by a constant. Therefore, these relative errors remain the same for both the $p$-norm and the mean $p$-norm,
$\|x\|_{p,\text{mean}}=\left(\sum\limits_{i=1}^{n}\frac{1}{n}|x_{i}|^{p}\right)^{\frac{1}{p}},\quad x\in\mathbb{C}^{n}.$
Using such mean norms avoids the expected growth of the $p$-norms with the dimension $n$.
Figure 8. The relative error $\delta(t)$ when $\|y(t)\|_{2}$ is the scale unit.
By writing the perturbed initial value as
$\widetilde{y}_{0}=y_{0}+\varepsilon\|y_{0}\|\widehat{z}_{0},$
where $\widehat{z}_{0}\in\mathbb{C}^{n}$ is a unit vector (i.e., $\|\widehat{z}_{0}\|=1$) indicating the direction of perturbation, we obtain
$\delta\left(t\right)=K\left(t,y_{0},\widehat{z}_{0}\right)\cdot\varepsilon,$ (16)
where
$K\left(t,y_{0},\widehat{z}_{0}\right):=\frac{\left\|\mathrm{e}^{tA}\widehat{z}_{0}\right\|}{\left\|\mathrm{e}^{tA}\widehat{y}_{0}\right\|}$
with $\widehat{y}_{0}:=\frac{y_{0}}{\left\|y_{0}\right\|}$ the normalized initial value. We define $K(t,y_{0},\widehat{z}_{0})$ as the directional pointwise condition number of the problem (2).
In general, we know nothing about the direction $\widehat{z}_{0}$ of the perturbation of $y_{0}$. Therefore, it is useful to introduce
$K\left(t,y_{0}\right):=\max\limits_{\begin{subarray}{c}\widehat{z}_{0}\in\mathbb{C}^{n}\\ \|\widehat{z}_{0}\|=1\end{subarray}}K\left(t,y_{0},\widehat{z}_{0}\right)=\frac{\left\|\mathrm{e}^{tA}\right\|}{\left\|\mathrm{e}^{tA}\widehat{y}_{0}\right\|},$
where $\left\|\mathrm{e}^{tA}\right\|$ is the matrix norm of $\mathrm{e}^{tA}$ induced by the vector norm $\left\|\ \cdot\ \right\|$. We define $K(t,y_{0})$ as the pointwise condition number of the problem (2) (see [5] for the definition of condition number of a general problem). It is the worst condition number $K(t,y_{0},\widehat{z}_{0})$ as the direction of perturbation $\widehat{z}_{0}$ varies.
2.1. Normwise and componentwise relative errors
We have introduced the normwise relative errors (14) and (15), but it may also be worthwhile to consider the componentwise relative errors
$\varepsilon_{l}=\frac{\left|\widetilde{y}_{0l}-y_{0l}\right|}{\left|y_{0l}\right|},\ l\in\{1,\ldots,n\},$
of the perturbed initial value and
$\delta_{l}\left(t\right)=\frac{\left|\widetilde{y}_{l}\left(t\right)-y_{l}\left(t\right)\right|}{\left|y_{l}\left(t\right)\right|},\ l\in\{1,\ldots,n\},$
of the perturbed solution, where $y_{0l}$, $\widetilde{y}_{0l}$, $y_{l}(t)$ and $\widetilde{y}_{l}(t)$, $l\in\{1,\ldots,n\}$, are the components of $y_{0}$, $\widetilde{y}_{0}$, $y(t)$ and $\widetilde{y}(t)$, respectively. Componentwise relative errors were considered in [11], for $A$ diagonalizable.
We can derive information regarding the componentwise relative errors from the normwise relative errors. In fact, if the vector norm $\left\|\ \cdot\ \right\|$ is a $p$-norm, then :
  • 1)
    $\varepsilon\leq\max\limits_{l\in\{1,\ldots,n\}}\varepsilon_{l};$
  • 2)
    $\delta(t)\leq\max\limits_{l\in\{1,\ldots,n\}}\delta_{l}\left(t\right);$
  • 3)
    $\delta_{l}\left(t\right)\leq\frac{\|y(t)\|}{|y_{l}(t)|}\delta(t),\ l\in\{1,\ldots,n\}.$
In particular, point 3) is useful for estimating the order of magnitude of the relative error of perturbed components $\widetilde{y}_{l}(t)$, once the order of magnitude of $\delta(t)$ is known (recall the observation regarding relative errors of the two components $Q(t)$ and $B(t)$ of the GDP-ND model of Subsection 1.2.1). In light of this, an important question is to understand how frequently components with a large ratio $\frac{\|y(t)\|}{|y_{l}(t)|}$ appear.
The next example demonstrates that for a randomly selected matrix $A$, the number of components with a large ratio $\frac{\|y(t)\|_{\infty}}{|y_{l}(t)|}$ is not a substantial percentage of the total number of components.
Example 2.
Figure 9. Maximum ratio $R(t)$ in (17) for 10 00010\ 000 random instances of (1) with $n=100$.
In Figure 9, we see for $t=0.1,1,10$ the maximum ratio
$R(t)=\max\limits_{l\in\{1,\ldots,n\}}\frac{\|y(t)\|_{\infty}}{|y_{l}(t)|}$ (17)
in logarithmic scale, for 10 00010\ 000 instances of (1), where $A$ of order $n=100$ and $y_{0}$ have entries sampled from the standard normal distribution. The ratio $R(t)$ is large for almost all instances.
However, in Figure 10, we see, for all 10 00010\ 000 instances, the fraction
$r(t,M)=\frac{\text{number of components $y_{l}(t)$, $l\in\{1,\ldots,n\}$, such that $\frac{\|y(t)\|_{\infty}}{|y_{l}(t)|}>M$}}{n}$ (18)
for $M=10$ on the left side and $M=100$ on the right side. For $M=10$, the fraction is less than $50\%$ in all instances and around $25\%$ on average. For $M=100$, it is less than $10\%$ in all instances and under $5\%$ on average.
Figure 10. Fraction $r(t,M)$ in (18) for the 10 00010\ 000 random instances of Figure 9.
In conclusion, analyzing the normwise relative error can provide valuable insights into the componentwise relative errors. In fact, we have evidence that a substantial percentage of the componentwise relative errors of the solution have an order of magnitude not larger than that of the normwise relative error, when the $\infty$-norm is used as vector norm.
3. Asymptotic behavior of condition numbers
We consider the spectrum of the matrix $A$ as partitioned into the sets $\Lambda_{j}$, $j\in\{1,\ldots,q\}$, where $\Lambda_{j}$ contains all the eigenvalues with the same real part $r_{j}$ and $r_{1}>r_{2}>\cdots>r_{q}$ holds. Observe that $\Lambda_{1}$ is the set of the rightmost eigenvalues of $A$.
The theorems presented in this section can be found in [19] and concern the generic case for the real matrix $A$ where the set $\Lambda_{1}$ of the rightmost eigenvalues of $A$ consists of either a real simple eigenvalue or a single pair of complex conjugate simple eigenvalues.
3.1. Notations
We introduce some notations.
  • 1)
    In the case where $\Lambda_{1}$ consists of a real simple eigenvalue $\lambda_{1}$, let $w^{(1)}$ (a row real vector) be a left eigenvector corresponding to $\lambda_{1}$ and let $\widehat{w}^{(1)}=\frac{w^{(1)}}{\|w^{(1)}\|}$ be its normalization. Here,
    $\|w^{\left(1\right)}\|=\max\limits_{\begin{subarray}{c}u\in\mathbb{R}^{n}\\ \|u\|=1\end{subarray}}|w^{\left(1\right)}u|$
    is the real induced norm of the real matrix $w^{(1)}$.
    We say that $\Lambda_{1}$ is simple single real when $\Lambda_{1}$ consists of a real simple eigenvalue.
  • 2)
    In the case where $\Lambda_{1}$ consists of a single pair of complex conjugate simple eigenvalues $\lambda_{1}$ and $\overline{\lambda_{1}}$, with $\lambda_{1}$ having positive imaginary part $\omega_{1}$, let $w^{(1)}$ (a complex row vector) and $v^{(1)}$ (a complex column vector) be left and right, respectively, eigenvectors corresponding to $\lambda_{1}$ such that $w^{(1)}v^{(1)}=1$ and let $\widehat{w}^{(1)}=\frac{w^{(1)}}{\|w^{(1)}\|}$ and $\widehat{v}^{(1)}=\frac{v^{(1)}}{\|v^{(1)}\|}$ be their normalizations. Here,
    $\|w^{\left(1\right)}\|=\max\limits_{\begin{subarray}{c}u\in\mathbb{C}^{n}\\ \|u\|=1\end{subarray}}|w^{\left(1\right)}u|$
    is the complex induced norm of the complex matrix $w^{(1)}$.
    Moreover, given the polar forms
    $\widehat{v}_{k}^{\left(1\right)}=\left|\widehat{v}_{k}^{\left(1\right)}\right|\mathrm{e}^{\sqrt{-1}\alpha_{1k}},\ k\in\{1,\ldots,n\},$
    $\widehat{w}_{l}^{\left(1\right)}=\left|\widehat{w}_{l}^{\left(1\right)}\right|\mathrm{e}^{\sqrt{-1}\beta_{1l}},\ l\in\{1,\ldots,n\},$
    of the components of the complex vectors $\widehat{v}^{\left(1\right)}$ and $\widehat{w}^{\left(1\right)}$ and the polar form
    $\widehat{w}^{\left(1\right)}u=\left|\widehat{w}^{\left(1\right)}u\right|\mathrm{e}^{\sqrt{-1}\gamma_{1}\left(u\right)}$
    of the complex scalar $\widehat{w}^{\left(1\right)}u$, where $u\in\mathbb{R}^{n}$, we introduce the vector
    $\widehat{\Theta}_{1}\left(t,u\right):=\left(\left|\widehat{v}_{k}^{\left(1\right)}\right|\cos\left(\omega_{1}t+\alpha_{1k}+\gamma_{1}\left(u\right)\right)\right)_{k=1,\ldots,n}\in\mathbb{R}^{n}$
    and the matrix
    $\widehat{\Theta}_{1}\left(t\right)=\left(\left|\widehat{v}_{k}^{\left(1\right)}\right|\left|\widehat{w}_{l}^{\left(1\right)}\right|\cos\left(\omega_{1}t+\alpha_{1k}+\beta_{1l}\right)\right)_{k,l=1,\ldots,n}\in\mathbb{R}^{n\times n}.$
    We say that $\Lambda_{1}$ is simple single complex when $\Lambda_{1}$ consists of a single complex conjugate pair of simple eigenvalues.
3.2. Asymptotic forms
For the description of the asymptotic behavior of the condition numbers, we use the following notion of asymptotic form. Let $a(t)$ and $b(t)$ be real functions of $t\in\mathbb{R}$. We say that $b(t)$ is an asymptotic form of $a(t)$ and write
$a\left(t\right)\sim b\left(t\right),\ t\rightarrow+\infty,$
if
$\lim\limits_{t\rightarrow+\infty}\frac{a(t)}{b(t)}=1.$
In other words, $b(t)$ is an asymptotic form of $a(t)$ if
$\lim\limits_{t\rightarrow+\infty}\chi(t)=0,$
where $\chi(t)$ is the relative error of $a(t)$ with respect to $b(t)$.
3.3. The asymptotic behavior of $K(t,y_{0},\widehat{z}_{0})$ and $K(t,y_{0})$
Next two results (see [19]) describe the asymptotic forms of the condition numbers $K(t,y_{0},\widehat{z}_{0})$ and $K(t,y_{0})$ in the generic case for $A$, described at the beginning of this Section 3, and the generic case for $y_{0}$ and $\widehat{z}_{0}$ given by $w^{(1)}y_{0}\neq 0$ and $w^{(1)}\widehat{z}_{0}\neq 0$.
Theorem 3.
Assume $\Lambda_{1}$ simple single real. For $y_{0}$ and $\widehat{z}_{0}$ such that $w^{(1)}y_{0}\neq 0$ and $w^{(1)}\widehat{z}_{0}\neq 0$, we have
$K\left(t,y_{0},\widehat{z}_{0}\right)\sim K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right)=K_{\infty}\left(y_{0},\widehat{z}_{0}\right):=\frac{\left|\widehat{w}^{(1)}\widehat{z}_{0}\right|}{\left|\widehat{w}^{(1)}\widehat{y}_{0}\right|},\ t\rightarrow+\infty,$
and
$K(t,y_{0})\sim K_{\infty}\left(t,y_{0}\right)=K_{\infty}\left(y_{0}\right):=\frac{1}{\left|\widehat{w}^{(1)}\widehat{y}_{0}\right|},\ t\rightarrow+\infty.$
Theorem 4.
Assume $\Lambda_{1}$ simple single complex. For $y_{0}$ and $\widehat{z}_{0}$ such that $w^{(1)}y_{0}\neq 0$ and $w^{(1)}\widehat{z}_{0}\neq 0$, we have
$K\left(t,y_{0},\widehat{z}_{0}\right)\sim K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right)=\mathrm{OSF}(y_{0},\widehat{z}_{0})\cdot\mathrm{OT}(t,y_{0},\widehat{z}_{0}),\ t\rightarrow+\infty,$
and
$K(t,y_{0})\sim K_{\infty}\left(t,y_{0}\right)=\mathrm{OSF}(y_{0})\cdot\mathrm{OT}(t,y_{0}),\ t\rightarrow+\infty,$
where
$\mathrm{OSF}(y_{0},\widehat{z}_{0}):=\frac{\left|\widehat{w}^{(1)}\widehat{z}_{0}\right|}{\left|\widehat{w}^{(1)}\widehat{y}_{0}\right|}\text{\ \ and\ \ }\mathrm{OSF}(y_{0}):=\frac{1}{\left|\widehat{w}^{(1)}\widehat{y}_{0}\right|}$
and
$\mathrm{OT}(t,y_{0},\widehat{z}_{0}):=\frac{\left\|\widehat{\Theta}_{1}\left(t,\widehat{z}_{0}\right)\right\|}{\left\|\widehat{\Theta}_{1}\left(t,\widehat{y}_{0}\right)\right\|}\text{\ \ and\ \ }\mathrm{OT}(t,y_{0}):=\frac{\left\|\widehat{\Theta}_{1}\left(t\right)\right\|}{\left\|\widehat{\Theta}_{1}\left(t,\widehat{y}_{0}\right)\right\|}.$ (19)
The constants $\mathrm{OSF}(y_{0},\widehat{z}_{0})$ and $\mathrm{OSF}(y_{0})$ are called oscillation scale factors and $\mathrm{OT}(t,y_{0},\widehat{z}_{0})$ and $\mathrm{OT}(t,y_{0})$, which are periodic functions of $t$ of period $\frac{\pi}{\omega_{1}}$, are called oscillating terms. Note that the oscillation scale factors depend on the moduli, and the oscillating terms on the angles, in the polar forms of the complex numbers $\widehat{w}^{(1)}\widehat{y}_{0}$ and $\widehat{w}^{(1)}\widehat{z}_{0}$.
3.4. The case $\Lambda_{1}$ simple single complex and the Euclidean norm
In this subsection, in the case of $\Lambda_{1}$ simple single complex, we analyze the oscillating terms when the vector norm is the Euclidean norm.
3.4.1. $V_{1}$ and $W_{1}$
We introduce
$V_{1}:=\left|\left(\widehat{v}^{(1)}\right)^{T}\widehat{v}^{(1)}\right|\text{\ \ and\ \ }W_{1}:=\left|\widehat{w}^{(1)}\left(\widehat{w}^{(1)}\right)^{T}\right|,$
moduli of the complex numbers $\left(\widehat{v}^{(1)}\right)^{T}\widehat{v}^{(1)}$ and $\widehat{w}^{(1)}\left(\widehat{w}^{(1)}\right)^{T}$. Here T, unlike H, denotes pure transposition without conjugation. We have $V_{1},W_{1}\in[0,1)$.
In Figure 11, we see, for $n=5,25,100$, the set of the pairs $(V_{1},W_{1})$ for 50 00050\ 000 random matrices $A$ of order $n$ such that $\Lambda_{1}$ is simple single complex. The elements of $A$ are sampled from the standard normal distribution. Observe that the distribution of the pairs in the square $0,1)^{2}$ is not uniform, since the pairs tend to accumulate around the diagonal $V_{1}=W_{1}$. The next table shows, for the 50 00050\ 000 random instances, the distribution of the distance $|V_{1}-W_{1}|$.
$|V_{1}-W_{1}|<\frac{1}{4}$ $\frac{1}{4}\leq|V_{1}-W_{1}|<\frac{1}{2}$ $\frac{1}{2}\leq|V_{1}-W_{1}|<\frac{3}{4}$ $\frac{3}{4}\leq|V_{1}-W_{1}|$
$n=5$ 88% 11% 1% 0.003%
$n=25$ 85% 14% 1% 0.002%
$n=100$ 86% 13% 1% 0.002%
Figure 11. Pairs $(V_{1},W_{1})$ for 50 00050\ 000 random matrices $A$ of order $n=5,25,100$.
3.4.2. $\mathrm{OT}(t,y_{0},\widehat{z}_{0})$ and $\mathrm{OT}(t,y_{0})$
The next two theorems (see [19]) specify upper and lower bounds of the oscillating terms $\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)$ in (19) and $\mathrm{OT}\left(t,y_{0}\right)$ in (19).
Theorem 5.
For $y_{0}$ and $\widehat{z}_{0}$ such that $w^{(1)}y_{0}\neq 0$ and $w^{(1)}\widehat{z}_{0}\neq 0$, we have
$\sqrt{\frac{1-V_{1}}{1+V_{1}}}\leq\mathrm{OT}\left(t,y_{0},\widehat{z}_{0}\right)\leq\sqrt{\frac{1+V_{1}}{1-V_{1}}},\ t\in\mathbb{R}.$
Observe that if $V_{1}$ is not close to 1, then, for any $y_{0}$ and $\widehat{z}_{0}$, $\mathrm{OT}(t,y_{0},\widehat{z}_{0})$ does not assume large or small values as $t$ varies. Therefore, if $V_{1}$ is not close to 1, then, for any $y_{0}$ and $\widehat{z}_{0}$, the values of $\mathrm{OT}(t,y_{0},\widehat{z}_{0})$ as $t$ varies have the order of magnitude $1$. Consequently, the values of $K_{\infty}(t,y_{0},\widehat{z}_{0})$ as $t$ varies have the same order of magnitude as $\mathrm{OSF}(y_{0},\widehat{z}_{0})$.
In light of Theorems 3, 4 and 5, we can state a fact B) regarding the relative error of the perturbed solution, analogous to fact A) concerning the absolute error of the perturbed solution and presented at the beginning of the paper.
  • B)
    For a generic $y_{0}$ and for a generic perturbation of $y_{0}$, the relative error of the perturbed solution asymptotically (as $t\rightarrow+\infty$) neither diverges nor decays to zero. Instead, it converges to a non-zero constant, in case of a real rightmost eigenvalue, and to a periodic oscillating function bounded and uniformly away from zero, in case of a complex conjugate pair of rightmost eigenvalues.
Theorem 6.
For $y_{0}$ such that $w^{(1)}y_{0}\neq 0$, we have
$a_{\min}(V_{1},W_{1})\leq\mathrm{OT}(t,y_{0})\leq a_{\max}(V_{1},W_{1}),\ t\in\mathbb{R},$
where
$a_{\min}(V_{1},W_{1}):=\left\{\begin{array}[]{l}\sqrt{\frac{(1+W_{1})(1-V_{1})}{2(1+V_{1})}}\text{\ if\ }V_{1}\leq W_{1}\\ \\ \sqrt{\frac{1-W_{1}}{2}}\text{\ if\ }V_{1}\geq W_{1},\end{array}\right$. (20)
and
$a_{\max}(V_{1},W_{1}):=\sqrt{\frac{(1+W_{1})(1+V_{1})}{2(1-V_{1})}}.$ (21)
Observe that if $V_{1}$ is not close to $1$, then, for any $y_{0}$, $\mathrm{OT}(t,y_{0})$ does not assume large values or small values as $t$ varies. Therefore, if $V_{1}$ is not close to 1, then, for any $y_{0}$, the values of $\mathrm{OT}(t,y_{0})$ as $t$ varies have the order of magnitude $1$. Consequently, the values of $K_{\infty}(t,y_{0})$ as $t$ varies have the same order of magnitude as $\mathrm{OSF}(y_{0})$.
3.4.3. Is $V_{1}$ close to $1$?
We have seen in the observations following Theorems 5 and 6, that $K_{\infty}(t,y_{0},\widehat{z}_{0})$ and $K_{\infty}(t,y_{0})$ have the same order of magnitude as the oscillation scale factors when $V_{1}$ is not close to 1, since, in this case, the oscillating terms have the order of magnitude $1$. Therefore, it is of interest to understand how frequently $V_{1}$ is close to 1.
The following table shows, for the 50 00050\ 000 random matrices of Figure 11, the percentages of cases with $V_{1}$ greater than $0.9$, $0.99$, $0.999$ and $0.9999$.
$V_{1}>0.9$ $V_{1}>0.99$ $V_{1}>0.999$ $V_{1}>0.9999$
$n=5$ 7% 0.7% 0.009% 0.008%
$n=25$ 5% 0.5% 0.03% 0.002%
$n=100$ 3% 0.3% 0.02% 0.002%
We have strong evidence that $V_{1}$ is rarely close to 1.
Regarding the order of magnitude of the oscillating terms when $V_{1}$ is close to $1$, note that
$\frac{1}{k}\leq\sqrt{\frac{1-V_{1}}{1+V_{1}}}\leq\mathrm{OT}(t,y_{0},\widehat{z}_{0})\leq\sqrt{\frac{1+V_{1}}{1-V_{1}}}\leq k$
and
$\frac{1}{k}\leq a_{\mathrm{min}}(V_{1},W_{1})\leq\mathrm{OT}(t,y_{0})\leq a_{\mathrm{max}}(V_{1},W_{1})\leq k,$
where
$k=\sqrt{\frac{2}{1-V_{1}}}.$
The values of $k$ for $V_{1}=0.9,0.99,0.999,0.9999$ (see table (22) above) are
$V_{1}=0.9$ $V_{1}=0.99$ $V_{1}=0.999$ $V_{1}=0.9999$
$k$ $4.4721$ $14.1421$ $44.7214$ $141.4214$
4. The onset of the asymptotic behavior
As mentioned in the introduction, it is important to understand when the condition numbers $K(t,y_{0},\widehat{z}_{0})$ and $K(t,y_{0})$ begin to get close to their asymptotic forms $K_{\infty}(t,y_{0},\widehat{z}_{0})$ and $K_{\infty}(t,y_{0})$. In this section, we address this question. The results presented here can be found in [19].
We assume that the matrix $A$ in (1) satisfies a condition a little bit stronger than the generic condition of the previous section, where we assumed that $\Lambda_{1}$ is single simple real. Here, we assume that, for any $j\in\{1,\ldots,q\}$, $\Lambda_{j}$ is simple single real or simple single complex, i.e., it consists of either a real simple eigenvalue or a single pair of complex conjugate simple eigenvalues. This latter condition, nonetheless, remains a generic condition for the matrix $A$.
4.1. Notations
We introduce for $\Lambda_{j}$, $j\in\{1,\ldots,q\}$, notations similar to the notations 1) and 2) introduced for $\Lambda_{1}$ in Subsection 3.1.
  • 1)
    In the case $\Lambda_{j}$ simple single real, let $w^{(j)}$ and $v^{(j)}$ be left and right, respectively, eigenvectors corresponding to the real eigenvalue in $\Lambda_{j}$ such that $w^{(j)}v^{(j)}=1$. The normalized vector $\widehat{w}^{(j)}$ is defined similarly to $\widehat{w}^{(1)}$.
  • 2)
    In the case $\Lambda_{j}$ simple single complex, let $w^{(j)}$ and $v^{(j)}$ be left and right, respectively, eigenvectors corresponding to the eigenvalue in the complex conjugate pair with positive imaginary part such that $w^{(j)}v^{(j)}=1$.
    Moreover, we introduce the normalized vectors $\widehat{w}^{(j)}$ and $\widehat{v}^{(j)}$, the vector $\widehat{\Theta}_{j}(t,u)\in\mathbb{R}^{n}$, with $u\in\mathbb{R}^{n}$, the matrix $\widehat{\Theta}_{j}(t)\in\mathbb{R}^{n\times n}$ and the scalars $V_{j},W_{j}\in[0,1)$, defined similarly to $\widehat{w}^{(1)}$, $\widehat{v}^{(1)}$, $\widehat{\Theta}_{1}(t,u)$, $\widehat{\Theta}_{1}(t)$, $V_{1}$, and $W_{1}$.
We also introduce the values
$f_{j}:=\|w^{(j)}\|\cdot\|v^{(j)}\|\in[1,+\infty),\ j\in\{1,\ldots,q\}.$ (23)
4.2. Approximation with a given precision
The closeness of the condition numbers to their asympototic forms is measured by using the following notion of approximation with a given precision. For $a,b\in\mathbb{R}$ and $\epsilon\geq 0$, we write
$a\approx b\text{\ with precision\ }\epsilon$
if
$a=b(1+\chi)\text{\ with\ }|\chi|\leq\epsilon,$
where $\chi$ is the relative error of $a$ with respect to $b$.
4.3. Closeness to the asymptotic behavior
The next theorem (see [19]) describes how the condition numbers get closer to their asymptotic versions, as $t$ increases.
Theorem 7.
Assume that, for any $j\in\{1,\ldots,q\}$, $\Lambda_{j}$ is simple single real or simple single complex. For $y_{0}$ and $\widehat{z}_{0}$ such that $w^{(1)}y_{0}\neq 0$ and $w^{(1)}\widehat{z}_{0}\neq 0$, we have
$K\left(t,y_{0},\widehat{z}_{0}\right)\approx K_{\infty}\left(t,y_{0},\widehat{z}_{0}\right)\text{\ with precision\ }\frac{\epsilon(t,\widehat{z}_{0})+\epsilon(t,\widehat{y}_{0})}{1-\epsilon(t,\widehat{y}_{0})}$
and
$K\left(t,y_{0}\right)\approx K_{\infty}\left(t,y_{0}\right)\text{\ with precision\ }\frac{\epsilon(t)+\epsilon(t,\widehat{y}_{0})}{1-\epsilon(t,\widehat{y}_{0})}$ (24)
whenever $\epsilon(t,\widehat{y}_{0})<1$, where
$\epsilon(t,u):=\sum\limits_{j=2}^{q}\mathrm{e}^{\left(r_{j}-r_{1}\right)t}\frac{f_{j}}{f_{1}}\cdot\frac{\left|\widehat{w}^{(j)}u\right|}{\left|\widehat{w}^{(1)}u\right|}G_{j}\left(t,u\right),\ u\in\mathbb{R}^{n},$
and
$\epsilon(t):=\sum\limits_{j=2}^{q}\mathrm{e}^{\left(r_{j}-r_{1}\right)t}\frac{f_{j}}{f_{1}}G_{j}(t).$
Here, $G_{j}(t,u)$ and $G_{j}(t)$, $j\in\{2,\ldots,q\}$, are defined as follows.
  • 1)
    If both $\Lambda_{j}$ and $\Lambda_{1}$ are simple single real, then
    $G_{j}(t,u)=1\text{\ \ and\ \ }G_{j}(t)=1.$
  • 2)
    If $\Lambda_{j}$ is simple single complex and $\Lambda_{1}$ is simple single real, then
    $G_{j}(t,u)=2\left\|\widehat{\Theta}_{j}(t,u)\right\|\text{\ \ and\ \ }G_{j}(t)=2\left\|\widehat{\Theta}_{j}(t)\right\|.$
    If the vector norm is a $p$-norm, then
    $G_{j}(t,u)\leq 2\text{\ \ and\ \ }G_{j}(t)\leq 2.$
  • 3)
    If $\Lambda_{j}$ is simple single real and $\Lambda_{1}$ is simple single complex, then
    $G_{j}(t,u)=\frac{1}{2\left\|\widehat{\Theta}_{1}(t,u)\right\|}\text{\ \ and\ \ }G_{j}(t)=\frac{1}{2\left\|\widehat{\Theta}_{1}(t)\right\|}.$
    If the vector norm is the Euclidean norm, then
    $G_{j}(t,u)\leq\sqrt{\frac{1}{2(1-V_{1})}}\text{\ \ and\ \ }G_{j}(t)\leq\sqrt{\frac{1}{a(V_{1},W_{1})}},$
    where
    $a(V_{1},W_{1}):=\left\{\begin{array}[]{l}(1-V_{1})(1+W_{1})\text{\ if\ }V_{1}\leq W_{1}\\ \\ (1+V_{1})(1-W_{1})\text{\ if\ }V_{1}\geq W_{1}.\end{array}\right$.
  • 4)
    If both $\Lambda_{j}$ and $\Lambda_{1}$ are simple single complex, then
    $G_{j}(t,u)=\frac{\left\|\widehat{\Theta}_{j}(t,u)\right\|}{\left\|\widehat{\Theta}_{1}(t,u)\right\|}\text{\ \ and\ \ }G_{j}(t)=\frac{\left\|\widehat{\Theta}_{j}(t)\right\|}{\left\|\widehat{\Theta}_{1}(t)\right\|}.$
    If the vector norm is the Euclidean norm, then
    $G_{j}(t,u)\leq\sqrt{\frac{1+V_{j}}{1-V_{1}}}\text{\ \ and\ \ }G_{j}(t)\leq\sqrt{\frac{(1+W_{j})(1+V_{j})}{a(V_{1},W_{1})}}.$
In conclusion, the onset of the asymptotic behavior, i.e., when the condition numbers begin to get close to their asymptotic forms, is determined by how rapidly $\epsilon(t,u)$, $u=\widehat{y}_{0},\widehat{z}_{0}$, and $\epsilon(t)$ reach small values. The factors influencing this can be found in the expressions for $\epsilon(t,u)$ and $\epsilon(t)$ in the previous Theorem 7. They are, for $j\in\{2,\ldots,q\}$:
  • the decreasing exponentials $\mathrm{e}^{(r_{j}-r_{1})t}$;
  • the ratios $\frac{f_{j}}{f_{1}}$;
  • the ratios $\frac{\left|\widehat{w}^{(j)}u\right|}{\left|\widehat{w}^{(1)}u\right|}$;
  • the numbers $G_{j}(t,u)$ and $G_{j}(t)$.
An exploration of such factors influencing how rapidly $K(t,y_{0})$ gets close to $K_{\infty}(t,y_{0})$ is presented in the next section.
5. Some important considerations
This section discusses two important issues concerning the asymptotic condition number $K_{\infty}(t,y_{0})$. It is the most important between the two asymptotic condition numbers since, in general, we have no information about the direction of perturbation $\widehat{z}_{0}$.
As in the previous section, we assume the generic case of a real matrix $A$ in (1) such that, for any $j\in\{1,\ldots,q\}$, $\Lambda_{j}$ is simple single real or simple single complex, and of an initial value $y_{0}$ such that $w^{(1)}y_{0}\neq 0$.
It is of interest to know:
  • AWC)
    for which initial values $y_{0}$ the problem (2) is asymptotically well-conditioned, i.e., $K_{\infty}(t,y_{0})$ does not assume large values as $t$ varies;
  • ONS)
    the onset of the asymptotic behavior, i.e., when $K(t,y_{0})$ begins to be close to $K_{\infty}(t,y_{0})$; as we have seen, this happens as soon as $\epsilon(t,\widehat{y}_{0})$ and $\epsilon(t)$ are small.
5.1. The case $\Lambda_{1}$ simple single real
Suppose $\Lambda_{1}$ simple single real. We have
$K_{\infty}(y_{0})=\frac{1}{\left|\widehat{w}^{(1)}\widehat{y}_{0}\right|}.$
Thus, regarding the issue AWC), we have the following fact (we use the term fact, rather than theorem, because the asymptotic well-conditioning has not a rigorous definition, since it is based on a vague term like ”large values”).
Fact 8.
Assume $\Lambda_{1}$ simple single real. The problem (2) is asymptotically well-conditioned if and only if $\left|\widehat{w}^{(1)}\widehat{y}_{0}\right|$ is not small.
Concerning the issue ONS), we have the following theorem. Here, $\widehat{t}$ is the time used as time unit. If $r_{1}\neq 0$, one possible choice for the time unit is the characteristic time $\frac{1}{|r_{1}|}$.
Theorem 9.
Assume $\Lambda_{1}$ simple single real and a $p$-norm as vector norm. For any $\overline{\epsilon}>0$, we have $K(t,y_{0})\approx K_{\infty}(y_{0})$ with precision $\overline{\epsilon}$ if
$\frac{t}{\widehat{t}}\geq\max\limits_{j\in\{2,\ldots,q\}}\frac{1}{\left(r_{1}-r_{j}\right)\widehat{t}}\left(\log 2+\log\frac{2+\overline{\epsilon}}{\overline{\epsilon}}+\log(q-1)\right$.
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\left.+\log\frac{f_{j}}{f_{1}}+\max\left\{0,\log\frac{\left|\widehat{w}^{(j)}\widehat{y}_{0}\right|}{\left|\widehat{w}^{(1)}\widehat{y}_{0}\right|}\right\}\right).$
Proof.
By Theorem 7, we have, when the vector norm is a $p$-norm,
$\max\{\epsilon(t,\widehat{y}_{0}),\epsilon(t)\}$ $\leq$ $2\sum\limits_{j=2}^{q}\mathrm{e}^{\left(r_{j}-r_{1}\right)t}\frac{f_{j}}{f_{1}}\max\left\{1,\frac{\left|\widehat{w}^{(j)}\widehat{y}_{0}\right|}{\left|\widehat{w}^{(1)}\widehat{y}_{0}\right|}\right\}$
$\leq$ $2(q-1)\max\limits_{j\in\{2,\ldots,q\}}\mathrm{e}^{\left(r_{j}-r_{1}\right)t}\frac{f_{j}}{f_{1}}\max\left\{1,\frac{\left|\widehat{w}^{(j)}\widehat{y}_{0}\right|}{\left|\widehat{w}^{(1)}\widehat{y}_{0}\right|}\right\}.$
Therefore, for $\epsilon>0$, we have $\max\{\epsilon(t,\widehat{y}_{0}),\epsilon(t)\}\leq\epsilon$ if
$\frac{t}{\widehat{t}}\geq\max\limits_{j\in\{2,\ldots,q\}}\frac{1}{\left(r_{1}-r_{j}\right)\widehat{t}}\left(\log\frac{2(q-1)}{\epsilon}+\log\frac{f_{j}}{f_{1}}+\max\left\{0,\log\frac{\left|\widehat{w}^{(j)}\widehat{y}_{0}\right|}{\left|\widehat{w}^{(1)}\widehat{y}_{0}\right|}\right\}\right).$
Now, given $\overline{\epsilon}>0$, consider $\epsilon>0$ such that
[…]
and use (24). ∎
Therefore, the smaller the right-hand side in (25), the earlier the onset of the asymptotic behavior.
5.2. The case $\Lambda_{1}$ simple single complex
Suppose that $\Lambda_{1}$ is simple single complex and the vector norm is the Euclidean norm.
We have
$K_{\infty}(t,y_{0})=\mathrm{OSF}(y_{0})\cdot\mathrm{OT}(t,y_{0}),$
where the oscillation scale factor $\mathrm{OSF}(y_{0})$ is given by
$\mathrm{OSF}(y_{0})=\frac{1}{\left|\widehat{w}^{(1)}\widehat{y}_{0}\right|}.$
If $V_{1}$ is not close to $1$, then the values of $K_{\infty}(t,y_{0})$ as $t$ varies have the same order of magnitude as $\mathrm{OSF}(y_{0})$, since the values of $\mathrm{OT}(t,y_{0})$ as $t$ varies have the order of magnitude $1$ (see the observation immediately after Theorem 6). Thus, regarding the issue AWC), we have the following fact
Fact 10.
Assume $\Lambda_{1}$ simple single complex and the Euclidean norms as vector norm. If $V_{1}$ is not close to $1$, then the problem (2) is asymptotically well-conditioned if and only if $\left|\widehat{w}^{(1)}\widehat{y}_{0}\right|$ is not small.
If $V_{1}$ is close to $1$, then $\left|\widehat{w}^{(1)}\widehat{y}_{0}\right|$ does not determine the asymptotic well-conditioning of the problem (2), unlike when $V_{1}$ is not close to $1$, since it is no longer guaranteed that the values of $\mathrm{OT}(t,y_{0})$ as $t$ varies maintain the order of magnitude of $1$.
Indeed, if $V_{1}$ is close to $1$, $W_{1}$ is not close to $1$ and $y_{0}$ is the second right singular vector of the matrix
$R_{1}=\left[\begin{array}[]{c}\mathrm{Re}\left(\widehat{w}^{(1)}\right)\\ \mathrm{Im}\left(\widehat{w}^{(1)}\right)\end{array}\right],$
then $\left|\widehat{w}^{(1)}\widehat{y}_{0}\right|$ is not small, but $K_{\infty}(t,y_{0})$ assumes large values as $t$ varies, i.e., the problem (2) is not asymptotically well-conditioned. In this situation, we have an oscillating term assuming large values along with a nonlarge oscillation scale factor. See [19].
However, by recalling Subsection 3.4.3, observe that $V_{1}$ is rarely close to $1$.
Concerning the issue ONS), we have the following theorem.
Theorem 11.
Assume $\Lambda_{1}$ simple single complex and the Euclidean norm as vector norm. For $\overline{\epsilon}>0$, we have $K(t,y_{0})\approx K_{\infty}(t,y_{0})$ with precision $\overline{\epsilon}$ if
$\frac{t}{\widehat{t}}\geq\max\limits_{j\in\{2,\ldots,q\}}\frac{1}{\left(r_{1}-r_{j}\right)\widehat{t}}\left(\log 2+\log\frac{2+\overline{\epsilon}}{\overline{\epsilon}}+\log(q-1)\right$.
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\left.+\frac{1}{2}\log\frac{1}{1-V_{1}}+\log\frac{f_{j}}{f_{1}}+\max\left\{0,\log\frac{\left|\widehat{w}^{(j)}\widehat{y}_{0}\right|}{\left|\widehat{w}^{(1)}\widehat{y}_{0}\right|}\right\}\right).$
Proof.
By Theorem 7, we have
$\max\{\epsilon(t,\widehat{y}_{0}),\epsilon(t)\}\leq\frac{2}{\sqrt{1-V_{1}}}\sum\limits_{j=2}^{q}\mathrm{e}^{\left(r_{j}-r_{1}\right)t}\frac{f_{j}}{f_{1}}\max\left\{1,\frac{\left|\widehat{w}^{(j)}\widehat{y}_{0}\right|}{\left|\widehat{w}^{(1)}\widehat{y}_{0}\right|}\right\}.$
The term $\frac{2}{\sqrt{1-V_{1}}}$ is obtained by comparing all the upper bounds for $G_{j}(t,u)$ and $G_{j}(t)$, $j\in\{2,\ldots,q\}$, provided in 3) and 4) of Theorem 7 and by noting that $a(V_{1},W_{1})\geq 1-V_{1}$. Now, the proof proceeds as in Theorem 9. ∎
Therefore, the smaller the right-hand side in (26), the earlier the onset of the asymptotic behavior. By comparing (25) and (26), we observe the presence of the term $\frac{1}{2}\log\frac{1}{1-V_{1}}$ in (26). If $V_{1}$ is close to 11, we could observe a delayed onset of the asymptotic behavior, compared to the case where $V_{1}$ is not close to $1$ or to the case where $\Lambda_{1}$ is simple single real.
5.3. Non-normal matrices
Suppose that the vector norm is the Euclidean norm. For a normal matrix $A$, we have $f_{j}=1,\ j\in\{1,\ldots,q\}$, and, in the case of $\Lambda_{1}$ simple single complex, $V_{1}=0$. On the other hand, a non-normal matrix $A$ can exhibit $f_{j}$ values that are arbitrarily large, as well as $V_{1}$ values that are arbitrarily close to $1$.
It is important to understand the impact of the non-normality of the matrix $A$ on the two aforementioned issues AWC) and ONS).
Regarding AWC), we can observe that the non-normality of $A$ has an impact only if $\Lambda_{1}$ is simple single complex. In fact, if $\Lambda_{1}$ is simple single real, or $\Lambda_{1}$ is simple single complex and $V_{1}$ is not close to $1$, then the problem (2) is asymptotically well-conditioned if and only if $|\widehat{w}^{(1)}\widehat{y}_{0}|$ is not small, where the quantity $|\widehat{w}^{(1)}\widehat{y}_{0}|$ is unrelated to the non-normality. On the other hand, if $\Lambda_{1}$ is simple single complex and $V_{1}$ is close to $1$, a case possible only for $A$ non-normal, the quantity $|\widehat{w}^{(1)}\widehat{y}_{0}|$ does not determine the asymptotic well-conditioning of the problem (2), as we have seen in the previous Subsection 5.2. In other words, for a non-normal matrix $A$, we could have an oscillating term assuming large values along with a nonlarge oscillation scale factor.
Regarding ONS), we can observe that if some ratio $\frac{f_{j}}{f_{1}}$, $j\in\{2,\ldots,q\}$, is large, or $\Lambda_{1}$ is simple single complex and $V_{1}$ is close to 1, cases possible only for $A$ non-normal, the onset of the asymptotic behavior could be delayed: recall (25) and (26).
It’s worth noting that the ratios $\frac{f_{j}}{f_{1}}$, $j\in\{2,\ldots,q\}$, can be much smaller than the large values of $f_{j}$ characterizing a high non-normality of the matrix $A$. More details are given in the next example.
Example 12.
In Figure 12, for the Euclidean norm as vector norm, we see the maxima
$M=\max\limits_{j\in\{2,\ldots,n\}}\log f_{j}\text{\ \ and\ \ }\widehat{M}=\max\limits_{j\in\{2,\ldots,n\}}\log\frac{f_{j}}{f_{1}},$ (27)
computed for 10 00010\ 000 matrices $A$ of order $n=500$, whose elements are sampled from the standard normal distribution. It appears that the values $\widehat{M}$ are smaller than the values $M$. The following table provides some statistical details:
maximum value minimum value median
$M$ $10.06$ $3.41$ $4.35$
$\widehat{M}$ $8.00$ $0.00$ $2.95$
Figure 12. Maxima $M$ and $\widehat{M}$ in (27) for 10 00010\ 000 random instances of $A$ of order $500$.
Moreover, the highly non-normal example in Subsection 6.5 also highlights that the ratios $\frac{f_{j}}{f_{1}}$ can be significantly smaller than the values of $f_{j}$.
5.4. Normal matrices
Once again, assume that the vector norm is the Euclidean norm and consider $A$ normal.
In the case $\Lambda_{1}$ simple single complex, we have $V_{1}=W_{1}=0$. Consequently, we have the constant over time values
$\mathrm{OT}(t,y_{0})=\sqrt{\frac{1}{2}}\text{\ \ and\ }K_{\infty}(t,y_{0})=\frac{1}{|\widehat{w}^{(1)}\widehat{y}_{0}|}\sqrt{\frac{1}{2}}.$
The asymptotic condition number $K_{\infty}(t,y_{0})$ is constant in time, as in the case $\Lambda_{1}$ simple single real.
Moreover, since $\frac{f_{j}}{f_{1}}=1$ for $j\in\{2,\ldots,q\}$, and $V_{1}=0$ in the case $\Lambda_{1}$ simple single complex, an earlier onset of the asymptotic behavior for $K(t,y_{0})$ is expected compared to non-normal matrices: see (25) and (26)).
Finally, the constant in time asymptotic condition number $K_{\infty}(t,y_{0})$ is not only the limit as $t\rightarrow+\infty$ of $K(t,y_{0})$, but it is also the supremum of $K(t,y_{0})$ for $t\geq 0$ (see [17]).
6. Examples
To illustrate the contents of the present paper, we consider five examples of ODEs (1), where the first two are the models of Section 1 revisited.
6.1. Gross Domestic Product and National Debt
The first example is the ODE (3) of the GDP-ND model of Subsection 1.2.1. Assume, as in the instance (4), that there are two real eigenvalues. Moreover, since we set $Q(0)=1$, there is uncertainty only on $B(0)$ and therefore the direction of perturbation is $\widehat{z}_{0}=(0,\pm 1)$ (in any vector $p$-norm). In the Euclidean norm, we have
$K_{\infty}(y_{0},\widehat{z}_{0})=\frac{\left|\widehat{w}^{(1)}\widehat{z}_{0}\right|}{\left|\widehat{w}^{(1)}\widehat{y}_{0}\right|}=\sqrt{1+B(0)^{2}}\cdot\frac{1}{\left|\frac{w^{(1)}_{1}}{w^{(1)}_{2}}+B(0)\right|}.$
In the instance (4), we have $\frac{w^{(1)}_{1}}{w^{(1)}_{2}}=-1$ and then
$K_{\infty}(y_{0},\widehat{z}_{0})=\sqrt{1+B(0)^{2}}\cdot\frac{1}{|-1+B(0)|}.$ (28)
For the initial ND $B(0)=0.61$ considered in Subsection 1.2.1, we have $K_{\infty}(y_{0},\widehat{z}_{0})=3.0035$.
In Figure 13, which extends Figure 5 to the interval $[0,150]$ (the characteristic time is $20$), we can observe how the relative error $\delta(t)$ asymptotically approaches the value $0.025641$. By recalling (16), we have
$\delta(t)=K(t,y_{0},\widehat{z}_{0})\cdot\delta(0)\text{\ \ and\ \ }\lim\limits_{t\rightarrow+\infty}\delta(t)=K_{\infty}(y_{0},\widehat{z}_{0})\cdot\delta(0)=3.0035\cdot\delta(0).$
The expression for $K_{\infty}(y_{0},\widehat{z}_{0})$ in (28) has a singularity at $B(0)=1$. In this case, where the initial value $(Q(0),B(0))=(1,1)$ has zero component along the eigenvector corresponding to the rightmost eigenvalue, the asymptotic condition number $K_{\infty}(t,y_{0},\widehat{z}_{0})$ grows exponentially in time (see [18]).
For $(Q(0),B(0))=(1,1)$ and $(\widetilde{Q}(0),\widetilde{B}(0))=(1,0.99)$, Figures 14, 15, and 16 depict the same information as Figures 3, 4, and 5 do for $(Q(0),B(0))=(1,0.61)$ and $(\widetilde{Q}(0),\widetilde{B}(0))=(1,0.60)$. The differences between the two cases are quite evident. After 50yr50\ $\mathrm${yr}, the relative error is about $12$ times the initial relative error.
Figure 13. Relative error $\delta(t)$ for the GDP-ND model.
Figure 14. $B(0)=1$ and $\widetilde{B}(0)=0.99$. Left: points $(Q(0),B(0))$ and $(\widetilde{Q}(0),\widetilde{B}(0))$. Right: points $(Q(50),B(50))$ and $(\widetilde{Q}(50),\widetilde{B}(50))$.
Figure 15. $B(0)=1$ and $\widetilde{B}(0)=0.99$. GDP $Q(t)$, perturbed GDP $\widetilde{Q}(t)$, ND $B(t)$ and perturbed ND $\widetilde{B}(t)$.
Figure 16. $B(0)=1$ and $\widetilde{B}(0)=0.99$. Relative error $\delta(t)$ for the GDP-ND model.
6.2. Building heating
The second example is the ODE (9) of the building heating model of Subsection 1.2.1. Assume, as in the instance (10), that there are three (negative) eigenvalues. We have
$K_{\infty}(y_{0},\widehat{z}_{0})=\frac{\left|\widehat{w}^{(1)}\widehat{z}_{0}\right|}{\left|\widehat{w}^{(1)}\widehat{y}_{0}\right|}\text{\ and\ }K_{\infty}(y_{0})=\frac{1}{\left|\widehat{w}^{(1)}\widehat{y}_{0}\right|}.$ (29)
For the particular instance (10) and the Euclidean norm as the vector norm, we have the normalized left eigenvector
$\widehat{w}^{(1)}=\left[\begin{array}[]{c}-0.4462\ -0.6111\ -0.6538\end{array}\right].$ (30)
corresponding to the rightmost eigenvalue. For the initial value and perturbed initial value in (11) and (12), we have $K_{\infty}(y_{0},\widehat{z}_{0})$ =11.8648=11.8648.
In Figure 7, we observe the relative error $\delta(t)=K(t,y_{0},\widehat{z}_{0})\cdot\delta(0)$ over approximately two characteristic times (the characteristic time is 3.17h3.17\ $\mathrm${h}). The relative error asymptotically approaches
$\lim\limits_{t\rightarrow+\infty}\delta(t)=K_{\infty}(y_{0},\widehat{z}_{0})\cdot\delta(0)=11.8648\cdot\delta(0)$
as $t$ increases. Moreover, we have $K_{\infty}(y_{0})=12.1330$. This indicates that the perturbation in (12) is nearly the worst-case scenario.
The expression for $K_{\infty}(y_{0})$ in (29) has a singularity when $y_{0}$ satisfies $\widehat{w}^{(1)}y_{0}=0$. For the instance (10), this occurs when (11) is replaced with
$y_{0}=\left(3.5\ ^{\circ}\mathrm{C},-5.2298\ ^{\circ}\mathrm{C},2.5\ ^{\circ}\mathrm{C}\right).$ (31)
In this case, where $y_{0}$ has zero component along the eigenvector corresponding to the rightmost eigenvalue, the asymptotic condition number $K_{\infty}(t,y_{0})$ grows exponentially in time (see [18]).
In Figure 17, for $y_{0}$ given in (31) and
$\widetilde{y}_{0}=\left(4\ ^{\circ}\mathrm{C},-5\ ^{\circ}\mathrm{C},3\ ^{\circ}\mathrm{C}\right),$
we observe the relative error $\delta(t)$ for $t\in[0,6\ \mathrm{h}]$. The difference compared to Figure 7 is quite evident. After 6h6\ $\mathrm${h}, the relative error is 235 times the initial relative error.
Figure 17. $\widehat{w}^{(1)}y_{0}=0$. Relative error $\delta(t)$ for the heating building model.
Remark 13.
In the instance (10), all the components of $\widehat{w}^{(1)}$ in (30) have the same sign. This fact permits the following observation. If all the components of $y_{0}$ have the same sign, then
$K_{\infty}(y_{0})=\frac{\|y_{0}\|_{2}}{\left|\widehat{w}^{(1)}_{1}\right||y_{01}|+\left|\widehat{w}^{(1)}_{2}\right||y_{02}|+\left|\widehat{w}^{(1)}_{3}\right||y_{03}|}$
and then
$K_{\infty}(y_{0})\leq\frac{1}{\min\left\{\left|\widehat{w}^{(1)}_{1}\right|,\left|\widehat{w}^{(1)}_{2}\right|,\left|\widehat{w}^{(1)}_{3}\right|\right\}}=\frac{1}{0.4462}=2.2411.$
This shows that if the temperatures of the three floors are all larger or all smaller than the equilibrium temperatures, then the relative error of the perturbed solution is not much larger than the initial relative error. This is not the case for the initial values (11) and (31), where the relative error grows significantly compared to its initial value.
Remark 14.
In the building heating model, as well as in the GDP-ND model, we have considered $y_{0}$ as the actual initial value and $\widetilde{y}_{0}$ as the initial value available to us for the simulation. Therefore, we have considered $\delta(t)$ as the relative error of the simulated solution with respect to the actual solution.
However, to obtain a computable asymptotic condition number $K_{\infty}(y_{0})$ for the heating building model, or a computable asymptotic condition number $K_{\infty}(y_{0},\widehat{z}_{0})$ for the GDP-ND model, one may consider $y_{0}$ as the available initial value and $\widetilde{y}_{0}$ as the actual initial value. Therefore, now we are considering $\delta(t)$ as the relative error of the actual solution with respect to the simulated solution.
In the case of the building heating model, we have $K_{\infty}(y_{0})=4.9195$ by considering the available initial value (12) as $y_{0}$, instead of $K_{\infty}(y_{0})=12.1330$ by considering the actual initial value (11) as $y_{0}$.
In the case of the GDP-ND model, we have $K_{\infty}(y_{0},\widehat{z}_{0})=2.29$ for $B(0)=0.60$, the available initial value, instead of $K_{\infty}(y_{0},\widehat{z}_{0})=3.00$ for $B(0)=0.61$, the actual initial value.
6.3. Another building heating model
The paper [23] presents a building heating model of dimension $4$, where the first state component is the indoor temperature and the other three are the temperatures in three layers of the wall separating the indoor space from the outdoor space. The matrix of the system of differential equations is
$A=\left[\begin{array}[]{rrrr}-5.7215&5.7215&0&0\\ 0.23076&-0.39276&0.162&0\\ 0&0.081&-0.162&0.081\\ 0&0&0.162&-0.91116\end{array}\right]$ (32)
[…]
$w^{(1)}=\left[-0.029943\ -0.73735\ -1.1058\ -0.1027\right]$
Consider initial temperatures $y_{0}$ (remind that these are temperatures with respect to the equilibrium temperatures) with $y_{02}=y_{03}=y_{04}=0$. We have, for a vector $p$-norm,
$K_{\infty}(y_{0})=\frac{1}{|\widehat{w}^{(1)}\widehat{y}_{0}|}=\frac{\|w^{(1)}\|}{|w^{(1)}_{1}|}$.
In Figure 18, for the $\infty$-norm as the vector norm, we see the condition numbers $K(t,y_{0})$ and $K_{\infty}(y_{0})=65.987$ for $t\in[0,30\ \mathrm{h}]$. The asymptotic behavior of the condition number is achieved within a characteristic time.
Figure 18. Condition numbers $K(t,y_{0})$ and $K_{\infty}(y_{0})$ for the matrix (32) and $y_{0}$ with $y_{02}=y_{03}=y_{04}=0$. The vector norm is the $\infty$-norm.
6.4. Charged particle subject to a magnetic fields with viscous frictional force
The first three examples described situations in which the matrix $A$ in (1) had real eigenvalues. This fourth example examines a situation in which the rightmost eigenvalues of $A$ form a complex conjugate pair.
We consider a charged particle subject to a magnetic field and moving in a medium with an anisotropic viscous frictional force (see the paper [6], which also considers Brownian motion and an electric field). Let $m$ and $q$ be the mass and the charge, respectively, of the particle, let $B=(B_{x},B_{y},B_{z})$ be the constant magnetic field and let $\Gamma_{x}$, $\Gamma_{y}$ and $\Gamma_{z}$ be the constant coefficients of the viscous frictional force in the three spatial directions. The motion of the particle is described by the first order linear ODE for its velocity $v=(v_{x},v_{y},v_{z})$:
$v^{\prime}(t)=Av(t),$
where
$A=\frac{1}{m}\left[\begin{array}[]{rrr}-\Gamma_{x}&qB_{z}&-qB_{y}\\ -qB_{z}&-\Gamma_{y}&qB_{x}\\ qB_{y}&-qB_{x}&-\Gamma_{z}\end{array}\right].$
We examine the instance (comparable with the examples given in [6]) where, in suitable units,
$m=1$
$qB_{x}=-0.5,\ qB_{y}=-1.3,\ qB_{z}=-0.4,$
$\Gamma_{x}=0.4,\ \Gamma_{y}=0.8,\ \Gamma_{z}=0.2.$
In such an instance, the matrix $A$ has the eigenvalues
$\lambda_{1}=-0.3433+\sqrt{-1}\cdot 1.4326,\ \overline{\lambda_{1}}=-0.3433-\sqrt{-1}\cdot 1.4326,\ \lambda_{2}=-0.7133,$
with a complex conjugate pair $\lambda_{1}$ and $\overline{\lambda_{1}}$ as rightmost eigenvalues.
In the Euclidean norm, the normalized left eigenvector corresponding to $\lambda_{1}$ is
$\widehat{w}^{(1)}=\left[\begin{array}[]{rrr}0.6543+\sqrt{-1}\cdot 0.0111&-0.2821+\sqrt{-1}\cdot 0.1288&-0.1031-\sqrt{-1}\cdot 0.6819\end{array}\right].$
The oscillation scale factor of the asymptotic condition number $K_{\infty}(t,v_{0})$ is
$\mathrm{OSF}(v_{0})=\frac{1}{\left|\widehat{w}^{(1)}\widehat{v}_{0}\right|}=\frac{\|v_{0}\|_{2}}{\sqrt{\left(\mathrm{Re}(\widehat{w}^{(1)})v_{0}\right)^{2}+\left(\mathrm{Im}(\widehat{w}^{(1)})v_{0}\right)^{2}}},$
where $v_{0}$ is the initial velocity.
In Figure 19, we see, for $v_{0}=(0.5,1,0.5)$, over $6$ characteristic times $\frac{1}{-r_{1}}=2.9125$: the condition numbers $K(t,v_{0})$ and $K_{\infty}(t,v_{0})$, the oscillation scale factor $\mathrm{OSF}(v_{0})=5.92$ and the lower and upper bounds for $K_{\infty}(t,y_{0})$ derived by the lower and upper bounds (20)-(21) for the oscillating term. The asymptotic behavior is achieved in approximately three characteristic times.
Since $V_{1}=0.0587$ and $W_{1}=0.0937$ in this example, the oscillations of $K_{\infty}(t,v_{0})$ are tight. In fact, the oscillating term $\mathrm{OT}(t,v_{0})$ varies only slightly around $\sqrt{\frac{1}{2}}$, the constant value of $\mathrm{OT}(t,v_{0})$ for $V_{1}=W_{1}=0$.
Figure 19. Condition numbers $K(t,v_{0})$ and $K_{\infty}(t,v_{0})$, $t\in[0,6\text{ characteristic time}]$, oscillating scale factor $\mathrm{OSF}(v_{0})$ and lower and upper bounds for $K_{\infty}(t,v_{0})$. The vector norm is the Euclidean norm.
6.5. A highly non-normal matrix AA
In this final example, we consider in the ODE (1) the highly non-normal matrix $A=VDV^{-1}$, where $V$ is the Hilbert matrix of order $8$ and
$D=\mathrm{diag}(-0.1,-0.2,\ldots,-0.8).$
The Euclidean norm is used as the vector norm.
The high non-normality of the matrix $A$ appears in the large values $f_{j}$ defined in (23): we have
$f_{1}=5.2554\cdot 10^{5},\ f_{2}=1.677\cdot 10^{7},\ f_{3}=1.6347\cdot 10^{8},\ f_{4}=7.1819\cdot 10^{8},$
$f_{5}=1.6407\cdot 10^{9},\ f_{6}=2.0252\cdot 10^{9},\ f_{7}=1.2815\cdot 10^{9},\ f_{8}=3.2603\cdot 10^{8}.$
In Figure 20, we see the norm $\|\mathrm{e}^{tA}\|_{2}$ for $t\in[0,100]$, i.e., for $t$ over ten characteristic times $\frac{1}{-r_{1}}=10$. Note that $\|\mathrm{e}^{tA}\|_{2}$ is the absolute condition number of the problem (2), i.e., when we consider absolute errors instead of relative errors. Despite $\|\mathrm{e}^{tA}\|_{2}\rightarrow 0$, $t\rightarrow+\infty$, this fact holds little significance. This is because, owing to the high non-normality of the matrix $A$, the norm remains large for numerous characteristic times and it only drops below $1$ after 13 characteristic times.
Hence, it is of little interest to ascertain that the absolute error of a perturbation in the initial value decays to zero in the long-time, as this “long-time” may be very distant. What we observe within a possible time span of interest is a very large amplification of the absolute error of the perturbation.
Figure 20. Norm $\|\mathrm{e}^{tA}\|_{2}$ for $A=VDV^{-1}$, where $V$ is the Hilbert matrix of order 88.
On the other hand, in Figure 21 we see, for two different initial values $y_{0}$ and for $t\in[0,100]$, the relative condition number $K(t,y_{0})$ of the problem (2), i.e., when we consider relative errors instead of absolute errors, and its asymptotic value $K_{\infty}(y_{0})=\frac{1}{|\widehat{w}^{(1)}\widehat{y}_{0}|}.$ Similar pictures are obtained by randomly varying the initial value.
(a) $y_{0}=(1,1,1,1,1,1,1,1)$
(b) $y_{0}=(1,1,1,1,-1,-1,-1,-1)$
Figure 21. Condition numbers $K(t,y_{0})$ and $K_{\infty}(y_{0})$ for $A=VDV^{-1}$, where $V$ is the Hilbert matrix of order 88. The vector norm is the Euclidean norm.
Note that the high non-normality does not have a critical impact, as is the case with the absolute condition number $\|\mathrm{e}^{tA}\|_{2}$. In fact, we can observe the following.
  • 1)
    Although $K(t,y_{0})$ oscillates similarly to $\|\mathrm{e}^{tA}\|_{2}$ as $t$ varies, the maximum of $K(t,y_{0})$ as $t$ varies and $K_{\infty}(y_{0})$ have the same order of magnitude. The high value of $K_{\infty}(y_{0})$ in the top figure is due to a small value of $|\widehat{w}^{(1)}\widehat{y}_{0}|$, rather than high non-normality.
  • 2)
    The constant asymptotic behavior is achieved after a few (approximately five) characteristic times, similar to many examples where the matrix is not highly non-normal. For comparison, consider the values of $\|\mathrm{e}^{tA}\|_{2}$ in Figure 20 after five characteristic times. Indeed, the fact that the high non-normality does not critically impact the attainment of asymptotic behavior can be explained by observing that, when the characteristic time $\frac{1}{-r_{1}}$ is used as the time unit $\widehat{t}$, we have
    $\max\limits_{j\in\{2,\ldots,8\}}\frac{1}{(r_{1}-r_{j})\widehat{t}}\log\frac{f_{j}}{f_{1}}=\max\limits_{j\in\{2,\ldots,8\}}\frac{1}{\underset{=j}{\underbrace{\frac{-r_{j}}{-r_{1}}}}-1}\log\frac{f_{j}}{f_{1}}=3.4629$
    in (25). Therefore, the high non-normality affects the achievement of the asymptotic behavior for only $3.4629$ characteristic times.
Whereas the fact that the absolute condition number $\|\mathrm{e}^{tA}\|_{2}$ decays to zero in the long-time is not significant, the fact that the relative condition number $K(t,y_{0})$ asymptotically becomes $K_{\infty}(y_{0})$, and its order of magnitude never exceeds that of $K_{\infty}(y_{0})$, is highly significant.
7. Asymptotic and non-asymptotic conditioning
In this paper, we have studied the asymptotic conditioning of the problem (2), i.e. the conditioning for large $t$. On the other hand, we are also interested in the conditioning of (2) for any $t$, not only for large $t$.
Of course, if the problem (2) is asymptotically ill-conditioned, i.e $K_{\infty}(t,y_{0})$ assumes large values as $t$ varies, then the problem (2) is also ill-conditioned, i.e. $K(t,y_{0})$ assumes large values as $t$ varies.
Vice versa, of course, we cannot say that if the problem (2) is asymptotically well-conditioned, i.e., $K_{\infty}(t,y_{0})$ does not assume large values as $t$ varies, then the problem (2) is also well-conditioned, i.e., $K(t,y_{0})$ does not assume large values as $t$ varies? In fact, the asymptotic behavior $K_{\infty}(t,y_{0})$ of $K(t,y_{0})$ cannot take into account an initial increase of $K(t,y_{0})$ to values much larger than those of $K_{\infty}(t,y_{0})$.
However, we now show experimentally that such an initial increase in $K(t,y_{0})$ is very rare, and then the asymptotic well conditioning can indeed be considered equivalent to the well conditioning, for the problem (2). This stresses the crucial and central importance of the asymptotic condition number.
For $n=5,25,100$, we consider 10 00010\ 000 instances of an ODE (1) of dimension $n$, where the entries of $A$ and the components of the initial value $y_{0}$ are sampled from the standard normal distribution. Of these 10 00010\ 000 instances, $5000$ have a matrix $A$ with a rightmost real eigenvalue and $5000$ have a matrix $A$ with a rightmost pair of complex conjugate eigenvalues. Since these are random instances, the rightmost eigenvalues are always simple.
We compute the ratios
$R=\frac{\max\limits_{t\in[0,T]}K(t,y_{0})}{\max\limits_{t\in[0,T]}K_{\infty}(t,y_{0})}$ (33)
for these 10 00010\ 000 instances, where $T$ is fifty times the characteristic time $\widehat{t}=\frac{1}{|r_{1}|}$. The maxima are determined using $1000$ equally spaced sampling points over the interval $[0,T]$, i.e. with a stepsize $\frac{1}{20}$ of the characteristic time. Observe that, for the case of a real rightmost eigenvalue, we have
$R=\frac{\max\limits_{t\in[0,T]}K(t,y_{0})}{K_{\infty}(y_{0})},$
since $K_{\infty}(t,y_{0})=K_{\infty}(y_{0})$ remains constant over time $t$.
In the case of the Euclidean norm as the vector norm, we see in Figures 22 and 23 the ratio $R$ for the $5000$ instances with a rightmost real eigenvalue and for the $5000$ instances with a rightmost pair of complex conjugate eigenvalues, respectively.
Figure 22. Ratio $R$ in (33) for $5000$ instances of $A$ with a rightmost real eigenvalue for $n=5,25,100$. The vector norm is the Euclidean norm.
Figure 23. Ratio $R$ in (33) for $5000$ instances of $A$ with a rightmost pair of complex conjugate eigenvalues for $n=5,25,100$. The vector norm is the Euclidean norm.
The next tables give statistical details for the ratio $R$: the first for the 5000 instances with the rightmost real eigenvalue, and the second for the 5000 instances with the rightmost pair of complex conjugate eigenvalues.
maximum value minimum value first decile median ninth decile
$n=5$ 5.4521 0.04036 1 1.0182 1.2906
$n=25$ 4.0616 0.00073135 1 1.039 1.3539
$n=100$ 3.8299 0.0027379 0.91861 1.0354 1.3156
maximum value minimum value first decile median ninth decile
$n=5$ 4.3163 0.30049 1 1.0039 1.1754
$n=25$ 4.8062 0.19181 0.99997 1.0175 1.2896
$n=100$ 5.9408 0.081863 0.8633 1.0175 1.3157
In the case of the $\infty$-norm and the $1$-norm as vector norms, we do not present figures but instead report the statistical details for $R$ in two similar tables as above.
$\infty$-norm:
maximum value minimum value first decile median ninth decile
$n=5$ 7.4913 0.034023 1 1.0777 1.5458
$n=25$ 5.3264 0.00088701 1 1.1189 1.6008
$n=100$ 5.4355 0.0022694 0.95464 1.1119 1.5304
maximum value minimum value first decile median ninth decile
$n=5$ 3.9183 0.28906 1 1.0001 1.2071
$n=25$ 4.5211 0.14 0.99848 1.0253 1.3403
$n=100$ 7.3818 0.065372 0.8556 1.0291 1.3511
11-norm:
maximum value minimum value first decile median ninth decile
$n=5$ 7.2499 0.052121 1 1.0534 1.4814
$n=25$ 4.1774 0.00067403 1 1.065 1.4568
$n=100$ 4.0398 0.0028903 0.93424 1.0521 1.408
maximum value minimum value first decile median ninth decile
$n=5$ 3.3708 0.32165 1 1.0019 1.2153
$n=25$ 4.7886 0.1652 0.99979 1.0238 1.3232
$n=100$ 6.1563 0.073164 0.87166 1.0251 1.3543
From these experiments, we can conclude that, at least for the $p$-norm with $p\in\{1,2,\infty\}$ as vector norm, the ratio $R$ is very rarely much larger than $1$. Indeed, it is never much larger than $1$ in these 10 00010\ 000 instances.
Therefore, we have the strong evidence of the following fact.
Fact 15.
Suppose that the $p$-norm with $p\in\{1,2,\infty\}$ is used as vector norm. In the vast majority of cases,
$\max\limits_{t\geq 0}K(t,y_{0})\ \text{and}\ \max\limits_{t\geq 0}K_{\infty}(t,y_{0})$
are of the same order of magnitude. Consequently, the problem (2) is well-conditioned if and only if it is asymptotically well-conditioned.
In the following two facts, we separate the cases of a rightmost real eigenvalue and a rightmost pair of complex conjugate eigenvalues. For the latter, recall the observation after Theorem 6.
Fact 16.
Suppose that the $p$-norm with $p\in\{1,2,\infty\}$ is used as vector norm. In the vast majority of cases of a rightmost real eigenvalue,
$\max\limits_{t\geq 0}K(t,y_{0})\ \text{and}\ K_{\infty}(y_{0})=\frac{1}{|\widehat{w}^{(1)}\widehat{y}_{0}|}=\frac{\left\|w^{(1)}\right\|\left\|y_{0}\right\|}{|w^{(1)}y_{0}|}$ (40)
are of the same order of magnitude. Consequently, the problem (2) is well-conditioned if and only if $|\widehat{w}^{1}\widehat{y}_{0}|$ is not small.
Fact 17.
Suppose that the Euclidean norm is used as vector norm. In the vast majority of cases of a rightmost complex conjugate pair of eigenvalues,
$\max\limits_{t\geq 0}K(t,y_{0})\ \text{and}\ \max\limits_{t\geq 0}K_{\infty}(t,y_{0})\text{\ \ and\ \ }\mathrm{OSF}(y_{0})=\frac{1}{|\widehat{w}^{(1)}\widehat{y}_{0}|}=\frac{\left\|w^{(1)}\right\|_{2}\left\|y_{0}\right\|_{2}}{|w^{(1)}y_{0}|}$
are of the same order of magnitude. Consequently, the problem (2) is well-conditioned if and only if $|\widehat{w}^{1}\widehat{y}_{0}|$ is not small.
7.1. Highly non-normal matrices
One might get the impression that the previous conclusion regarding what happens in the vast majority of cases does not hold for highly non-normal matrices, since sampling the entries of the matrix $A$ from the standard normal distribution rarely produces highly non-normal matrices. However, this is not the case, as demonstrated by the example of a highly non-normal matrix in Subsection 6.5, where the quantities in (40) are of the same order of magnitude.
To reinforce this conclusion, we consider, for $n=5,25,100$, a total of $5000$ instances of an ODE (1) of dimension $n$, where the initial value $y_{0}$ has components drawn from the standard normal distribution, and the matrix $A$ is given by $A=QUQ^{T}$ with $Q$ being the orthogonal matrix obtained from the QR factorization of a $n\times n$ random matrix with entries from the standard normal distribution; and $U$ is a $n\times n$ random upper triangular matrix with entries also drawn from the standard normal distribution. Generating matrices $A$ in this manner yields highly non-normal matrices with real eigenvalues.
Figure 24 and the following tables are analogous to Figures 22 and 23 and the subsequent tables. We now observe larger values of the ratio $R$, but they are rare. Overall, the global picture remains unchanged compared to the previous experiment.
Euclidean norm:
maximum value minimum value first decile median ninth decile
$n=5$ 53.1093 0.053073 1 1.0254 1.7962
$n=25$ 72.8536 0.019093 0.99592 1.1645 3.0518
$n=100$ 414.7996 0.0029776 0.86314 1.5107 4.9485
$\infty$-norm:
maximum value minimum value first decile median ninth decile
$n=5$ 43.0974 0.065916 1 1.0974 2.1424
$n=25$ 101.0679 0.013189 0.99805 1.297 3.4043
$n=100$ 502.1649 0.0028646 0.87874 1.6603 5.287
11-norm:
maximum value minimum value first decile median ninth decile
$n=5$ 58.3547 0.051422 1 1.0719 2.0066
$n=25$ 69.126 0.020669 0.99736 1.2393 3.2585
$n=100$ 434.7223 0.0035079 0.85826 1.5892 5.092
Figure 24. Ratio $R$ in (33) for the $5000$ instances of $A=QUQ^{T}$. The vector norm is the Euclidean norm.
8. Non-normal dynamics
Suppose we want to simulate the transient phase of a real-world system, where this transient $y$ satisfies an ODE (1) with a non-normal stable matrix $A$, where stable means that all the eigenvalues of $A$ have negative real part. For a non-linear real-world system, this ODE is the linearization around an asymptotically stable equilibrium of a non-linear ODE.
Due to the non-normality of the matrix $A$, it is expected that the transient exhibits an initial (possibly large) growth before decaying to zero. This phenomenon is what characterizes non-normal dynamics and can destabilize a non-linear system subject to small perturbations from the asymptotically stable equilibrium (see [22] and [4]).
Suppose that, through simulation, we need to determine the maximum growth of the transient, i.e. $\max\limits_{t\geq 0}\|y(t)\|$. Additionally, suppose there is uncertainty in the initial value of our simulation. Here, we consider $y_{0}$ as the initial value available for the simulation and $\widetilde{y}_{0}$ as the actual initial value.
Since there are uncertainties in the initial value, it is important to assess how close the simulation result $\max\limits_{t\geq 0}\|y(t)\|$ is to the actual maximum growth of the transient, i.e. $\max\limits_{t\geq 0}\|\widetilde{y}(t)\|$.
Information on relative closeness can be obtained by the next result.
Theorem 18.
Suppose that the initial value of an ODE (1) with $A$ stable is perturbed by a normwise relative error $\varepsilon$. We have
$\left|\frac{\max\limits_{t\geq 0}\|\widetilde{y}(t)\|}{\max\limits_{t\geq 0}\|y(t)\|}-1\right|\leq E:=\max\limits_{t\geq 0}K(t,y_{0})\cdot\varepsilon,$ (44)
whenever $E<1$.
Proof.
Since
$\|\widetilde{y}(t)\|=\|y(t)\|(1+\chi(t)),$
where
$|\chi(t)|\leq\delta(t)=K(t,y_{0},\widehat{z}_{0})\varepsilon\leq K(t,y_{0})\varepsilon\leq E$
(see (16)), we have
$\max\limits_{t\geq 0}\|\widetilde{y}(t)\|\leq\max\limits_{t\geq 0}\|y(t)\|(1+E).$
Moreover, since
$\|y(t)\|=\frac{\|\widetilde{y}(t)\|}{1+\chi(t)},$
we have
$\max\limits_{t\geq 0}\|y(t)\|\leq\frac{\max\limits_{t\geq 0}\|\widetilde{y}(t)\|}{1-E},$
whenever $E<1$. ∎
The quantity $E$ is an upper bound for the relative error of the actual maximum transient growth with respect to the simulated maximum transient growth, demonstrating how the condition number $K(t,y_{0})$ introduced in the present paper can be useful in the study of non-normal dynamics.
By recalling Fact 15, we can conclude what follows.
Fact 19.
Suppose that the $p$-norm with $p\in\{1,2,\infty\}$ is used as vector norm. In the vast majority of cases, $E$ in (44) and
$E_{\infty}:=\max\limits_{t\geq 0}K_{\infty}(t,y_{0})\cdot\varepsilon$
are of the same order of magnitude.
Therefore, also the asymptotic condition number $K_{\infty}(t,y_{0})$ introduced and analyzed in this paper can be useful in the study of non-normal dynamics.
Example 20.
Consider an ODE (1) with
$A=\left[\begin{array}[]{rr}-1&a\\ 0&-2\end{array}\right],$ (45)
where $a>0$ is large. In this example, the $\infty$-norm is used as the vector norm.
The matrix $A$ represents a classical example of a highly non-normal stable matrix (see the “quiz” at the beginning of [22]), exhibiting a large initial growth of the solution of (1). In Figure 25, for $a=50,500,5000$, we see that $\|\mathrm{e}^{tA}\|_{\infty}$ initially grows significantly over time before eventually decaying to zero.
Since
$\left|\max_{t\geq 0}\|\widetilde{y}(t)\|_{\infty}-\max_{t\geq 0}\|y(t)\|_{\infty}\right|\leq\max_{t\geq 0}\|\mathrm{e}^{tA}\|_{\infty}\|\widetilde{y}_{0}-y_{0}\|_{\infty},$
it is expected that $\max\limits_{t\geq 0}\|\widetilde{y}(t)\|_{\infty}$ will not be as close to $\max\limits_{t\geq 0}\|y(t)\|_{\infty}$ as $\widetilde{y}_{0}$ is close to $y_{0}$, when the closeness is measured in terms of absolute error.
On the other hand, as we will see below, this is not the case when the closeness is measured in terms of relative error. Therefore, if controlling the relative error of $\max\limits_{t\geq 0}\|\widetilde{y}(t)\|_{\infty}$ is important, considering the growth of $\|\mathrm{e}^{tA}\|_{\infty}$ can be misleading, as it pertains to the absolute error rather than the relative error.
Figure 25. $\|\mathrm{e}^{tA}\|_{\infty}$, $t\in[0,10]$, for $a=50,500,5000$ in (45).
For $a=50,500,5000$, we present in the table below statistical details of the ratio
$R=\frac{\max\limits_{t\in[0,T]}K(t,y_{0})}{K_{\infty}(y_{0})}$
computed for 10 00010\ 000 instances of $y_{0}$ sampled from the standard normal distribution. As in Section 7, we set $T$ to fifty times the characteristic time, i.e. $T=50$, and the maximum is determined using $1000$ equally spaced sampling points.
median ninth decile 99th percentile maximum value
$a=50$ $1.2629$ $3.3095$ $16.2544$ $46.806$
$a=500$ $1.039$ $1.1803$ $9.5118$ $275.9292$
$a=5000$ $1.0039$ $1.0159$ $1.1308$ $53.393$
This table confirms that, in the vast majority of cases for $y_{0}$, the maxima in (40) are of the same order of magnitude. Consequently, it validates Fact 19.
We explain this example in more detail. The left eigenvector corresponding to the rightmost eigenvalue 1-1 is
$w^{(1)}=\left[\begin{array}[]{rr}1&a\end{array}\right].$
Therefore,
$E_{\infty}$ $=$ $\frac{1}{|\widehat{w}^{(1)}\widehat{y}_{0}|}\cdot\varepsilon=\frac{\left\|w^{(1)}\right\|_{\infty}}{|w^{(1)}y_{0}|}\cdot\|\widetilde{y}_{0}-y_{0}\|_{\infty}$
$=$ $\frac{1+a}{|y_{01}+ay_{02}|}\cdot\max\{|\widetilde{y}_{01}-y_{01}|,|\widetilde{y}_{02}-y_{02}|\}.$
For $\left|\frac{y_{01}}{y_{02}}\right|$ not large (this is true in the vast majority of cases for $y_{0}$), $E_{\infty}$ is close to
$\frac{1}{|y_{02}|}\cdot\max\{|\widetilde{y}_{01}-y_{01}|,|\widetilde{y}_{02}-y_{02}|\}=\max\left\{\left|\frac{y_{01}}{y_{02}}\right|\cdot\frac{|\widetilde{y}_{01}-y_{01}|}{|y_{01}|},\frac{|\widetilde{y}_{02}-y_{02}|}{|y_{02}|}\right\},$
since $a$ is large. Thus, $E_{\infty}$ has the same order of magnitude as the initial relative errors of the components. We conclude that, in the vast majority of cases for $y_{0}$, the relative error of $\max\limits_{t\geq 0}\|\widetilde{y}(t)\|_{\infty}$ with respect to $\max\limits_{t\geq 0}\|y(t)\|_{\infty}$ is not much larger than the initial relative errors of the components.
To confirm this conclusion, consider
$y_{0}=(1,1)\quad\text{and}\quad\widetilde{y}_{0}=(1.01,0.99).$
The following table presents the signed absolute and relative errors of $\max\limits_{t\geq 0}\|\widetilde{y}(t)\|_{\infty}$ with respect to $\max\limits_{t\geq 0}\|y(t)\|_{\infty}$.
Absolute error Relative error
$a=50$ 0.1198-0.1198 0.0092-0.0092
$a=500$ 1.2450-1.2450 0.0099-0.0099
$a=5000$ 12.4949-12.4949 0.0100-0.0100
While the absolute error grows significantly with respect to the absolute errors $\pm 0.01$ of the components of $\widetilde{y}_{0}$, due to the large values attained by $\|\mathrm{e}^{tA}\|_{\infty}$ , the relative error, as expected, remains close to the relative errors $\pm 0.01$ of the components of $\widetilde{y}_{0}$.
9. Conclusion
The present paper explores how a perturbation of the initial value of the ODE (1) is propagated to the solution by examining the relative error, i.e., by comparing the perturbation of the solution to the solution itself. Considering the relative error, rather than the absolute error, provides a clearer perspective on the perturbation as the solution $y(t)$ evolves with time $t$: at any given time $t$, the perturbation of the solution $y(t)$ can be large (small) when compared to the initial value $y_{0}$, and simultaneously small (large) when compared to $y(t)$.
Understanding the behavior of the relative error of the perturbation can offer new insights and impact our approach to understanding the propagation of uncertainties. For example, the non-normality of the matrix $A$ appears to have a lesser effect on the relative error compared to the absolute error.
When we examine the absolute conditioning of the problem (2), the pointwise condition number is
$K_{\mathrm{abs}}(t)=\|\mathrm{e}^{tA}\|.$
Therefore, in the context of absolute error, a perturbation of the initial value grows in the worst-case scenario as a solution of $y^{\prime}(t)=Ay(t)$ grows in the worst-case scenario. There is nothing new: by knowing how the solutions grow, we also know how the perturbations grow, since the perturbations are themselves solutions. Moreover, the initial value $y_{0}$ does not play any role, since the pointwise condition number is independent of $y_{0}$.
On the other hand, when we consider the relative conditioning of the problem (2), the pointwise condition number is
$K_{\mathrm{rel}}(t,y_{0})=K(t,y_{0})=\frac{\|\mathrm{e}^{tA}\|\|y_{0}\|}{\|\mathrm{e}^{tA}y_{0}\|}.$
In the context of relative error, the perturbations do not grow in the same manner as the solutions, but in a completely different way: compare fact A) at page 1 with fact B) at page 17. Unlike the absolute conditioning, the initial value $y_{0}$ plays a role in the relative conditioning. Indeed, the time growth of $\|\mathrm{e}^{tA}\|$ can be mitigated by the time growth of $\|\mathrm{e}^{tA}y_{0}\|$.
These considerations represent the novelties in focusing on relative error rather than absolute error, offering new perspectives in linear dynamics, particularly in non-normal dynamics, when simulating the transient behavior of a real-world system and there is uncertainty in the initial value of the transient.
The present paper analyzes the asymptotic (long-time) behavior of the relative error of the perturbed solution. This asymptotic behavior also provides information on the non-asymptotic behavior. In fact, the strong experimental evidence included in the paper suggests that, in the vast majority of cases, the maximum values over time $t$ of the condition number $K(t,y_{0})$ and its asymptotic form $K_{\infty}(t,y_{0})$ share the same order of magnitude.
The main result of the paper can be summarized as follows. Consider the Euclidean norm as vector norm. In the vast majority of cases for the ODE (1), the relative error of the perturbed initial value is magnified in the perturbed solution, in the worst-case scenario for the initial perturbation, by a factor of the order of magnitude of
$\frac{\left\|w^{(1)}\right\|_{2}\left\|y_{0}\right\|_{2}}{|w^{(1)}y_{0}|},$ (46)
where the row vector $w^{(1)}$ is a left eigenvector of the matrix $A$ corresponding to the rightmost real or complex eigenvalue. The quantity in (46) is the asymptotic magnification factor.
Acknowledgements: the research was supported by the INdAM Research group GNCS (Gruppo Nazionale di Calcolo Scientifico).